### v3

# Asymptotically Unbiased Synthetic Control Methods by Distribution Matching

Abstract
Synthetic Control Methods (SCMs) have become an essential tool for comparative case studies. The fundamental idea of SCMs is to estimate the counterfactual outcomes of a treated unit using a weighted sum of the observed outcomes of untreated units. The accuracy of the synthetic control (SC) is critical for evaluating the treatment effect of a policy intervention; therefore, the estimation of SC weights has been the focus of extensive research. In this study, we first point out that existing SCMs suffer from an endogeneity problem, the correlation between the outcomes of untreated units and the error term of the synthetic control, which yields a bias in the treatment effect estimator. We then propose a novel SCM based on density matching, assuming that the density of outcomes of the treated unit can be approximated by a weighted average of the joint density of untreated units (i.e., a mixture model). Based on this assumption, we estimate SC weights by matching the moments of treated outcomes with the weighted sum of moments of untreated outcomes. Our proposed method has three advantages over existing methods: first, our estimator is asymptotically unbiased under the assumption of the mixture model; second, due to the asymptotic unbiasedness, we can reduce the mean squared error in counterfactual predictions; third, our method generates full densities of the treatment effect, not merely expected values, which broadens the applicability of SCMs. We provide experimental results to demonstrate the effectiveness of our proposed method. [^1] This study was presented at the Workshop on Counterfactuals in Minds and Machines at the International Conference on Machine Learning in July 2023 and at the International Conference on Econometrics and Statistics in August 2023.
1 Introduction
Synthetic control methods (SCMs, Abadie and Gardeazabal, 2003; Abadie et al., 2010) have gained attention as a crucial tool for program evaluation in comparative case studies across fields such as economics, statistics, and machine learning. According to Athey and Imbens (2017), SCMs are “arguably the most important innovation in the policy evaluation literature in the last 15 years” due to their “simplicity of the idea, and the obvious improvement over standard methods.”
SCMs posit a scenario in which multiple units exist, with one unit subjected to a policy intervention (treated unit) at a certain time period, while the rest remain untreated. The focus is on estimating the treatment effect of the treated unit, i.e., the difference between its observed factual outcome and unobserved counterfactual outcome. Untreated units are used to account for unobserved trends in the outcome over time that are unrelated to the effect of the policy intervention. The key insight is that an optimally weighted average of the untreated units, referred to as the synthetic control (SC) unit, provides a more appropriate estimate of the counterfactual than estimating the counterfactual using only the treated unit. See Figure 1 for an illustration. SCMs have been applied in a variety of settings, such as studying the effect of terrorism on GDP (Abadie and Gardeazabal, 2003) and the decriminalization of indoor prostitution (Cunningham and Shah, 2017).
Figure 1: Illustration of SCMs
For instance, Abadie et al. (2010) used SCM to study the treatment effect of a comprehensive tobacco control program implemented in California in 1988. Smoking rates in California decreased following the program’s implementation, but it was unclear if this decrease was a result of the program or a long-term trend. To investigate the treatment effect, annual per-capita cigarette sales data across several states that did not implement a similar tobacco control program were used as untreated units. They then applied SCM to estimate California’s counterfactual outcome: the smoking rate in California in 1989 had the tobacco control program not been implemented. For further details, also see Abadie (2021).
A significant challenge in SCMs is the typical assumption that the expected outcome of the treated unit can be represented by the weighted sum of untreated units. However, Ferman and Pinto (2021) points out that there exists a non-negligible asymptotic bias in the obtained estimator. According to Ferman and Pinto (2021), this asymptotic bias is caused by a kind of endogeneity in the linear regression model deduced from the assumption, due to a correlation between the outcomes of untreated units and the error term of the linear model. This problem is akin to endogeneity caused by measurement error (Greene, 2003). We refer to this endogeneity as implicit endogeneity.
To address the aforementioned issue, we propose to assume a mixture model instead of a linear relationship among expected outcomes. Specifically, we assume that the density of the outcome of the treated unit can be approximated by a linear combination of those of the untreated units. This mixture model assumption also implies that each moment of the outcome of the treated unit can be represented by a linear combination of those of the untreated units. Therefore, we can estimate the SC weights by matching all moments between the outcome of the treated unit and those of the untreated units. We refer to our method as the Density Matching SCM (DMSCM). Unlike existing methods, our proposed estimator is asymptotically unbiased. We validate the efficacy of our proposed method through simulation studies and empirical analysis.
Our proposed method corresponds to a simplification of the distributional SCMs by Gunsilius (2023). Although Gunsilius (2023) does not explicitly point out the implicit endogeneity problem, it observes that their result does not coincide with the result under the standard SCMs. This is because under the mixture model assumption, while Gunsilius (2023) returns a consistent estimator for the treatment effect, existing methods do not. We connect the distributional SCMs to the asymptotic bias problem and propose a simplified estimation method for SCMs by distribution matching.
In summary, our contributions are as follows:
  • Reconsidering the assumptions necessary for treatment effect estimation in SCMs.
  • Proposing a density-matching method for asymptotically unbiased SCMs.
  • Simplifying the density-matching method.
1.1 Related Work
Following the pioneering work by Abadie and Gardeazabal (2003), numerous studies have advanced the use of SCMs in comparative studies (Abadie et al., 2010, 2015). These SCMs, including their variants such as those by Doudchenko and Imbens (2016) and Ben-Michael et al. (2019), are based on minimizing the pre-treatment sum of squared errors between the outcome of the treated unit and the weighted sum of those of the untreated units, with some regularization and constraints regarding the weights.
Gunsilius (2023) explores distribution SCMs (DiSCo), which estimate SC weights to match the distribution of the outcomes of the treated unit and the weighted sum of the distributions of the untreated units. [^2] The method of Gunsilius (2023) is called DiSCo by Dijcke (2024). His method is based on minimizing the Wasserstein distance, which can be simplified as matching the quantiles in the setting of SCMs. Their assumption about identification can be interpreted as a variant of mixture models.
From a theoretical viewpoint, Abadie et al. (2010) demonstrate that, if there are weights that provide a perfect pre-treatment fit, then the bias of the SC estimator is bounded by a function that converges to zero asymptotically as the length of the pre-treatment periods $T_{0}$ increases, even when the number of untreated units $J$ is fixed. However, Ferman and Pinto (2021) finds that if the pre-treatment fit is imperfect, then the conventional SCMs do not yield a consistent estimator of the treatment effect due to the bias caused by the correlation between the outcomes of the untreated units and the error term in the posited linear models. We review this problem in Section 4 and name it as the implicit endogeneity.
To address this bias issue, existing studies have proposed several approaches. Ferman (2021) demonstrates that we can obtain an asymptotically unbiased estimator under the regime where the number of untreated units increases ($J\to\infty$) as well as the length of the pre-intervention periods $T_{0}$ also increases.
The use of mixture models has a close relationship with linear factor models, assumed in existing studies (Abadie, 2021). We justify the use of the former by employing arguments from Shi et al. (2022) and Nazaret et al. (2023). Our proposed method can be considered a new interpretation of existing SCMs. By reinterpreting existing SCMs through density matching, in addition to addressing the implicit endogeneity problem, we address two problems in the SCM literature. First, our proposed method improves the pre-treatment fit by debiasing the estimator. Second, our method generates full counterfactual densities of the treated unit, rather than merely the expected value. As a result, we can conduct statistical inference on the distribution, or any function of the distribution, such as quantiles (Gunsilius, 2023), Lorenz curves (Gastwirth, 1971), and expected social welfare (Abadie, 2002).
An excellent illustration of the relevance of estimating distribution functions is Ropponen (2011), who applies the changes-in-changes estimator by Athey and Imbens (2006) to estimate the effect of minimum wage changes on employment levels in order to reconcile the classical conflicting results about the minimum wage debate in Card and Krueger (2000) and Neumark and Wascher (2000). Spiess et al. (2023) investigates high-dimensional linear regression models in SCMs.
The exploration of identification assumptions has been attempted in the existing literature. Ferman and Pinto (2021) conducts a comprehensive investigation to discern the conditions under which SCMs can accurately predict counterfactual outcomes.
This study also answers an important open question raised by Gunsilius (2023). Gunsilius (2023) discusses that an optimal weight obtained by the density matching differs from that obtained by the classical least squares approach. Gunsilius (2023) conjectures that the classical least squares approach estimates a weight to match the first moment, while the density matching approach estimates a weight to match all moments. However, we point out that the classical least squares approach estimates a weight not only to match the first moment but also tends to incur bias due to the implicit endogeneity.
Concerning the assumptions about SCMs, Wan et al. (2018) points out that the SCMs essentially assume mixture models, which distinguish SCMs from traditional panel data analysis.
On another note, Shi et al. (2022) introduces fine-grained models to investigate hidden assumptions inherent to SCMs, which also explores a connection between the linear combination of expected outcomes and the linear factor models. Nazaret et al. (2023) extends these findings into the realm of model misspecification. Their formulation is also based on the mixture models.
Independently of us, Fry (2024) also discusses the endogeneity problem in SCMs as a cause of the asymptotic bias pointed out by Ferman and Pinto (2021). [^3] Note that his study became public on arXiv after our study became public on arXiv and at the ICML workshop, and we and he conducted our studies completely independently. While our and his formulations about the endogeneity are similar, the approaches are significantly different. We address the problem from the viewpoint of density matching, which is an approach similar to Gunsilius (2023). In contrast, Fry (2024) uses an instrumental variable method by employing outcomes of untreated units that are not used in SCMs as instrumental variables.
Table 1: Comparison of SCMs. See the following sections for the definitions of variables and methods.
Method Assumption Asymptotic unbiasedness Remark
Ours DMSCM $F_{Y^{N}_{0,t}}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F_{Y^{N}_{j,t}}(y)\ \ \ \forall y\in\mathbb{R}$ Unbiased
Abadie and Gardeazabal (2003) Constrained least squares $\mathbb{E}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}[Y^{N}_{j,t}]$ Biased
Ferman (2021) Constrained least squares $\mathbb{E}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}[Y^{N}_{j,t}]$ Unbiased $J\to\infty$
Gunsilius (2023) DiSCo $F^{-1}_{Y^{N}_{0,t}}(q)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F^{-1}_{Y^{N}_{j,t}}(q)\ \ \ \forall q\in(0,1)$ Unbiased
Fry (2024) IV method (GMM) $\mathbb{E}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}[Y^{N}_{j,t}]$ Unbiased Use IVs
1.2 Organizations
In Section 2, we formulate the problem. In Section 3, we review an idea of SCMs. In Section 4, we point out the problem of implicit endogeneity. Then, in Section 5, we propose our method under the assumption of mixture models. In Section 6, we investiate the asymptotic property of our proposed method. In Section 7, we present a statistical inference method for our proposed method. In Section 8, we discuss related topics and existing studies. We numerically investigate the performance of the proposed estimators through extensive simulations in Section 9.
2 Problem Setting
Suppose that there are $J+1$ units, indexed by $j\in\mathcal{J}\coloneqq\{0,1,2,\dots,J\}$, and a time series indexed by $t\in\mathcal{T}=\{1,2,\dots,T\}$ where $3\leq T\in\mathbb{N}$. Without loss of generality, we assume that the first unit ($j=0$) is the treated unit, i.e., the unit affected by the policy intervention of interest. A set of potential comparisons, $j\in\mathcal{J}^{U}\coloneqq\mathcal{J}\backslash\{0\}=\{1,2,\dots,J\}$, consists of untreated units not affected by the intervention. We also assume that our data span $T$ periods and that the intervention occurs at $t=T_{0}$, such that $2\leq T_{0}<T$. Observations at period $t\in\mathcal{T}_{0}\coloneqq\{1,2,\dots,T_{0}\}$ are from before the intervention, and observations at period $t\in\mathcal{T}_{1}\coloneqq\mathcal{T}\backslash\mathcal{T}_{0}$ are from after the intervention. Let $T_{1}\coloneqq|\mathcal{T}_{1}|=T-T_{0}$.
2.1 Potential Outcomes
First, following the Neyman-Rubin causal model (Neyman, 1923; Rubin, 1974), we define the potential outcomes of the policy intervention. For each unit $j\in\mathcal{J}$ and period $t\in\mathcal{T}$, there are two outcomes of interest, $\left(Y^{I}_{j,t},Y^{N}_{j,t}\right)\in\mathbb{R}^{2}$, corresponding to scenarios with and without the intervention. For each $j\in\mathcal{J}$ and $t\in\mathcal{T}$, let $F^{I}_{j,t}(y)$ and $F^{N}_{j,t}(y)$ be the cumulative distribution functions (CDFs) for $Y^{I}_{j,t}$ and $Y^{N}_{j,t}$, respectively.
2.2 Observations
We then define our observations. For each unit $j\in\mathcal{J}$, we can only observe one of the outcomes corresponding to the actual intervention; that is, we observe $Y_{j,t}\in\mathbb{R}$ such that
$Y_{0,t}=\begin{cases}Y^{N}_{0,t}&\forall t\in\mathcal{T}_{0}\\ Y^{I}_{0,t}&\forall t\in\mathcal{T}_{1}\end{cases},\qquad\mathrm{and}\qquad Y_{j,t}=Y^{N}_{j,t}\qquad\forall j\in\mathcal{J}^{U},\forall t\in\mathcal{T}.$
We aim to estimate the effect of the policy intervention by predicting $Y^{I}_{0,t}$ from $(Y^{N}_{j,t})_{j\in\mathcal{J}^{U}}$ and by comparing this prediction to the observed $Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$. We will define the treatment effect more formally in the following section.
2.3 Treatment Effect
In SCMs, there are several ways to define the treatment effect $\tau_{0,t}\in\mathbb{R}$ for each $t\in\mathcal{T}_{1}$ of the intervention for the treated unit. First, for each $t\in\mathcal{T}_{1}$, we define the treatment effect on the treated unit $j=0$ as
$\tau_{0,t}\coloneqq\mathbb{E}\big{[}Y^{I}_{0,t}\big{]}-\mathbb{E}\big{[}Y^{N}_{0,t}\big{]}.$
Another definition is based on the following linear factor model:
Definition 2.1.
We say that $(Y^{I}_{j,t},Y^{N}_{j,t})$ follows the linear factor model if:
$Y^{N}_{j,t}=c_{j}+\delta_{t}+\lambda_{t}\mu_{j}+\epsilon_{j,t},\quad\mathbb{E}_{j,t}[\epsilon_{j,t}]=0,\quad\forall t\in\mathcal{T},\quad\forall j\in\mathcal{J},$ (1)
$Y^{I}_{j,t}=\tau_{j,t}+Y^{N}_{j,t}.$
where $\delta_{t}$ is an unobserved common factor with constant factor loadings across units, $c_{j}$ is an unknown time-invariant fixed effect, $\lambda_{t}$ is a $(1\times F)$ vector of unobserved common factors, $\mu_{j}$ is a $(F\times 1)$ vector of unknown factor loadings, and the error terms $\epsilon_{j,t}$ are unobserved idiosyncratic shocks. Here, while $c_{j}$, $\mu_{j}$, and $\tau_{0,t}$ are non-random variables, $\delta_{t}$, $\lambda_{t}$, and $\epsilon_{j,t}$ are random variables.
Under the linear factor model, the treatment effect is given by $\tau_{0,t}=Y^{I}_{0,t}-Y^{N}_{0,t}$, a non-random variable by definition of the factor model. Therefore, the expression $\tau_{0,t}\coloneqq\mathbb{E}\big{[}Y^{I}_{0,t}\big{]}-\mathbb{E}\big{[}Y^{N}_{0,t}\big{]}$ also holds under the factor model.
Comparative case studies aim to evaluate $\tau_{0,t}$ by predicting $Y^{N}_{0,t}$ for $t>T_{0}$, which is the counterfactual outcome that would have been observed for the treated unit in the absence of the intervention. By definition, $Y^{N}_{0,t}$ is a counterfactual outcome because of the intervention after $t=T_{0}$. We usually use one unit or a small number of units that have similar characteristics to the treated unit at the time of the intervention but are not exposed to the intervention. However, when the data consist of a few units, such as regions or countries, it is often difficult to find a single untreated unit that provides a suitable comparison for the unit treated by the policy intervention of interest.
Distributional Treatment Effects. In program evaluation, we are often interested not only in the scalar treatment effect but also in the distributions of outcomes (Maier, 2011) or their functionals, such as the quantile treatment effects (QTEs) and expected social welfare (Abadie, 2002). In this study, we consider estimating $p^{I}_{0,t}(y)$ and $p^{N}_{0,t}(y)$ for $t\in\mathcal{T}_{1}$, where outcomes corresponding to $p^{I}_{0,t}(y)$ are unobservable. The treatment effect is then captured by $p^{I}_{0,t}(y)$ and $p^{N}_{0,t}(y)$, or functionals of these distributions, such as QTEs and expected social welfare. Depending on the application, we may adjust the target density, for instance, $\overline{p}^{N}_{0}(y)=\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p^{N}_{0,t}(y)$. We refer to the treatment effects related to these distributions as distributional treatment effects (Park et al., 2021; Kallus and Oprescu, 2022; Chikahara et al., 2022). In the context of SCMs, Gunsilius (2023) and Chen (2020) have estimated the treatment effect.
3 Recap of SCMs
This section reviews the original SCM proposed by Abadie and Gardeazabal (2003). To estimate the treatment effect $\tau_{0,t}$, we predict the counterfactual outcome $Y^{N}_{0,t}$ by using the weighted sum of $\{Y^{N}_{j,t}\}^{J}_{j=1}$ for $t\in\mathcal{T}_{1}$; that is, SCMs predict $Y^{N}_{0,t}$ as
$\widehat{Y}^{N}_{0,t}\coloneqq w_{0}+\sum_{j\in\mathcal{J}^{U}}w_{j}Y^{N}_{j,t},$
where $w_{0}\in\mathbb{R}$ and $w_{j}\in\mathbb{R}$ $(j\in\mathcal{J}^{U})$ are the weights. We refer to these weights as SC weights.
Several methods exist for estimating the SC weights. In the SCM proposed by Abadie and Gardeazabal (2003), SC weights are determined by minimizing the squared error between $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{1t},\dots,Y^{N}_{J,t}$. Let us define
$\Delta^{J}\coloneqq\{\bm{w}=(w_{1}\ w_{2}\ \dots\ w_{J})^{\top}\in[0,1]^{J}\mid\sum_{j\in\mathcal{J}^{U}}w_{j}=1\}.$
Then, Abadie and Gardeazabal (2003) estimates the SC weights as
$\bm{w}^{\mathrm{Abadie}}=\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{0,t}-\sum_{j\in\mathcal{J}^{U}}w_{j}Y^{N}_{j,t}\right)^{2},$
The counterfactual outcome is then predicted as
$\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}=\widehat{w}^{\mathrm{Abadie}}_{0}+\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{Abadie}}_{j}Y^{N}_{j,t}.$
Using the prediction $\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}$, we estimate the treatment effect as $\widehat{\tau}^{\mathrm{Abadie}}_{t}=Y^{I}_{0,t}-\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}$.
The literature on inference for SCMs has grown significantly, extending beyond the original proposals of Abadie and Gardeazabal (2003) and Abadie et al. (2010). Notable contributions include works by Li (2020), Shaikh and Toulis (2021), Cattaneo et al. (2021), and Chernozhukov et al. (2019). Doudchenko and Imbens (2016) proposes estimating the SC weight as
$\widetilde{\bm{w}}^{\mathrm{DI}}=\operatorname*{arg\,min}_{\widetilde{\bm{w}}=(w_{0}\ w_{1}\ \cdots\ w_{J})^{\top}\in\mathbb{R}^{J+1}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{0,t}-w_{0}-\sum_{j\in\mathcal{J}^{U}}w_{j}Y_{j,t}\right)^{2}+\lambda_{1}\|\widetilde{\bm{w}}\|_{1}+\lambda_{2}\|\widetilde{\bm{w}}\|_{2},$
where $\lambda_{1}>0$ and $\lambda_{2}>0$ are regularization parameters. The parameter space for $\widetilde{\bm{w}}$ can be restricted in various ways, such as setting $w_{0}=0$ and $\bm{w}=(w_{1}\ \cdots\ w_{J})^{\top}\in\Delta^{J}$.
4 Identification and Endogeneity
This section addresses the identification problem of treatment effects from the perspective of endogeneity.
4.1 Linear Models in Expected Outcomes
We reconsider the underlying assumptions in SCMs. Existing studies often posit some linear relationship between $Y^{N}_{0,t}$ and $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$. One such relationship assumes the existence of $\bm{w}^{*}\in\Delta^{J}$ (Abadie et al., 2010):
$Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}.$
This assumption, known as the perfect pre-treatment fit (Ferman and Pinto, 2021), implies that the stochastic process of $Y^{N}_{0,1},\dots,Y^{N}_{0,T}$ can be completely replicated by $\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$. In linear factor models, $Y^{N}_{j,t}=c_{j}+\delta_{t}+\lambda_{t}\mu_{j}+\epsilon_{j,t}$, and this assumption might imply $\epsilon_{0,t}=\sum_{j\in[J]}w^{*}_{j}\epsilon_{j,t}$, which is considered unrealistic (note that this is a specific case and not a necessary condition for $Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$).
A more natural assumption is that each term in the linear factor models of $Y^{N}_{0,t}$ can be replicated by the weighted sum of corresponding terms in the models of $Y^{N}_{0,1},\dots,Y^{N}_{0,T}$. A weaker assumption still is a linear relationship between $\mathbb{E}_{0,t}\left[Y^{N}_{0,t}\right]$ and $(\mathbb{E}_{j,t}\left[Y^{N}_{j,t}\right])_{j\in\mathcal{J}^{U}}$, which includes the replication of each term in linear factor models as a specific case.
Assumption 4.1.
There exists a set $\widetilde{\Phi}\subset\Delta^{J}$ such that for all $\bm{w}^{*}=(w^{*}_{j})_{j\in\mathcal{J}^{U}}\in\widetilde{\Phi}$, it holds that
$\mathbb{E}_{0,t}\left[Y^{N}_{0,t}\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[Y^{N}_{j,t}\right],$ (2)
where $\Delta^{J}\coloneqq\{(w_{1},\dots,w_{J})\in\mathbb{R}^{J}\mid w_{j}\geq 0\ \forall j\ \mathrm{and}\ \sum_{j\in\mathcal{J}^{U}}w_{j}=1\}$.
Compared to Assumption 4.1, assuming $Y_{0,t}=\sum_{j\in\mathcal{J}^{U}}Y_{j,t}$ is a stronger and less convincing assumption than (2) since $Y_{j,t}$ is a random variable. Therefore, in this study, we discuss SCMs based on (2). Ferman and Pinto (2021) explicitly considers Assumption 4.1.
4.2 Asymptotic Bias in Linear Factor Models
We consider the linear factor model as defined in Definition 2.1, along with Assumption 4.1. However, under the linear factor model, we adopt the following more specific assumption:
Assumption 4.2.
Under the linear factor model from Definition 2.1, there exists a set $\widetilde{\Phi}\subset\Delta^{J}$ such that for all $\bm{w}^{*}=(w^{*}_{j})_{j\in\mathcal{J}^{U}}\in\widetilde{\Phi}$, the following conditions hold:
$c_{0}=\sum_{j\in[J]}w^{*}_{j}c_{j}\quad\text{and}\quad\mu_{0}=\sum_{j\in[J]}w^{*}_{j}\mu_{j}\ \ \forall t\in\mathcal{T},\ \forall j\in\mathcal{J}.$
Assumption 4.2 is a specific case of Assumption 4.1; that is, if Assumption 4.1 holds, then Assumption 4.2 also holds under the linear factor model defined in Definition 2.1.
Under the linear factor model with Assumption 4.2, Ferman and Pinto (2021) finds that the SC estimator $\widehat{\tau}^{\mathrm{Abadie}}_{t}$ is biased and not consistent for $\tau_{0,t}$ when $c_{j}$ is non-zero. We further make the following assumption for the linear factor models.
Assumption 4.3 (Common and idiosyncratic shocks).
For each $j\in\mathcal{J}$, the linear factor models 2.1 satisfy the following conditions as $T_{0}\to\infty$:
  • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\lambda_{t}\xrightarrow{\mathrm{p}}0$.
  • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\epsilon_{j,t}\xrightarrow{\mathrm{p}}0$.
  • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\lambda^{\top}_{t}\lambda_{t}\xrightarrow{\mathrm{p}}\Omega_{0}$, where $\Omega_{0}$ is positive semidefinite.
  • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\epsilon_{j,t}\epsilon^{\top}_{j,t}\xrightarrow{\mathrm{p}}\sigma^{2}_{j,\epsilon}I_{J+1}$, where $\sigma^{2}_{j,\epsilon}\geq 0$ is constant.
  • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\epsilon_{j,t}\lambda_{t}\xrightarrow{\mathrm{p}}0$.
Ferman and Pinto (2021) then demonstrates that the SCM proposed by Abadie and Gardeazabal (2003) does not yield a consistent estimator for $\tau_{0,t}$ under these conditions.
Proposition 4.4 (Proposition 1 in Ferman and Pinto (2021)).
Under Assumptions 4.34.2, it holds that
$\bm{w}^{\mathrm{Abadie}}\xrightarrow{\mathrm{p}}\overline{\bm{w}}^{\mathrm{Abadie}}\coloneqq\left(\overline{w}^{\mathrm{Abadie}}_{j}\right)_{j\in\mathcal{J}^{U}}$
as $T_{0}\to\infty$, where $(c_{0},\mu_{0})\neq(\sum_{j\in\mathcal{J}^{U}}c_{j}\overline{w}^{\mathrm{Abadie}}_{j},\sum_{j\in\mathcal{J}^{U}}\mu_{j}\overline{w}^{\mathrm{Abadie}}_{j})$, unless $\sigma^{2}_{\epsilon}=0$ or $\widetilde{\Phi}\cap\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\{\bm{w}^{\top}\bm{w}\}\neq\emptyset$. Moreover, for $t\in\mathcal{T}_{1}$, it holds that
$\widehat{\tau}^{\mathrm{Abadie}}_{0,t}$ $\xrightarrow{\mathrm{p}}\tau_{0,t}+\lambda_{t}\left(\mu_{0}-\sum_{j\in\mathcal{J}^{U}}\mu_{j}\overline{w}^{\mathrm{Abadie}}_{j}\right)+\left(c_{0}-\sum_{j\in\mathcal{J}^{U}}c_{j}\overline{w}^{\mathrm{Abadie}}_{j}\right)+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}\epsilon_{j,t}\overline{w}^{\mathrm{Abadie}}_{j}\right)$
as $T_{0}\to\infty$.
To mitigate this issue, Ferman and Pinto (2021) proposes the demeaned SC estimator
$\widehat{\bm{w}}^{\mathrm{FP}}=\operatorname*{arg\,min}_{\widehat{\bm{w}}^{\mathrm{FP}}\in\Delta^{J}}\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left\{\left(Y^{N}_{0,t}-\overline{Y}^{N}_{0}\right)-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)\right\}^{2},$
where $\overline{Y}^{N}_{j}=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}Y_{j,t}$. They estimate $\tau_{0,t}$ as $\widehat{\tau}^{\mathrm{FP}}_{0,t}=Y^{I}_{0,t}-\overline{Y}^{N}_{0}-\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{FP}}_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)$. They show the following result.
4.3 Implicit Endogeneity
In this section, we interpret the asymptotic bias identified by Ferman and Pinto (2021) from the perspective of endogeneity.
To clarify this issue, we consider a simpler case using a least-squares estimator rather than $\bm{w}^{\mathrm{Abadie}}$:
$\widehat{\bm{w}}^{\mathrm{LS}}=\operatorname*{arg\,min}_{\bm{w}\in\mathbb{R}^{J}}\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{0,t}-\sum_{j\in\mathcal{J}^{U}}w_{j}Y_{j,t}\right)^{2}.$
We demonstrate that $\widehat{\bm{w}}^{\mathrm{LS}}$ is asymptotically biased due to endogeneity. Define $\epsilon_{j,t}=Y_{j,t}-\mathbb{E}_{j,t}\left[Y_{j,t}\right]$. Assuming that (2) holds, the following linear regression model is applicable:
$Y_{0,t}=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}Y_{j,t}+\nu_{t},\quad\text{where}\quad\nu_{t}=-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}+\epsilon_{0,t}.$
Here, $Y_{0,t}=\mathbb{E}_{0,t}[Y_{0,t}]+\epsilon_{0,t}=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[Y_{j,t}\right]+\epsilon_{0,t}=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}Y_{j,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}+\epsilon_{0,t}$. Note that $\mathbb{E}\left[\nu_{t}Y_{j,t}\right]\neq 0$, implying endogeneity due to a correlation between the regressor and the error term. Endogeneity is known to introduce bias in the least-squares estimator. Define $Q^{*}$ as a $(J\times J)$-matrix whose $(i,j)$-element is $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\mathbb{E}_{i,t}[Y_{i,t}]\mathbb{E}_{j,t}[Y_{j,t}]$. Let $\Sigma$ be a $J\times J$ diagonal matrix whose $(j,j)$-element is $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\sigma^{2}_{j,t}$.
Theorem 4.5.
As $T_{0}\to\infty$, $\widehat{\bm{w}}^{\mathrm{LS}}\xrightarrow{\mathrm{p}}\overline{\bm{w}}^{\mathrm{LS}}$, where
$\overline{\bm{w}}^{\mathrm{LS}}=\bm{w}^{*}-\left(Q^{*}+\Sigma\right)^{-1}\Sigma\bm{w}^{*}.$
This theorem is derived from Section 5.6 of Greene (2003). The result implies that under the assumption of (2), there is an asymptotic bias in $\widehat{\bm{w}}^{\mathrm{LS}}$; that is, least-squares-type estimators do not converge to the true value $\bm{w}^{*}$.
This section highlights the identification of treatment effects and the problem of implicit endogeneity. Though the estimator $\bm{w}^{\mathrm{Abadie}}$ is widely used in existing studies, we find that it suffers from implicit endogeneity. This issue has been explicitly or implicitly discussed in previous studies (Chernozhukov et al., 2019; Ferman and Pinto, 2021; Ferman, 2021). Independently of our work, Fry (2024) explicitly addresses the bias as an endogeneity problem and applies the instrumental variable method to resolve the issue.
5 SCM by Density Matching
As discussed in Section 4, Assumption 4.1 is considered weak or incompatible for identification of $\bm{w}^{*}$ using least squares. Therefore, we consider adopting another assumption for identification. Specifically, we propose that the density of $Y^{N}_{0,t}$ can be replicated by a weighted sum of the densities of $Y^{N}_{j,t}$ and introduce a SCM based on the density-matching approach.
5.1 Identification and Estimation Strategy
Section 4 and the results in Ferman and Pinto (2021) imply that combining the assumption $\mathbb{E}_{0,t}\left[Y_{0,t}\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[Y_{j,t}\right]$ with least squares estimators does not yield consistent estimators for $\bm{w}^{*}$. Hence, we consider another more realistic and theoretically favorable assumption.
We begin with the premise that the distribution of the target unit’s outcome, $Y^{N}_{0,t}$, is a linear combination of those of the control units’ outcomes, $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$. For simplicity, in this subsection, we assume that $Y^{N}_{j,t}$ is a continuous random variable with a probability density function (PDF). We then consider a scenario where the density of $Y^{N}_{0,t}$ can be replicated by a weighted sum of the densities of $Y^{N}_{j,t}$:
$p^{N}_{0,t}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p^{N}_{j,t}(y)\qquad\forall t\in\mathcal{T}.$ (3)
We discuss how to estimate SC weights that satisfy (3). One approach is to employ the optimal-transport argument, as used by Gunsilius (2023), which estimates SC weights by minimizing the Wasserstein distance. However, in this study, we provide a simpler algorithm for estimating SC weights since the method of Gunsilius (2023) requires a procedure that is more complex than the original SCMs.
Because the mixture model (3) holds, we establish the following equation:
$\mathbb{E}_{0,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{0,t}\Big{)}\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{j,t}\Big{)}\right]\quad\forall t\in\mathcal{T},\ \ \forall\xi\in\mathbb{R},$ (4)
where $\sqrt{-1}$ denotes the imaginary number. From this equation, we estimate $w^{*}$ using the following methodology. We consider the upper bound of the difference between the expected exponential transforms of $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{j,t}$:
$\int_{\xi\in(-h,h)}\left|\mathbb{E}_{0,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{0,t}\Big{)}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{j,t}\Big{)}\right]\right|\mathrm{d}\xi$
$=\int_{\xi\in(-h,h)}\left|\sum_{\gamma\in\{0,1,2,\dots,G\}}\frac{\sqrt{-1}^{\gamma}\xi^{\gamma}}{\gamma!}\left(\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right)+\epsilon|\xi^{G}|\right|\mathrm{d}\xi$
$\leq\int_{\xi\in(-h,h)}\left(\sum_{\gamma\in\{1,2,\dots,G\}}\frac{|\sqrt{-1}^{\gamma}||\xi|^{\gamma}}{\gamma!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|+\epsilon|\xi^{G}|\right)\mathrm{d}\xi$
$=\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|+\frac{2\epsilon}{G+1}h^{G+1}.$
This formula implies that by minimizing $\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|+\frac{2\epsilon}{G+1}h^{G+1}$, we can accurately estimate $\bm{w}^{*}$ such that $\int_{\xi in(-h,h)}\left|\mathbb{E}_{0,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{0,t}\Big{)}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{j,t}\Big{)}\right]\right|\mathrm{d}\xi=0$. This estimation strategy can be interpreted as a variant of the method of moments, where we estimate SC weights to match all moments between the treated unit and the weighted sum of the untreated units.
5.2 Core Assumptions
We make the following core assumption for identification of the SC weights.
Assumption 5.1.
The outcomes $(Y^{N}_{j,t})_{j\in\mathcal{J},t\in\mathcal{T}}$ are bounded.
Assumption 5.2.
There exists $\widetilde{\Phi}^{*}\in\Delta^{J}$ such that for all $\bm{w}^{*}\in\widetilde{\Phi}^{*}$, it holds that
$F_{Y^{N}_{0,t}}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F_{Y^{N}_{j,t}}(y)\quad\forall y\in\mathbb{R}.$
Assumption 5.1 allows us to guarantee the existence of the $G$-th moment of $Y^{N}_{j,t}$ for any $G\in\mathbb{N}$. Assumption 5.2 also has been utilized in existing studies such as Wan et al. (2018), Gunsilius (2023) and Shi et al. (2022). Wan et al. (2018) points out that the SCMs are based on Assumption 5.2. Gunsilius (2023) assumes Assumption 5.2 and estimates $\bm{w}^{*}$ by using minimizing the Wasserstein distance. Shi et al. (2022) investigates the underlying assumptions in the classical SCMs and then find that the combination of fine-grained models and Assumption 5.2 justifies the standard SCMs. We introduce the fine-grained models proposed by Shi et al. (2022) in Section 8.1.
Although the mixture model and linear factor models are different, they are closely related, as pointed out by Shi et al. (2022) and Nazaret et al. (2023). In terms of fine-grained models, Shi et al. (2022) identifies implicit assumptions underlying linear factor models, and we argue that the result justifies the use of a mixture model in the context of linear factor models. Furthermore, Nazaret et al. (2023) independently report this finding. For further details, see Section 8.1.
5.3 Estimation of SC Weights
Under the assumption of mixture models, we estimate the SC weight to match the characteristic functions (CHFs). We refer to our method as the Density Matching SCM (DMSCM).
Let $h\in(0,1)$ be a parameter determined and fixed by us. Then, we estimate SC weights as
$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\coloneqq\left(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0},G}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right)\right|.$ (5)
By using the estimate $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0}}$, we can estimate the treatment effect and distributional treatment effect.
Treatment effect estimation.
Once estimating the weights, as well as the standard SCMs, we can estimate the treatment effect as
$\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0},G}Y_{j,t}.$ (6)
We refer to this estimator as the Density Matching SCM (DMSCM) estimator. In Algorithm 1, we summarize our proposed DMSCM estimator with the D2MSCM estimator introduced in the following section.
Distributional treatment effect estimation.
We can also estimate the distributional treatment effect as the counterfactual distribution of $p_{0,t}(y)$ by a bootstrap method. We re-sample $L>1$ observations $\{Y^{l,\dagger}_{0}\}^{L}_{l=1}$ from $\{Y_{0,t},Y_{1,t},\dots,Y_{J,t}\}_{t\in\mathcal{T}_{1}}$ as follows: for each $l\in\mathcal{L}\coloneqq\{1,2,\dots,L\}$, (i) for each $j\in\mathcal{J}$, we re-sample $Y^{l,\dagger}_{j}$ from $\{Y_{j,t}\}_{t\in\mathcal{T}_{1}}$; (ii) we choose $Y^{l,\dagger}_{0}$ from $(Y^{l,\dagger}_{1},\dots,Y^{l,\dagger}_{J})$ with probability $(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0}})_{j\in\mathcal{J}^{U}}$. In this case, our focus is on $\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p_{0,t}(y)$.
Algorithm 1 DMSCM.
  Parameter: The number of moments $G$.
  Obtain trained weights $\widehat{\bm{w}}^{\mathrm{DMSCM}}$ as (5).
  Estimate the treatment effect as (6).
6 Convergence Analysis
We provide the consistency results about our proposed SC weight estimator. The following theorem describes the results including the consistency for the treatment effect $\tau_{0}$.
Assumption 6.1.
Assume that the density function of $\varepsilon_{j,t}$ exists and is bounded for each $j\in\mathcal{J}$ and $t\in\mathcal{T}_{0}$, and that the stochastic process $\{\varepsilon_{j,t}\}^{T_{0}}_{t=1}$ is stationary, strongly mixing, with sum of mixing coefficient bounded by $M$.
Theorem 6.2.
Suppose that Assumptions 5.15.2 and 6.1 hold. Fix $h\in(0,1)$. Define
$\widetilde{\Phi}^{\dagger}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|$
Then, it holds that $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}$ as $T_{0}\to\infty$. Additionally, we have $\bm{w}^{*}_{G}\to\bm{w}^{*}\in\widetilde{\Phi}^{*}$ as $G\to\infty$ and
$\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}\xrightarrow{\mathrm{p}}\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right),$
as $T_{0}\to\infty$ and $G\to\infty$.
The proof is shown in Appendix A. Note that this result can be applied for $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ as a special case. Thus, $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ and $\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}$ are asymptotically unbiased under each assumption.
7 Inference
This section presents methods for statistical inference using our proposed DMSCM estimator.
7.1 Conformal Inference on the Treatment Effect
We consider statistical inference on the treatment effect using our proposed DMSCM and D2SCM to test a sharp null hypothesis, $H_{0}:\tau_{0,t}=\alpha_{t}$ for some $t\in\mathcal{T}_{t}$ and $\bm{\alpha}=(\alpha_{t})_{t\in\mathcal{T}_{1}}$. Here, we focus on the conformal inference of Chernozhukov et al. (2021), which has been utilized in existing studies, such as Ben-Michael et al. (2021) and Ferman and Pinto (2021).
Our goal is to establish a confidence interval whose $p$-value $\widehat{p}$ satisfies that for any $a\in(0,1)$,
$\mathrm{Pr}\left(\widehat{p}\leq a\right)\to a$ (7)
as $T_{0}\to\infty$ and $T_{1}$ is fixed. To achieve this goal, the conformal inference takes the following three steps. First, for the sharp null hypothesis
$H_{0}:\tau_{0,t}=\alpha_{t},$
we create adjusted post-treatment outcomes for the treated unit $Y^{N\sharp}_{0,t}=Y^{I}_{0,t}-\alpha_{t}$ and add it to the original dataset as $\left(Y_{0,1},\dots,Y_{0,T_{0}},Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$. Then, we apply our proposed method to the extended datasets consisting of
$\left(Y_{0,1},Y_{0,2},\dots,Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$
and $(Y_{j,1})_{t\in\mathcal{T},j\in\mathcal{J}^{U}}$ to obtain adjusted weights $\widehat{\bm{w}}(\alpha)$ and a corresponding predictor of the counterfactual outcome, denoted by $\widehat{Y}^{N}_{0,t}(\bm{\alpha})$. Finally, we compute the $p$-value by assessing whether the adjusted residual “conforms” with the pretreatment residuals. Let $u_{t}(\bm{\alpha})$ be the residual defined as $u_{t}(\bm{\alpha})=Y_{0,t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{0}$ and $u_{t}(\bm{\alpha})=Y^{I}_{0,t}-\alpha_{t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{1}$. For $j\in\mathcal{T}$, let $\pi_{j}$ be a permutation of indexes defined as $\pi_{j}(t)=t+j$ if $t+j\leq T$ and $\pi_{j}(t)=t+j-T$ otherwise. Let $\Pi=\{\pi_{1},\dots,\pi_{T-1}\}$. Then, the $p$-value is defined as $\widehat{p}(\bm{\alpha})=1-\widehat{F}(\bm{\alpha})$, where
$\widehat{F}(\bm{\alpha})=\frac{1}{T}\sum_{j\in\mathcal{T}}\mathbbm{1}\left[\frac{1}{T}\sum_{t\in\mathcal{T}}\left|u_{t}(\bm{\alpha})\right|<\frac{1}{T}\sum_{\pi_{j}\in\mathcal{T}}\left|u_{\pi_{j}}(\bm{\alpha})\right|\right].$
Because $Y^{N}_{0,t}$ is random, by inverting this test, we can construct a confidence interval for $\tau_{0,t}$, which is equivalent to constructing a conformal prediction set (Vovk et al., 2005) for $Y^{N}_{0,T}$. The $(1-\xi)$ confidence interval is given as
$\widehat{C}^{\mathrm{conf}}_{1-\xi}=\left\{\bm{\alpha}\in\mathcal{A}\mid\widehat{p}(\bm{\alpha})>\xi\right\},$
where $\mathcal{A}$ is a set of candidates of $\bm{a}$. Theorem 1 of Chernozhukov et al. (2021) guarantees (7).
7.2 Two-Sample Homogeneity Test
We define the null and alternative hypotheses as $H_{0}:p^{I}_{t}=p^{N}_{t}$ and $H_{1}:p^{I}_{t}\neq p^{N}_{t}$. Then, we can conduct hypothesis testing by applying two-sample homogeneity test, and our distributional treatment effect estimator. For example, we can employ maximum mean discrepancy (MMD) for testing (Gretton et al., 2012).
8 Discussion
This section discusses some remaining problems.
8.1 Justification of Mixture Models
To elucidates assumptions in SCMs, Shi et al. (2022) posits fine-grained models, where each unit has its components, and the outcomes of treated and untreated units are averages of the outcomes of the components. For example, in the empirical study of the tobacco control program in California, the outcome of interest is annual per capita cigarette consumption at the state level, which is an average of the consumption of people in the state. Based on their arguments, we can justify the assumption of mixture models by connecting them to linear factor models.
Assume that there are $U_{j}$ components in unit $j\in\mathcal{J}$. For each $u\in\mathcal{U}_{j}\coloneqq\{1,2,\dots,U_{j}\}$, let $(Y^{I}_{j,t,u},Y^{N}_{j,t,u})$ an outcome of a component $u$ in unit $j$ at time $t$. We also suppose an existence of a $d_{W}$-dimensional unobserved random variable $W_{j,t,u}\in\mathbb{Z}^{d_{W}}$, contributing to individual $u$’s outcome, such as age, education level, or income. We refer to $W_{j,t,u}$ as causes. Assume that $(Y^{I}_{j,t,u},Y^{N}_{j,t,u},W_{j,t,u})$ is an i.i.d. random variable across $u\in\mathcal{U}_{j}$.
The relationship between the causes $W_{j,t,u}$ and potential outcomes $(Y^{I}_{j,t,u},Y^{U}_{j,t})$ can be linear or non-linear, and can also be time-variant. Let $p_{j,t}(w,y)$ be the PDF of $(W_{j,t},Y^{N}_{j,t})$, $p_{j,t}(y\mid w)$ be the PDF of $Y^{N}_{j,t}$ given $W_{j,t}$, and $p_{j,t}$ be the marginalized PDF of $W_{j,t}$. The authors assume that interventions are made at a group-level and individuals in each group comply with their group-level intervention.
Under this setting, Shi et al. (2022) finds that the following assumptions are required for the linearity of SCMs (2).
Assumption 8.1 (Independent Causal Mechanism; ICM).
Conditional on the causes $W_{j,t,u}$, the potential outcome $Y^{N}_{j,t,u}$ is independent of the unit $j\in\mathcal{J}$. For population distribution $j\in\in\mathcal{J}$ at time $t\leq\mathcal{T}$, the joint distribution of $(Y^{N}_{j,t,u},W_{j,t,u})$ is given as $p_{j,t}(w,y)=r_{j,t}(w)p_{t}(y|w)$.
Assumption 8.2 (Stable distributions).
Decompose the causes into two subsets $W_{j,t,u}=\{U_{j,t,u},S_{j,t,u}\}$. Let $S$ denote the subset that differentiates the treated unit from the selected untreated units (minimal invariant set); that is, its distribution in the target unit is different from its distribution in the selected untreated units. We assume that for all groups, the distribution of $S$ does not change for all time periods $t\leq T$, $r_{j,t}(w)=\zeta_{j}(s)\xi_{t}(u\mid s)$, where $\zeta_{j}(s)$ is the PDF of $S_{j,t,u}$, and $\xi_{t}(u\mid s)$ is the PDF of $U_{j,t,u}$ given $S_{j,t,u}$.
Assumption 8.3 (Sufficiently similar untreated units).
Let $S$ be the minimal invariant set for the target and untreated units. The untreated units are sufficiently similar if the cardinality of the minima invariant set, $\big{|}\mathcal{J}^{U}\big{|}\geq|S|$.
Assumption 8.4 (Common support).
Let $(s_{1},\dots,s_{R})$ be the support of $S$. For each $s$, there exists at least one untreated unit $j$, where $\zeta_{j}(S=s)>0$.
Proposition 8.5 (Causal identifiability. From Theorem 1 of Shi et al. (2022)).
Suppose that Assumptions 8.18.4 hold. Then, there exists $\{w^{*}_{j}\}_{j\in[J]}$ such that $p^{N}_{0,t}(y)=\sum_{j\in\mathcal{U}^{U}}w^{*}_{j}p^{N}_{j,t}(y)$ holds.
This proposition also implies that the mixture model holds (Nazaret et al., 2023), which is consistent to the linear factor models. Therefore, based on this result, we can obtain the following result for linear factor models as well as Ferman and Pinto (2021).
Corollary 8.6.
Suppose that Assumptions 5.1, 6.1, and 8.18.4 hold. Then, $\bm{w}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}$ holds as $T_{0}\to\infty$, and $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$. Additionally, $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ converge to $\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)$ in probability as $T_{0}\to\infty$ and $G\to\infty$.
The proof is straightforward. From Proposition 8.5, under Assumptions 8.18.4, (3) holds. Therefore, from Theorem 6.2, the claim holds under (1) with $c_{j}=0$ for $j\in\mathcal{J}$.
8.2 Auxiliary Covariates
We can consider a case where auxiliary covariates are observable. For each unit $j\in\mathcal{J}$, let
$X_{j,t}\coloneqq(X_{j,t,1},X_{j,t,1},\dots,X_{j,t,K})\in\mathbb{R}^{K}$
be $K$-dimensional covariates, which is not affected by treatment for all $j\in\mathcal{J}$. Let
^$\widehat{\ell}_{k}(\bm{w})=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left\{X_{0,t,k}-\sum_{j\in\mathcal{J}^{U}}w_{j}X_{j,t,k}\right\}.$
When the auxiliary covariates are observable, we extend the DMSCM estimator as
$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0}}\coloneqq v_{0}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right)\right|+\sum^{K}_{k=1}v_{k}\widehat{\ell}^{2}_{k}(\bm{w}),$
where $\bm{v}\coloneqq(v_{k})_{k\in\{0,1,2,\dots,K\}}$ are some weights such that $\sum^{K}_{k=0}v_{0}=1$.
8.3 Distributional SCMs and Optimal Transports
Gunsilius (2023) establies the framework of distributional SCMs. In the framework, that study assumes the following:
$F^{-1}_{Y_{0,t},N}(q)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F^{-1}_{Y_{j,t},N}(q),\ \ \ \forall q\in(0,1),$
where $F^{-1}_{Y_{j,t},N}(q)$ is the quantile function defined as $F^{-1}_{Y_{j,t},N}(q)\coloneqq\inf_{y\in\mathbb{R}}\big{\{}F_{Y_{j,t},N}(q)\geq q\big{\}}$. Then, Gunsilius (2023) estimates the SC weight to match the distribution under the Wasserstein distance. Their method is called DiSCo. Specifically, they estimate the SC weights as
$\widehat{\bm{w}}^{\mathrm{DiSCo}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\frac{1}{M}\sum^{M}_{m=1}\left|\widehat{F}^{-1}_{Y^{N}_{0,t}}(V_{m})-\sum_{j\in\mathcal{J}^{U}}w_{j}\widehat{F}^{-1}_{Y^{N}_{j,t}}(V_{m})\right|,$
where $\widehat{F}^{-1}_{Y^{N}_{j,t}}(V_{m})$ is an empirical quantile function of $Y^{N}_{j,t}$, and $\{V_{m}\}$ is a set of $M$ independent draws from the uniform distribution with support $[0,1]$.
8.4 Demeaned Density Matching SCM (D2MSCM)
Next, we consider a case where the target distribution does not follow mixture models, but the deviation is zero when it is marginalized. Instead of (3), We consider the following another mixture model
$p^{N}_{0,t}(y-c_{0})=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p^{N}_{j,t}(y-c_{j})\qquad\forall t\in\mathcal{T}.$
From the viewpoints of linear factor models 1 and fine-grained models (Shi et al., 2022), this assumption allows us $c_{j}\neq 0$ for some $j\in\mathcal{J}$. Under the previous mixture model, we only consider a case where $c_{j}=0$ for all $j\in\mathcal{J}$, which is restrictive in practice.
To estimate the SC weight, we first estimate $c_{j}$ by taking the average of $(Y_{j,t})_{t\in\mathcal{T}_{0}}$ and subtract it from $Y_{j,t}$. This manipulation corresponds to the demeaned estimation in Ferman and Pinto (2021). Let $h\in(0,1)$ be a parameter determined and fixed by us. Then, we estimate SC weights as
$\widehat{\bm{w}}^{\mathrm{D2MSCM}}_{T_{0}}\coloneqq\left(\widehat{w}^{\mathrm{D2MSCM}}_{j,T_{0}}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{0,t}-\overline{Y}^{N}_{0}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}-\overline{Y}^{N}_{0}\Big{)}^{\gamma}\right)\right|.$
We refer to this method as the demeaned DMSCM (D2SCM). By using the D2MSCM estimator $(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0}})_{j\in\mathcal{J}^{U}}$, we estimate the treatment effect as
$\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\left(\overline{Y}^{N}_{0}+\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{D2SCM}}_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)\right).$ (8)
9 Experiments
We conduct experiments with simulated and real-world data. Additional results are in Appendix 9.1.
9.1 Simulation Studies
In this section, we perform simulation studies to show how our proposed SCMs compared to traditional SCMs. Let $J\in\{10,30,60\}$ and $G\in\{2,3,5,10,20,30,40,50,60,70,80,90,100\}$, $T_{0}=30$, $T_{1}=100$, and $K=5$. We perform a simulation on SCMs for $p_{0,t}(\bm{z})=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p_{0,t}(\bm{z})$, where $\bm{z}$ is ($K+1$)-dimensional ($6$-dimensional) vector, where the first element corresponds to $Y_{j,t}$ and the rest corresponds to $X_{j,t}$. For each $j\in\mathcal{J}$, let $p_{j,t}$ be a density of a multinomial normal distribution $\mathcal{N}(\bm{\mu}_{j,t},\bm{\sigma}^{2}_{j,t})$, where $\bm{\mu}_{j,t}=(\mu_{j,t,1},\mu_{j,t,2},\mu_{j,t,3},\mu_{j,t,4},\mu_{j,t,5},\mu_{j,t,6})$, and $\bm{\sigma}^{2}_{j,t}$ is a $(6\times 6)$-diagonal matrix with a $(k,k)$-element denoted by $\sigma^{2}_{j,t,k}$. Each mean parameter is drawn from a standard normal; distribution. Each variance parameter is drawn from a uniform distribution with a support $[1,20]$. Let us assume that for each $j\in\mathcal{J}$, $\mu_{j,t,1}=\mu_{j,t,1}+\epsilon_{j,t}$ and $\sigma^{2}_{j,t,1}=\sigma^{2}_{j,t,1}+\widetilde{\epsilon}_{j,t}$, where $\epsilon_{j,t}$ and $\widetilde{\epsilon}_{j,t}$ follow a normal densities $\mathcal{N}(0,10)$, respectively. Let us correct the variance as $\widetilde{\epsilon}_{j,t}=0.1$ if $\widetilde{\epsilon}_{j,t}\leq 0.1$ occurs. For the treated unit, we assume that $Y^{N}_{0,t}\sim p_{0,t}(\bm{z})$ for $t\in[t_{0},T_{1}]$ and $Y^{I}_{0,t}=20+Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$; that is $\tau_{0,t}=20$. We draw $\bm{w}^{*}$ from a uniform distribution and normalize the sum to be $1$. We apply DMSCM and Abadie $1,000$ times and compute the error of the treatment effect estimation. We show the result with boxplot in Figure 4. Our proposed method successfully estimate the treatment effect with smaller estimation error. The estimation decreases as $G$ increases.
We conduct simulation studies to illustrate the comparative performance of our proposed SCMs and traditional SCMs within a setting analogous to Section 9.1. The only modifications introduced involve the parameters. Specifically, $J\in\{1,5,10,15,20,\dots,35,40,45,50\}$, $G\in\{2,3,5,10\}$, $T_{0}=30$, $T_{1}=1000$, and $K=5$. All other conditions align with those detailed in Section 9.1.
We apply DMSCM and Abadie $1,000$ times and compute the error of the treatment effect estimation in Figure 7. Our focus lies on the fluctuation in mean squared estimation errors as the number of units, $J$, increases. Despite the method by Abadie (2002) incurring escalating estimation errors with the increase of $J$, estimation errors do not exhibit a similar growth under our proposed methods. Furthermore, our proposed methods exhibit lower estimation errors compared to those under the method of Abadie (2002).
A comparable phenomenon is evident in distributional treatment effect estimation. We compare the true and estimated distributions using the Maximum Mean Discrepancy (MMD, Gretton et al., 2012), a probability distance computed through kernel embedding. The upper figure in Figure 8 displays how MMDs alter as $J$ increases. As illustrated, our proposed methods outperform even in this task. The lower figure in Figure 8 presents results for $G\in\{2,10,50,100\}$. As the results indicate, the predictive capability augments as $G$ (the number of moment conditions) increases.
Our findings might have a relationship with those of Spiess et al. (2023), who investigates high-dimensional SCMs. Future investigations will be devoted to exploring these connections.
Figure 2: Results of simulation studies.
Figure 3: Counterfactual outcomes in empirical analysis.
Figure 4: Counterfactual distributions in empirical anaylsis.
Table 2: Results of conformal inference. Estimated treatment effect $\widehat{\tau}_{0,t}$ and $90\%$ confidence interval $(\underline{\tau}_{0,t},\overline{\tau}_{0,t})$.
Terrorist conflict in Basque country Tobacco control program in California Reunification of Germany
Abadie DMSCM Abadie DMSCM Abadie DMSCM
$\gamma$ $2$ $5$ $10$ $100$ $2$ $5$ $10$ $100$ $2$ $5$ $10$ $100$
$\widehat{\tau}_{0,t}$0.800 -0.845 -1.054 -0.279 -0.138 -26.148 -27.744 -32.306 -35.080 -36.879 -1492.374 -1253.616 -1419.120 -1324.448 -510.765
$\overline{\tau}_{0,t}$ 0.522 0.699 0.495 0.521 0.578 8.698 7.852 8.898 9.544 11.748 2635.583 -1008.822 -591.420 -716.376 2327.944
$\underline{\tau}_{0,t}$2.587 -2.150 -2.605 -2.616 -2.590 -36.798 -29.268 -34.783 -37.308 -35.243 -2390.413 -2738.232 -1989.323 -2018.878 -2625.129
Figure 5: Results of simulation studies on ATE prediction. The $x$-axis is the number of untreated units $J$, and the $y$-axis is the mean squared prediction error.
Figure 6: Results of simulation studies on distributional treatment effect prediction. The $x$-axis is the number of untreated units $J$, and the $y$-axis is the MMD.
Figure 7: Results of simulation studies on ATE prediction. The $x$-axis is the number of untreated units $J$, and the $y$-axis is the mean squared prediction error.
Figure 8: Results of simulation studies on distributional treatment effect prediction. The $x$-axis is the number of untreated units $J$, and the $y$-axis is the MMD.
9.2 Empirical Analysis
We conduct empirical studies discussed in Abadie and Gardeazabal (2003), Abadie et al. (2010), and Abadie et al. (2015).
Terrorist conflict in the Basque Country. Abadie and Gardeazabal (2003) investigates the effect of terrorist conflict in the Basque Country on gross domestic product (GDP). In this example, $J=16$, $K=14$, and $(t_{0},T_{0},T_{1})=(1955,1970,1997)$. Basque Euskadi Ta Askatasuna (Basque Homeland and Liberty; ETA) was founded in 1959 as a Basque separatist organization in Spain. ETA promoted an independent Basque state by carrying out terrorist activities mainly in the Basque Country region of Spain. Abadie and Gardeazabal (2003) utilizes annual Spanish regional-level panel data from 1955 to 1997, with a particular focus on the 1970s when ETA conducted significant terrorist activity. In the data, 15 years of pre-intervention data are included. The set of untreated units comprises $17$ other Spanish regions. Each region’s datum contains the outcome of interest (GDP per capita) at the regional level, with covariates including investment rate (percentage), population density, five sectoral productions, and four human capitals, all averaged over the period from 1955 to 1968.
Tobacco control program in California. Abadie et al. (2010) study the effect of a large tobacco control program adopted in California in $1988$, which we explained in our introduction. In this example, $J=20$, $K=5$, and $(t_{0},T_{0},T_{1})=(1970,1989,2000)$. Abadie et al. (2015) investigates the treatment effect of the reunification of Germany. East Germany was absorbed into West Germany, the result being a reunited Germany in 1990. They utilize annual panel data at the country level from 1960 to 2003. The outcome of interest is GDP per capita at the country level. The pre-intervention period spans $30$ years. The set of untreated units comprises $16$ OECD countries that make up synthetic West Germany. Covariates include inflation rate, industry share, investment rate, schooling, and a measure of trade openness.
Reunification of Germany. Abadie et al. (2015) studies the effect of the reunification of Germany on GDP in 1990. Because the countries were dissimilar from any one country to make a comparison group, they apply SCMs to create a composite comparison group. In this example, $J=8$, $K=31$, and $(t_{0},T_{0},T_{1})=(1960,1990,2003)$. Abadie et al. (2010) investigate the causal effect of Proposition 99, a comprehensive tobacco control program adopted by California. They utilize annual state-level panel data from 1970 to 2000. The program was implemented in January 1989, providing 18 years of pre-intervention data. Four states that have adopted large-scale tobacco control programs (Massachusetts, Arizona, Oregon, and Florida) and states that increased state cigarette taxes by $50$ cents or more between 1989 and 2000 (Alaska, Hawaii, Maryland, Michigan, New Jersey, New York, and Washington) and the District of Columbia are excluded from the untreated units of 38 states. The outcome of interest is annual per capita cigarette consumption at the state level, measured as per capita cigarette sales in packs. Covariates include the average retail price of cigarettes, state income per capita (log), percentage of the population aged 15-24, and beer consumption per capita; all averaged over the period from 1980 to 1988.
Results. We show the result of the estimated treatment effect and $90$% confidence interval in Table 2. We show the results of treatment effect estimation in Figure 4 and distributional treatment effect estimation in Figure 4.
For each dataset, we plot the counterfactual outcomes and distributions in Figures 4 and 4, respectively. We also conduct conformal inference on the sharp null hypothesis $H_{0}:\tau_{0,(T_{0}+1)}=\cdots=\tau_{0,T_{1}}=0$ for the treatment effect. For the results of Abadie, DMSCM, We show the $90\%$ confidence interval of the treatment effect in Table 2. Our proposed method has a good fit in terms of the observed outcomes in pretreatment periods $t\in\mathcal{T}_{0}$. Therefore, we consider that our proposed method also predicts counterfactual outcomes and distributions well. In conformal inference, our method does not reject the null hypothesis as with Abadie, except for the DMSCM in the tobacco control problem in California.
10 Conclusion
In this study, we identified the problem of implicit endogeneity in the linear regression models commonly used in existing SCM studies. To address this issue, we proposed a novel approach to SCMs using density matching. Our proposed method enables consistent estimation of SC weights and improves the accuracy of counterfactual outcome predictions compared to existing methods. We validated the effectiveness of our approach through simulation experiments and empirical analyses.
Appendix A Proof of Theorem 6.2
First, we confirm the basic fact about the uniqueness of the CHF.
Proposition A.1 (Theorem 26.2 in Billingsley (1995)).
Distinct measures cannot have the same characteristic function.
Let us define
$\widehat{Q}_{G}(\bm{w})\coloneqq\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right)\right|,$
$Q_{G}(\bm{w})\coloneqq\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|.$
Proof.
First, from the definition of $\Delta^{J}$, $\Delta^{J}$ is a compact parameter space. Next, $\widehat{Q}(\bm{w})$ converges uniformly to $Q(\bm{w})$ as $T_{0}\to\infty$ under Assumption 6.1. Furthermore, $Q(\bm{w})$ is continuous with respect to $\bm{w}$.
Under these conditions, for any $\varepsilon>0$, we have with probability approaching one, we have
  1. 1.
    $\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>\widehat{Q}(\bm{w}_{0})-\varepsilon/3$;
  2. 2.
    $\widehat{Q}_{G}(\widehat{\bm{w}}^{\mathrm{DMSCM}})>\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)-\varepsilon/3$;
  3. 3.
    $\widehat{Q}_{G}\left(\bm{w}_{0}\right)>Q_{G}(\bm{w}_{0})-\varepsilon/3$.
Therefore, with probability approaching one,
$Q_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)-\varepsilon/3>\widehat{Q}_{G}\left(\bm{w}_{0}\right)-2\varepsilon/3>Q_{G}\left(\bm{w}_{0}\right)-\varepsilon.$
Thus, for any $\varepsilon>0$, $Q_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>Q_{G}\left(\bm{w}_{0}\right)-\varepsilon$ with probability approaching one. Let $\mathcal{N}$ be any open subset of $\Delta^{J}$ containing $\widetilde{\Phi}^{*}$. Since $\Theta\cap\mathcal{N}^{c}$ is compact, and $Q_{G}(\bm{w})$ is continuous with respect to $\bm{w}$, we have
$\sup_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w})=Q_{G}(\bm{w}^{\dagger})<Q_{G}(\bm{w}_{0})$
for some $\bm{w}^{\dagger}\in\Theta\cap\mathcal{N}^{c}$. Thus, choosing $\varepsilon=Q_{G}(\bm{w}_{0})-\sup_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w})$, it follows that
$Q_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>\sup_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w}),$
and hence $\widehat{\bm{w}}^{\mathrm{DMSCM}}\in\mathcal{N}$, where recall that $\mathcal{N}$ is any open subset of $\Delta^{J}$ containing $\widetilde{\Phi}^{*}$. Thus, $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}$ holds as $T_{0}\to\infty$. Additionally, from Proposition A.1, we have $\bm{w}^{*}_{G}\to\bm{w}^{*}\in\widetilde{\Phi}^{*}$ as $G\to\infty$. ∎

### v4

@@
-Asymptotically Unbiased Synthetic Control Methods by Distribution Matching
+Asymptotically Unbiased Synthetic Control Methods by Density Matching
 Abstract
-Synthetic Control Methods (SCMs) have become an essential tool for comparative case studies.
-The fundamental idea of SCMs is to estimate the counterfactual outcomes of a treated unit using a weighted sum of the observed outcomes of untreated units.
-The accuracy of the synthetic control (SC) is critical for evaluating the treatment effect of a policy intervention; therefore, the estimation of SC weights has been the focus of extensive research.
-In this study, we first point out that existing SCMs suffer from an endogeneity problem, the correlation between the outcomes of untreated units and the error term of the synthetic control, which yields a bias in the treatment effect estimator.
-We then propose a novel SCM based on density matching, assuming that the density of outcomes of the treated unit can be approximated by a weighted average of the joint density of untreated units (i.e., a mixture model).
-Based on this assumption, we estimate SC weights by matching the moments of treated outcomes with the weighted sum of moments of untreated outcomes.
-Our proposed method has three advantages over existing methods: first, our estimator is asymptotically unbiased under the assumption of the mixture model; second, due to the asymptotic unbiasedness, we can reduce the mean squared error in counterfactual predictions; third, our method generates full densities of the treatment effect, not merely expected values, which broadens the applicability of SCMs.
-We provide experimental results to demonstrate the effectiveness of our proposed method.
+Synthetic Control Methods (SCMs) have become a fundamental tool for comparative case studies.
+The core idea behind SCMs is to estimate treatment effects by predicting counterfactual outcomes for a treated unit using a weighted combination of observed outcomes from untreated units.
+The accuracy of these predictions is crucial for evaluating the treatment effect of a policy intervention.
+Subsequent research has therefore focused on estimating SC weights.
+In this study, we highlight a key endogeneity issue in existing SCMs—namely, the correlation between the outcomes of untreated units and the error term of the synthetic control, which leads to bias in both counterfactual outcome prediction and treatment effect estimation.
+To address this issue, we propose a novel SCM based on density matching, assuming that the outcome density of the treated unit can be approximated by a weighted mixture of the joint density of untreated units.
+Under this assumption, we estimate SC weights by matching the moments of the treated outcomes with the weighted sum of the moments of the untreated outcomes.
+Our method offers three advantages: first, under the mixture model assumption, our estimator is asymptotically unbiased; second, this asymptotic unbiasedness reduces the mean squared error in counterfactual predictions; and third, our method provides full densities of the treatment effect rather than just expected values, thereby broadening the applicability of SCMs.
+Finally, we present experimental results that demonstrate the effectiveness of our approach.
 [^1] This study was presented at the Workshop on Counterfactuals in Minds and Machines at the International Conference on Machine Learning in July 2023 and at the International Conference on Econometrics and Statistics in August 2023.
 1 Introduction
-Synthetic control methods (SCMs, Abadie and Gardeazabal, 2003; Abadie et al., 2010) have gained attention as a crucial tool for program evaluation in comparative case studies across fields such as economics, statistics, and machine learning.
-According to Athey and Imbens (2017), SCMs are “arguably the most important innovation in the policy evaluation literature in the last 15 years” due to their “simplicity of the idea, and the obvious improvement over standard methods.”
-SCMs posit a scenario in which multiple units exist, with one unit subjected to a policy intervention (treated unit) at a certain time period, while the rest remain untreated.
-The focus is on estimating the treatment effect of the treated unit, i.e., the difference between its observed factual outcome and unobserved counterfactual outcome.
-Untreated units are used to account for unobserved trends in the outcome over time that are unrelated to the effect of the policy intervention.
-The key insight is that an optimally weighted average of the untreated units, referred to as the synthetic control (SC) unit, provides a more appropriate estimate of the counterfactual than estimating the counterfactual using only the treated unit.
+Synthetic control methods (SCMs, Abadie and Gardeazabal, 2003; Abadie et al., 2010) have emerged as a crucial tool for program evaluation in comparative case studies across disciplines such as economics, statistics, and machine learning.
+As noted by Athey and Imbens (2017), SCMs represent “arguably the most important innovation in the policy evaluation literature in the last 15 years” due to their “simplicity of the idea, and the obvious improvement over standard methods.”
+In SCMs, there are multiple units, where one unit receives a policy intervention (treated unit) at a certain time while the remaining units remain untreated.
+The objective is to estimate the treatment effect for the treated unit, defined as the difference between its observed factual outcome and its unobserved counterfactual outcome.
+Untreated units help account for unobserved trends in the outcome over time that are unrelated to the policy intervention.
+The key idea is that an optimally weighted combination of the untreated units, referred to as the synthetic control (SC) unit, serves as an appropriate estimate of the counterfactual outcome.
 See Figure 1 for an illustration.
-SCMs have been applied in a variety of settings, such as studying the effect of terrorism on GDP (Abadie and Gardeazabal, 2003) and the decriminalization of indoor prostitution (Cunningham and Shah, 2017).
+d SCMs have been widely applied in various empirical studies, such as evaluating the impact of terrorism on GDP (Abadie and Gardeazabal, 2003) and the decriminalization of indoor prostitution (Cunningham and Shah, 2017).
+For instance, Abadie et al.
+(2010) applies SCM to study the treatment effect of a comprehensive tobacco control program implemented in California in 1988.
+While smoking rates in California declined following the program’s implementation, it remained unclear whether this decrease was attributable to the policy.
+For estimating the treatment effect, annual per-capita cigarette sales data from states that did not implement a similar tobacco control program were used as untreated units.
+SCM was then applied to construct California’s counterfactual outcome, representing the smoking rate in the absence of the tobacco control program.
+A key challenge in SCMs is the assumption, made in most existing studies (Abadie, 2002), that the expected outcome of the treated unit can be represented as a weighted sum of the outcomes of untreated units.
+Ferman and Pinto (2021) highlights that this assumption leads to a non-negligible bias in the estimator.
+According to Ferman and Pinto (2021), this bias arises from endogeneity in the linear regression model implied by the SCM assumption due to the correlation between the outcomes of untreated units and the error term.
+This issue is analogous to the endogeneity caused by measurement error (Greene, 2003).
+We refer to this problem as implicit endogeneity.
+To address this issue, we propose an alternative approach based on a mixture model assumption rather than a linear relationship among expected outcomes.
+Specifically, we assume that the density of the treated unit’s outcome can be approximated by a weighted combination of the densities of the untreated units.
+This assumption further implies that each moment of the treated unit’s outcome can be represented as a weighted sum of the moments of the untreated units.
+Leveraging this insight, we estimate SC weights by matching the moments of the treated unit’s outcome to those of the untreated units.
+We refer to our method as the Density Matching SCM (DMSCM).
+Unlike existing methods, our proposed estimator is asymptotically unbiased if the mixture model assumption is correct.
+We validate the effectiveness of our approach through both simulation studies and empirical analysis.
+The mixture model assumption has appeared in previous and concurrent studies, including Wan et al.
+(2018), Gunsilius (2023), and Shi et al.
+(2022).
+Our proposed method can be viewed as a simplification of the distributional SCMs introduced by Gunsilius (2023).
+While Gunsilius (2023) does not explicitly address the implicit endogeneity problem, they note that their results diverge from those obtained under standard SCMs.
+This discrepancy arises because, under the mixture model assumption, Gunsilius (2023) provides a consistent estimator for the treatment effect, whereas existing SCM methods do not.
+We establish a connection between distributional SCMs and the asymptotic bias problem and propose a simplified estimation method based on density matching.
+In summary, our contributions are as follows:
+  • Reexamining the assumptions necessary for treatment effect estimation in SCMs.
+  • Proposing a density-matching approach for asymptotically unbiased SCMs.
+  • Simplifying the implementation of density matching in SCMs.
+We formulate the problem in Section 2.
+Next, in Section 3, we review the basic idea of SCMs.
+In Section 4, we highlight the problem of implicit endogeneity.
+In Section 5, we then propose our method under the mixture model assumption.
+We investigate the asymptotic properties of this method in Section 6 and present a statistical inference method in Section 7.
+In Section 8, we discuss related topics and existing studies.
+Finally, in Section 9, we examine the performance of our estimators through extensive simulations.
 Figure 1: Illustration of SCMs
-For instance, Abadie et al.
-(2010) used SCM to study the treatment effect of a comprehensive tobacco control program implemented in California in 1988.
-Smoking rates in California decreased following the program’s implementation, but it was unclear if this decrease was a result of the program or a long-term trend.
-To investigate the treatment effect, annual per-capita cigarette sales data across several states that did not implement a similar tobacco control program were used as untreated units.
-They then applied SCM to estimate California’s counterfactual outcome: the smoking rate in California in 1989 had the tobacco control program not been implemented.
-For further details, also see Abadie (2021).
-A significant challenge in SCMs is the typical assumption that the expected outcome of the treated unit can be represented by the weighted sum of untreated units.
-However, Ferman and Pinto (2021) points out that there exists a non-negligible asymptotic bias in the obtained estimator.
-According to Ferman and Pinto (2021), this asymptotic bias is caused by a kind of endogeneity in the linear regression model deduced from the assumption, due to a correlation between the outcomes of untreated units and the error term of the linear model.
-This problem is akin to endogeneity caused by measurement error (Greene, 2003).
-We refer to this endogeneity as implicit endogeneity.
-To address the aforementioned issue, we propose to assume a mixture model instead of a linear relationship among expected outcomes.
-Specifically, we assume that the density of the outcome of the treated unit can be approximated by a linear combination of those of the untreated units.
-This mixture model assumption also implies that each moment of the outcome of the treated unit can be represented by a linear combination of those of the untreated units.
-Therefore, we can estimate the SC weights by matching all moments between the outcome of the treated unit and those of the untreated units.
-We refer to our method as the Density Matching SCM (DMSCM).
-Unlike existing methods, our proposed estimator is asymptotically unbiased.
-We validate the efficacy of our proposed method through simulation studies and empirical analysis.
-Our proposed method corresponds to a simplification of the distributional SCMs by Gunsilius (2023).
-Although Gunsilius (2023) does not explicitly point out the implicit endogeneity problem, it observes that their result does not coincide with the result under the standard SCMs.
-This is because under the mixture model assumption, while Gunsilius (2023) returns a consistent estimator for the treatment effect, existing methods do not.
-We connect the distributional SCMs to the asymptotic bias problem and propose a simplified estimation method for SCMs by distribution matching.
-In summary, our contributions are as follows:
-  • Reconsidering the assumptions necessary for treatment effect estimation in SCMs.
-  • Proposing a density-matching method for asymptotically unbiased SCMs.
-  • Simplifying the density-matching method.
-1.1 Related Work
+Related work.
 Following the pioneering work by Abadie and Gardeazabal (2003), numerous studies have advanced the use of SCMs in comparative studies (Abadie et al., 2010, 2015).
@@
 (2019), are based on minimizing the pre-treatment sum of squared errors between the outcome of the treated unit and the weighted sum of those of the untreated units, with some regularization and constraints regarding the weights.
-Gunsilius (2023) explores distribution SCMs (DiSCo), which estimate SC weights to match the distribution of the outcomes of the treated unit and the weighted sum of the distributions of the untreated units.
-[^2] The method of Gunsilius (2023) is called DiSCo by Dijcke (2024).
-His method is based on minimizing the Wasserstein distance, which can be simplified as matching the quantiles in the setting of SCMs.
-Their assumption about identification can be interpreted as a variant of mixture models.
-From a theoretical viewpoint, Abadie et al.
-(2010) demonstrate that, if there are weights that provide a perfect pre-treatment fit, then the bias of the SC estimator is bounded by a function that converges to zero asymptotically as the length of the pre-treatment periods $T_{0}$ increases, even when the number of untreated units $J$ is fixed.
-However, Ferman and Pinto (2021) finds that if the pre-treatment fit is imperfect, then the conventional SCMs do not yield a consistent estimator of the treatment effect due to the bias caused by the correlation between the outcomes of the untreated units and the error term in the posited linear models.
-We review this problem in Section 4 and name it as the implicit endogeneity.
-To address this bias issue, existing studies have proposed several approaches.
-Ferman (2021) demonstrates that we can obtain an asymptotically unbiased estimator under the regime where the number of untreated units increases ($J\to\infty$) as well as the length of the pre-intervention periods $T_{0}$ also increases.
-The use of mixture models has a close relationship with linear factor models, assumed in existing studies (Abadie, 2021).
-We justify the use of the former by employing arguments from Shi et al.
-(2022) and Nazaret et al.
+We refer to such SCMs as Least-Squares SCMs (LS-SCMs).
+The bias of the LS-SCMs strongly depends on whether there exist weights that can predict the outcome of the treated unit without error.
+In other words, a random outcome of the treated unit must be replicable by the sum of outcomes of untreated units.
+Since such weights yield a perfect fit in the estimation process using observations from the pre-treatment period, this situation is called pretreatment-fit.
+Abadie et al.
+(2010) demonstrates that under a pre-treatment fit, the bias of the LS-SCM estimator converges to zero asymptotically as the length of the pre-treatment periods $T_{0}$ increases.
+However, Ferman and Pinto (2021) finds that if the pre-treatment fit does not hold, then the LS-SCM estimator has a bias and is not consistent.
+The cause of this bias is the correlation between the outcomes of the untreated units and the error term in the posited linear model, which we refer to as implicit endogeneity.
+Several methods have been proposed to address this bias issue.
+Ferman (2021) demonstrates that we can obtain an asymptotically unbiased estimator under the regime where both the number of untreated units ($J\to\infty$) and the length of the pre-intervention periods ($T_{0}$) grow large.
+Concurrently with us, Fry (2024) also discusses the endogeneity problem in SCMs as a cause of the asymptotic bias reported by Ferman and Pinto (2021).
+[^2] Note that his study became public on arXiv after our study became public on arXiv and at the ICML workshop, and he and we conducted our studies completely independently.
+While our formulation and his regarding endogeneity are similar, our approach differs significantly.
+We address the problem through density matching, similar to the strategy in Gunsilius (2023), whereas Fry (2024) employs an instrumental variable approach using outcomes of untreated units not used in SCMs as instruments.
+In this study, we focus on mixture models among outcome distributions, an assumption also employed in Wan et al.
+(2018), Gunsilius (2023), Shi et al.
+(2022), and Nazaret et al.
 (2023).
-Our proposed method can be considered a new interpretation of existing SCMs.
-By reinterpreting existing SCMs through density matching, in addition to addressing the implicit endogeneity problem, we address two problems in the SCM literature.
-First, our proposed method improves the pre-treatment fit by debiasing the estimator.
-Second, our method generates full counterfactual densities of the treated unit, rather than merely the expected value.
-As a result, we can conduct statistical inference on the distribution, or any function of the distribution, such as quantiles (Gunsilius, 2023), Lorenz curves (Gastwirth, 1971), and expected social welfare (Abadie, 2002).
-An excellent illustration of the relevance of estimating distribution functions is Ropponen (2011), who applies the changes-in-changes estimator by Athey and Imbens (2006) to estimate the effect of minimum wage changes on employment levels in order to reconcile the classical conflicting results about the minimum wage debate in Card and Krueger (2000) and Neumark and Wascher (2000).
-Spiess et al.
-(2023) investigates high-dimensional linear regression models in SCMs.
-The exploration of identification assumptions has been attempted in the existing literature.
-Ferman and Pinto (2021) conducts a comprehensive investigation to discern the conditions under which SCMs can accurately predict counterfactual outcomes.
-This study also answers an important open question raised by Gunsilius (2023).
-Gunsilius (2023) discusses that an optimal weight obtained by the density matching differs from that obtained by the classical least squares approach.
-Gunsilius (2023) conjectures that the classical least squares approach estimates a weight to match the first moment, while the density matching approach estimates a weight to match all moments.
-However, we point out that the classical least squares approach estimates a weight not only to match the first moment but also tends to incur bias due to the implicit endogeneity.
-Concerning the assumptions about SCMs, Wan et al. (2018) points out that the SCMs essentially assume mixture models, which distinguish SCMs from traditional panel data analysis.
-On another note, Shi et al.
-(2022) introduces fine-grained models to investigate hidden assumptions inherent to SCMs, which also explores a connection between the linear combination of expected outcomes and the linear factor models.
+Although these studies introduce mixture models with motivations other than the bias problem, we highlight that such models can effectively circumvent bias under imperfect pre-treatment fit.
+Here, we briefly review existing work using mixture models.
+Wan et al.
+(2018) points out that SCMs essentially rely on mixture models, distinguishing them from traditional panel data analysis.
+Gunsilius (2023) proposes distribution SCMs (DiSCo), which estimate SC weights by matching the outcome distributions of the treated unit and the weighted sum of the distributions of the untreated units.
+Shi et al.
+(2022) introduces fine-grained models to elucidate hidden assumptions in SCMs and explores connections to linear factor models.
 Nazaret et al.
-(2023) extends these findings into the realm of model misspecification.
-Their formulation is also based on the mixture models.
-Independently of us, Fry (2024) also discusses the endogeneity problem in SCMs as a cause of the asymptotic bias pointed out by Ferman and Pinto (2021).
-[^3] Note that his study became public on arXiv after our study became public on arXiv and at the ICML workshop, and we and he conducted our studies completely independently.
-While our and his formulations about the endogeneity are similar, the approaches are significantly different.
-We address the problem from the viewpoint of density matching, which is an approach similar to Gunsilius (2023).
-In contrast, Fry (2024) uses an instrumental variable method by employing outcomes of untreated units that are not used in SCMs as instrumental variables.
+(2023) extends these insights by analyzing model misspecification under a mixture model formulation.
+This study also addresses a question raised by Gunsilius (2023).
+While Gunsilius (2023) suggests that standard least squares SCMs try to match the first moment of the outcome distribution (and therefore differ from distributional approaches), we stress that least squares approaches are additionally vulnerable to implicit endogeneity, which can generate bias.
+Our density-matching approach avoids this problem by matching a broader set of distributional features.
 Table 1: Comparison of SCMs. See the following sections for the definitions of variables and methods.
@@
 Fry (2024) IV method (GMM) $\mathbb{E}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}[Y^{N}_{j,t}]$ Unbiased Use IVs
-1.2 Organizations
-In Section 2, we formulate the problem.
-In Section 3, we review an idea of SCMs.
-In Section 4, we point out the problem of implicit endogeneity.
-Then, in Section 5, we propose our method under the assumption of mixture models.
-In Section 6, we investiate the asymptotic property of our proposed method.
-In Section 7, we present a statistical inference method for our proposed method.
-In Section 8, we discuss related topics and existing studies.
-We numerically investigate the performance of the proposed estimators through extensive simulations in Section 9.
 2 Problem Setting
-Suppose that there are $J+1$ units, indexed by $j\in\mathcal{J}\coloneqq\{0,1,2,\dots,J\}$, and a time series indexed by $t\in\mathcal{T}=\{1,2,\dots,T\}$ where $3\leq T\in\mathbb{N}$.
-Without loss of generality, we assume that the first unit ($j=0$) is the treated unit, i.e., the unit affected by the policy intervention of interest.
-A set of potential comparisons, $j\in\mathcal{J}^{U}\coloneqq\mathcal{J}\backslash\{0\}=\{1,2,\dots,J\}$, consists of untreated units not affected by the intervention.
-We also assume that our data span $T$ periods and that the intervention occurs at $t=T_{0}$, such that $2\leq T_{0}<T$.
-Observations at period $t\in\mathcal{T}_{0}\coloneqq\{1,2,\dots,T_{0}\}$ are from before the intervention, and observations at period $t\in\mathcal{T}_{1}\coloneqq\mathcal{T}\backslash\mathcal{T}_{0}$ are from after the intervention.
+Suppose there are $J+1$ units, indexed by $j\in\mathcal{J}\coloneqq\{0,1,2,\dots,J\}$, and a time series indexed by $t\in\mathcal{T}=\{1,2,\dots,T\}$, where $3\leq T\in\mathbb{N}$.
+Without loss of generality, assume that the first unit ($j=0$) is the treated unit, i.e., the unit affected by the policy intervention of interest.
+The set of potential control units is $\mathcal{J}^{U}\coloneqq\mathcal{J}\backslash\{0\}=\{1,2,\dots,J\}$, consisting of units not affected by the intervention.
+We also assume that our data span $T$ periods and that the intervention occurs at $t=T_{0}$, with $2\leq T_{0}<T$.
+Observations in period $t\in\mathcal{T}_{0}\coloneqq\{1,2,\dots,T_{0}\}$ are from before the intervention, and observations in period $t\in\mathcal{T}_{1}\coloneqq\mathcal{T}\backslash\mathcal{T}_{0}$ are from after the intervention.
 Let $T_{1}\coloneqq|\mathcal{T}_{1}|=T-T_{0}$.
 2.1 Potential Outcomes
-First, following the Neyman-Rubin causal model (Neyman, 1923; Rubin, 1974), we define the potential outcomes of the policy intervention.
+Following the Neyman-Rubin causal model (Neyman, 1923; Rubin, 1974), we define the potential outcomes of the policy intervention.
 For each unit $j\in\mathcal{J}$ and period $t\in\mathcal{T}$, there are two outcomes of interest, $\left(Y^{I}_{j,t},Y^{N}_{j,t}\right)\in\mathbb{R}^{2}$, corresponding to scenarios with and without the intervention.
@@
 2.2 Observations
-We then define our observations. For each unit $j\in\mathcal{J}$, we can only observe one of the outcomes corresponding to the actual intervention; that is, we observe $Y_{j,t}\in\mathbb{R}$ such that
-$Y_{0,t}=\begin{cases}Y^{N}_{0,t}&\forall t\in\mathcal{T}_{0}\\ Y^{I}_{0,t}&\forall t\in\mathcal{T}_{1}\end{cases},\qquad\mathrm{and}\qquad Y_{j,t}=Y^{N}_{j,t}\qquad\forall j\in\mathcal{J}^{U},\forall t\in\mathcal{T}.$
-We aim to estimate the effect of the policy intervention by predicting $Y^{I}_{0,t}$ from $(Y^{N}_{j,t})_{j\in\mathcal{J}^{U}}$ and by comparing this prediction to the observed $Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$.
-We will define the treatment effect more formally in the following section.
+We next define our observations.
+For each unit $j\in\mathcal{J}$, only one of the two potential outcomes can be observed, corresponding to what actually happened.
+Specifically, for each $j\in\mathcal{J}$, we observe $Y_{j,t}\in\mathbb{R}$ such that
+$Y_{0,t}=\begin{cases}Y^{N}_{0,t}&\forall t\in\mathcal{T}_{0},\\ Y^{I}_{0,t}&\forall t\in\mathcal{T}_{1},\end{cases}\quad\mathrm{and}\quad Y_{j,t}=Y^{N}_{j,t}\quad\forall j\in\mathcal{J}^{U},\forall t\in\mathcal{T}.$
+We aim to estimate the effect of the policy intervention by predicting $Y^{I}_{0,t}$ from $\left(Y^{N}_{j,t}\right)_{j\in\mathcal{J}^{U}}$ and comparing that prediction to the observed $Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$.
+We define the treatment effect more formally in the next subsection.
 2.3 Treatment Effect
-In SCMs, there are several ways to define the treatment effect $\tau_{0,t}\in\mathbb{R}$ for each $t\in\mathcal{T}_{1}$ of the intervention for the treated unit.
-First, for each $t\in\mathcal{T}_{1}$, we define the treatment effect on the treated unit $j=0$ as
-$\tau_{0,t}\coloneqq\mathbb{E}\big{[}Y^{I}_{0,t}\big{]}-\mathbb{E}\big{[}Y^{N}_{0,t}\big{]}.$
+In SCMs, there are several ways to define the treatment effect $\tau_{0,t}\in\mathbb{R}$ for each $t\in\mathcal{T}_{1}$ (i.e., after the intervention).
+First, for each $t\in\mathcal{T}_{1}$, we can define the treatment effect for the treated unit $j=0$ as
+$\tau_{0,t}\coloneqq\mathbb{E}[Y^{I}_{0,t}]-\mathbb{E}[Y^{N}_{0,t}].$
 Another definition is based on the following linear factor model:
@@
 $Y^{N}_{j,t}=c_{j}+\delta_{t}+\lambda_{t}\mu_{j}+\epsilon_{j,t},\quad\mathbb{E}_{j,t}[\epsilon_{j,t}]=0,\quad\forall t\in\mathcal{T},\quad\forall j\in\mathcal{J},$ (1)
-$Y^{I}_{j,t}=\tau_{j,t}+Y^{N}_{j,t}.$
-where $\delta_{t}$ is an unobserved common factor with constant factor loadings across units, $c_{j}$ is an unknown time-invariant fixed effect, $\lambda_{t}$ is a $(1\times F)$ vector of unobserved common factors, $\mu_{j}$ is a $(F\times 1)$ vector of unknown factor loadings, and the error terms $\epsilon_{j,t}$ are unobserved idiosyncratic shocks.
-Here, while $c_{j}$, $\mu_{j}$, and $\tau_{0,t}$ are non-random variables, $\delta_{t}$, $\lambda_{t}$, and $\epsilon_{j,t}$ are random variables.
-Under the linear factor model, the treatment effect is given by $\tau_{0,t}=Y^{I}_{0,t}-Y^{N}_{0,t}$, a non-random variable by definition of the factor model.
-Therefore, the expression $\tau_{0,t}\coloneqq\mathbb{E}\big{[}Y^{I}_{0,t}\big{]}-\mathbb{E}\big{[}Y^{N}_{0,t}\big{]}$ also holds under the factor model.
-Comparative case studies aim to evaluate $\tau_{0,t}$ by predicting $Y^{N}_{0,t}$ for $t>T_{0}$, which is the counterfactual outcome that would have been observed for the treated unit in the absence of the intervention.
-By definition, $Y^{N}_{0,t}$ is a counterfactual outcome because of the intervention after $t=T_{0}$.
-We usually use one unit or a small number of units that have similar characteristics to the treated unit at the time of the intervention but are not exposed to the intervention.
-However, when the data consist of a few units, such as regions or countries, it is often difficult to find a single untreated unit that provides a suitable comparison for the unit treated by the policy intervention of interest.
+$Y^{I}_{j,t}=\tau_{j,t}+Y^{N}_{j,t},$
+where $\delta_{t}$ is an unobserved common factor with constant factor loadings across units, $c_{j}$ is an unknown time-invariant fixed effect, $\lambda_{t}$ is a $(1\times F)$ vector of unobserved common factors, $\mu_{j}$ is a $(F\times 1)$ vector of unknown factor loadings, and $\epsilon_{j,t}$ is an unobserved idiosyncratic shock.
+Here, $c_{j}$, $\mu_{j}$, and $\tau_{0,t}$ are non-random variables, while $\delta_{t}$, $\lambda_{t}$, and $\epsilon_{j,t}$ are random variables.
+Under the linear factor model, the treatment effect is given by $\tau_{0,t}=Y^{I}_{0,t}-Y^{N}_{0,t}$, a non-random variable by definition.
+Consequently, the expression $\tau_{0,t}\coloneqq\mathbb{E}[Y^{I}_{0,t}]-\mathbb{E}[Y^{N}_{0,t}]$ also holds under the factor model.
+Comparative case studies aim to evaluate $\tau_{0,t}$ by predicting $Y^{N}_{0,t}$ for $t>T_{0}$, the counterfactual outcome that would have been observed for the treated unit had the intervention not occurred.
+Because the intervention takes place after $t=T_{0}$, $Y^{N}_{0,t}$ is a counterfactual outcome.
+In settings where data consist of only a few units (e.g., regions or countries), it can be difficult to find a single untreated unit that provides a suitable comparison.
+For this problem, SCMs enable us to construct a counterfactual outcome using a predictor based on the outcomes of the untreated units.
+Under suitable model assumptions that relate $Y^{N}_{0,t}$ to $\left(Y^{N}_{j,t}\right)_{j\in\mathcal{J}^{U}}$, we can construct the predictor using observations from the pre-intervention period and then predict the unobserved counterfactual outcome in the post-intervention period.
+Consequently, we estimate the treatment effect by comparing the predicted $Y^{N}_{0,t}$ with the observed $Y^{I}_{0,t}$.
 Distributional Treatment Effects.
-In program evaluation, we are often interested not only in the scalar treatment effect but also in the distributions of outcomes (Maier, 2011) or their functionals, such as the quantile treatment effects (QTEs) and expected social welfare (Abadie, 2002).
-In this study, we consider estimating $p^{I}_{0,t}(y)$ and $p^{N}_{0,t}(y)$ for $t\in\mathcal{T}_{1}$, where outcomes corresponding to $p^{I}_{0,t}(y)$ are unobservable.
-The treatment effect is then captured by $p^{I}_{0,t}(y)$ and $p^{N}_{0,t}(y)$, or functionals of these distributions, such as QTEs and expected social welfare.
-Depending on the application, we may adjust the target density, for instance, $\overline{p}^{N}_{0}(y)=\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p^{N}_{0,t}(y)$.
+In program evaluation, we are often interested not only in the scalar treatment effect but also in the distributions of outcomes (Maier, 2011) or their functionals, such as quantile treatment effects (QTEs) and expected social welfare (Abadie, 2002).
+By applying our method, we can also estimate the distribution of the counterfactual outcome, $p^{N}_{0,t}(y)$, for $t\in\mathcal{T}_{1}$.
+The treatment effect is then captured by the difference between $p^{I}_{0,t}(y)$ (the distribution of the actual outcome) and $p^{N}_{0,t}(y)$, or by functionals of these distributions, such as QTEs and expected social welfare.
+Depending on the application, we may also adjust the target density, for example, by averaging across all post-intervention periods: $\overline{p}^{N}_{0}(y)=\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p^{N}_{0,t}(y)$.
 We refer to the treatment effects related to these distributions as distributional treatment effects (Park et al., 2021; Kallus and Oprescu, 2022; Chikahara et al., 2022).
-In the context of SCMs, Gunsilius (2023) and Chen (2020) have estimated the treatment effect.
+In the context of SCMs, Gunsilius (2023) and Chen (2020) have proposed methods for estimating such distributional treatment effects.
 3 Recap of SCMs
 This section reviews the original SCM proposed by Abadie and Gardeazabal (2003).
-To estimate the treatment effect $\tau_{0,t}$, we predict the counterfactual outcome $Y^{N}_{0,t}$ by using the weighted sum of $\{Y^{N}_{j,t}\}^{J}_{j=1}$ for $t\in\mathcal{T}_{1}$; that is, SCMs predict $Y^{N}_{0,t}$ as
+To estimate the treatment effect $\tau_{0,t}$, we predict the counterfactual outcome $Y^{N}_{0,t}$ using a weighted sum of $\{Y^{N}_{j,t}\}_{j=1}^{J}$ for $t\in\mathcal{T}_{1}$.
+Specifically, SCMs predict $Y^{N}_{0,t}$ as
 $\widehat{Y}^{N}_{0,t}\coloneqq w_{0}+\sum_{j\in\mathcal{J}^{U}}w_{j}Y^{N}_{j,t},$
-where $w_{0}\in\mathbb{R}$ and $w_{j}\in\mathbb{R}$ $(j\in\mathcal{J}^{U})$ are the weights. We refer to these weights as SC weights.
+where $w_{0}\in\mathbb{R}$ and $w_{j}\in\mathbb{R}$ $(j\in\mathcal{J}^{U})$ are the SC weights.
 Several methods exist for estimating the SC weights.
-In the SCM proposed by Abadie and Gardeazabal (2003), SC weights are determined by minimizing the squared error between $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{1t},\dots,Y^{N}_{J,t}$.
-Let us define
-$\Delta^{J}\coloneqq\{\bm{w}=(w_{1}\ w_{2}\ \dots\ w_{J})^{\top}\in[0,1]^{J}\mid\sum_{j\in\mathcal{J}^{U}}w_{j}=1\}.$
-Then, Abadie and Gardeazabal (2003) estimates the SC weights as
-$\bm{w}^{\mathrm{Abadie}}=\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{0,t}-\sum_{j\in\mathcal{J}^{U}}w_{j}Y^{N}_{j,t}\right)^{2},$
-The counterfactual outcome is then predicted as
+In the SCM proposed by Abadie and Gardeazabal (2003), these weights are determined by minimizing the squared error between $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$.
+Define
+$\Delta^{J}\coloneqq\left\{\bm{w}=(w_{1}\ w_{2}\ \dots\ w_{J})^{\top}\in[0,1]^{J}\mid\sum_{j\in\mathcal{J}^{U}}w_{j}=1\right\}.$
+Then, Abadie and Gardeazabal (2003) estimate the SC weights as
+$\bm{w}^{\mathrm{Abadie}}=\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{0,t}-\sum_{j\in\mathcal{J}^{U}}w_{j}Y^{N}_{j,t}\right)^{2}.$
+Recall that we referred to SCMs using this approach as the Least-Squares-SCMs (LS-SCMs). The counterfactual outcome is then predicted as
 $\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}=\widehat{w}^{\mathrm{Abadie}}_{0}+\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{Abadie}}_{j}Y^{N}_{j,t}.$
-Using the prediction $\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}$, we estimate the treatment effect as $\widehat{\tau}^{\mathrm{Abadie}}_{t}=Y^{I}_{0,t}-\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}$.
-The literature on inference for SCMs has grown significantly, extending beyond the original proposals of Abadie and Gardeazabal (2003) and Abadie et al.
+Using $\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}$, we estimate the treatment effect as
+$\widehat{\tau}^{\mathrm{Abadie}}_{t}=Y^{I}_{0,t}-\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}.$
+The literature on inference for SCMs has grown significantly, extending the original proposals by Abadie and Gardeazabal (2003) and Abadie et al.
 (2010).
@@
 where $\lambda_{1}>0$ and $\lambda_{2}>0$ are regularization parameters.
-The parameter space for $\widetilde{\bm{w}}$ can be restricted in various ways, such as setting $w_{0}=0$ and $\bm{w}=(w_{1}\ \cdots\ w_{J})^{\top}\in\Delta^{J}$.
-4 Identification and Endogeneity
-This section addresses the identification problem of treatment effects from the perspective of endogeneity.
+4 Identification and Implicit Endogeneity
+This section addresses the identification problem of treatment effects and points out the problem of implicit endogeneity in LS-SCMs.
 4.1 Linear Models in Expected Outcomes
-We reconsider the underlying assumptions in SCMs.
-Existing studies often posit some linear relationship between $Y^{N}_{0,t}$ and $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$.
-One such relationship assumes the existence of $\bm{w}^{*}\in\Delta^{J}$ (Abadie et al., 2010):
+We revisit the assumptions underlying SCMs.
+In many existing studies, a linear relationship is assumed between $Y^{N}_{0,t}$ and $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$.
+One possibility is positing the existence of $\bm{w}^{*}\in\Delta^{J}$ (Abadie et al., 2010) such that
 $Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}.$
-This assumption, known as the perfect pre-treatment fit (Ferman and Pinto, 2021), implies that the stochastic process of $Y^{N}_{0,1},\dots,Y^{N}_{0,T}$ can be completely replicated by $\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$.
-In linear factor models, $Y^{N}_{j,t}=c_{j}+\delta_{t}+\lambda_{t}\mu_{j}+\epsilon_{j,t}$, and this assumption might imply $\epsilon_{0,t}=\sum_{j\in[J]}w^{*}_{j}\epsilon_{j,t}$, which is considered unrealistic (note that this is a specific case and not a necessary condition for $Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$).
-A more natural assumption is that each term in the linear factor models of $Y^{N}_{0,t}$ can be replicated by the weighted sum of corresponding terms in the models of $Y^{N}_{0,1},\dots,Y^{N}_{0,T}$.
-A weaker assumption still is a linear relationship between $\mathbb{E}_{0,t}\left[Y^{N}_{0,t}\right]$ and $(\mathbb{E}_{j,t}\left[Y^{N}_{j,t}\right])_{j\in\mathcal{J}^{U}}$, which includes the replication of each term in linear factor models as a specific case.
+This assumption, often called the perfect pre-treatment fit (Ferman and Pinto, 2021), implies that the stochastic process $\left(Y^{N}_{0,t}\right)_{t\in\mathcal{T}_{0}}$ can be fully replicated by $\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$ for $t\in\mathcal{T}_{0}$.
+In linear factor models where $Y^{N}_{j,t}=c_{j}+\delta_{t}+\lambda_{t}\mu_{j}+\epsilon_{j,t}$, this assumption might imply $\epsilon_{0,t}=\sum_{j\in[J]}w^{*}_{j}\epsilon_{j,t}$, which is often considered unrealistic.
+Note that this is a specific case and not necessary for $Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$.
+A more natural assumption is a linear model between $\mathbb{E}_{0,t}[Y^{N}_{0,t}]$ and $(\mathbb{E}_{j,t}[Y^{N}_{j,t}])_{j\in\mathcal{J}^{U}}$.
 Assumption 4.1.
-There exists a set $\widetilde{\Phi}\subset\Delta^{J}$ such that for all $\bm{w}^{*}=(w^{*}_{j})_{j\in\mathcal{J}^{U}}\in\widetilde{\Phi}$, it holds that
-$\mathbb{E}_{0,t}\left[Y^{N}_{0,t}\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[Y^{N}_{j,t}\right],$ (2)
-where $\Delta^{J}\coloneqq\{(w_{1},\dots,w_{J})\in\mathbb{R}^{J}\mid w_{j}\geq 0\ \forall j\ \mathrm{and}\ \sum_{j\in\mathcal{J}^{U}}w_{j}=1\}$.
-Compared to Assumption 4.1, assuming $Y_{0,t}=\sum_{j\in\mathcal{J}^{U}}Y_{j,t}$ is a stronger and less convincing assumption than (2) since $Y_{j,t}$ is a random variable.
-Therefore, in this study, we discuss SCMs based on (2).
-Ferman and Pinto (2021) explicitly considers Assumption 4.1.
+There exists a set $\widetilde{\Phi}\subset\Delta^{J}$ such that for all $\bm{w}^{*}=(w^{*}_{j})_{j\in\mathcal{J}^{U}}\in\widetilde{\Phi}$,
+$\mathbb{E}_{0,t}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}[Y^{N}_{j,t}].$ (2)
+Compared to Assumption 4.1, the assumption $Y_{0,t}=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}Y_{j,t}$ is stronger and less plausible, given that $Y_{j,t}$ is random.
 4.2 Asymptotic Bias in Linear Factor Models
-We consider the linear factor model as defined in Definition 2.1, along with Assumption 4.1. However, under the linear factor model, we adopt the following more specific assumption:
+We now consider the linear factor model introduced in Definition 2.1, along with Assumption 4.1.
+This is a specific case of (2).
+Under the linear factor model, we adopt the following more specialized assumption:
 Assumption 4.2.
-Under the linear factor model from Definition 2.1, there exists a set $\widetilde{\Phi}\subset\Delta^{J}$ such that for all $\bm{w}^{*}=(w^{*}_{j})_{j\in\mathcal{J}^{U}}\in\widetilde{\Phi}$, the following conditions hold:
-$c_{0}=\sum_{j\in[J]}w^{*}_{j}c_{j}\quad\text{and}\quad\mu_{0}=\sum_{j\in[J]}w^{*}_{j}\mu_{j}\ \ \forall t\in\mathcal{T},\ \forall j\in\mathcal{J}.$
-Assumption 4.2 is a specific case of Assumption 4.1; that is, if Assumption 4.1 holds, then Assumption 4.2 also holds under the linear factor model defined in Definition 2.1.
-Under the linear factor model with Assumption 4.2, Ferman and Pinto (2021) finds that the SC estimator $\widehat{\tau}^{\mathrm{Abadie}}_{t}$ is biased and not consistent for $\tau_{0,t}$ when $c_{j}$ is non-zero.
-We further make the following assumption for the linear factor models.
+Under the linear factor model from Definition 2.1, there exists a set $\widetilde{\Phi}\subset\Delta^{J}$ such that for all $\bm{w}^{*}=(w^{*}_{j})_{j\in\mathcal{J}^{U}}\in\widetilde{\Phi}$,
+$c_{0}=\sum_{j\in[J]}w^{*}_{j}c_{j}\quad\text{and}\quad\mu_{0}=\sum_{j\in[J]}w^{*}_{j}\mu_{j}\quad\forall t\in\mathcal{T},\ \forall j\in\mathcal{J}.$
+Assumption 4.2 is a specific case of Assumption 4.1; if Assumption 4.1 holds, then Assumption 4.2 also holds under the linear factor model in Definition 2.1.
+Under these assumptions, Ferman and Pinto (2021) shows that the SC estimator $\widehat{\tau}^{\mathrm{Abadie}}_{0,t}$ suffers from bias and is not consistent for $\tau_{0,t}$ when $c_{j}$ is non-zero.
+We also adopt the following assumption regarding the linear factor model.
 Assumption 4.3 (Common and idiosyncratic shocks).
-For each $j\in\mathcal{J}$, the linear factor models 2.1 satisfy the following conditions as $T_{0}\to\infty$:
+For each $j\in\mathcal{J}$, the linear factor model in Definition 2.1 satisfies the following as $T_{0}\to\infty$:
   • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\lambda_{t}\xrightarrow{\mathrm{p}}0$.
@@
   • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\epsilon_{j,t}\lambda_{t}\xrightarrow{\mathrm{p}}0$.
-Ferman and Pinto (2021) then demonstrates that the SCM proposed by Abadie and Gardeazabal (2003) does not yield a consistent estimator for $\tau_{0,t}$ under these conditions.
+Under Assumptions 4.2 and 4.3, Ferman and Pinto (2021) concludes that the LS-SCM proposed by Abadie and Gardeazabal (2003) does not yield a consistent estimator for $\tau_{0,t}$, as the following proposition.
 Proposition 4.4 (Proposition 1 in Ferman and Pinto (2021)).
-Under Assumptions 4.34.2, it holds that
+Under Assumptions 4.34.2, we have
 $\bm{w}^{\mathrm{Abadie}}\xrightarrow{\mathrm{p}}\overline{\bm{w}}^{\mathrm{Abadie}}\coloneqq\left(\overline{w}^{\mathrm{Abadie}}_{j}\right)_{j\in\mathcal{J}^{U}}$
-as $T_{0}\to\infty$, where $(c_{0},\mu_{0})\neq(\sum_{j\in\mathcal{J}^{U}}c_{j}\overline{w}^{\mathrm{Abadie}}_{j},\sum_{j\in\mathcal{J}^{U}}\mu_{j}\overline{w}^{\mathrm{Abadie}}_{j})$, unless $\sigma^{2}_{\epsilon}=0$ or $\widetilde{\Phi}\cap\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\{\bm{w}^{\top}\bm{w}\}\neq\emptyset$.
-Moreover, for $t\in\mathcal{T}_{1}$, it holds that
+as $T_{0}\to\infty$, where
+$(c_{0},\mu_{0})\neq\left(\sum_{j\in\mathcal{J}^{U}}c_{j}\overline{w}^{\mathrm{Abadie}}_{j},\sum_{j\in\mathcal{J}^{U}}\mu_{j}\overline{w}^{\mathrm{Abadie}}_{j}\right),$
+unless $\sigma^{2}_{\epsilon}=0$ or $\widetilde{\Phi}\cap\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\{\bm{w}^{\top}\bm{w}\}\neq\emptyset$. Furthermore, for $t\in\mathcal{T}_{1}$,
 $\widehat{\tau}^{\mathrm{Abadie}}_{0,t}\xrightarrow{\mathrm{p}}\tau_{0,t}+\lambda_{t}\left(\mu_{0}-\sum_{j\in\mathcal{J}^{U}}\mu_{j}\overline{w}^{\mathrm{Abadie}}_{j}\right)+\left(c_{0}-\sum_{j\in\mathcal{J}^{U}}c_{j}\overline{w}^{\mathrm{Abadie}}_{j}\right)+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}\epsilon_{j,t}\overline{w}^{\mathrm{Abadie}}_{j}\right)$
 as $T_{0}\to\infty$.
-To mitigate this issue, Ferman and Pinto (2021) proposes the demeaned SC estimator
+To alleviate this bias problem, Ferman and Pinto (2021) proposes the demeaned SC estimator:
 $\widehat{\bm{w}}^{\mathrm{FP}}=\operatorname*{arg\,min}_{\widehat{\bm{w}}^{\mathrm{FP}}\in\Delta^{J}}\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left\{\left(Y^{N}_{0,t}-\overline{Y}^{N}_{0}\right)-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)\right\}^{2},$
-where $\overline{Y}^{N}_{j}=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}Y_{j,t}$.
-They estimate $\tau_{0,t}$ as $\widehat{\tau}^{\mathrm{FP}}_{0,t}=Y^{I}_{0,t}-\overline{Y}^{N}_{0}-\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{FP}}_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)$.
-They show the following result.
+where $\overline{Y}^{N}_{j}=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}Y_{j,t}$. They then estimate $\tau_{0,t}$ as
+$\widehat{\tau}^{\mathrm{FP}}_{0,t}=Y^{I}_{0,t}-\overline{Y}^{N}_{0}-\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{FP}}_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right).$
+They show that this approach can reduce the bias seen in the classical LS-SCM estimator. However, there still remains bias in the estimator, which does not vanish even if $T_{0}\to\infty$.
 4.3 Implicit Endogeneity
-In this section, we interpret the asymptotic bias identified by Ferman and Pinto (2021) from the perspective of endogeneity.
-To clarify this issue, we consider a simpler case using a least-squares estimator rather than $\bm{w}^{\mathrm{Abadie}}$:
+We now interpret the asymptotic bias found by Ferman and Pinto (2021) as an endogeneity problem.
+We consider (2), that is, $\mathbb{E}_{0,t}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}[Y^{N}_{j,t}]$, which is more general than the linear factor model.
+To illustrate, we consider a simpler case with the least-squares estimator rather than $\bm{w}^{\mathrm{Abadie}}$:
 $\widehat{\bm{w}}^{\mathrm{LS}}=\operatorname*{arg\,min}_{\bm{w}\in\mathbb{R}^{J}}\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{0,t}-\sum_{j\in\mathcal{J}^{U}}w_{j}Y_{j,t}\right)^{2}.$
-We demonstrate that $\widehat{\bm{w}}^{\mathrm{LS}}$ is asymptotically biased due to endogeneity.
-Define $\epsilon_{j,t}=Y_{j,t}-\mathbb{E}_{j,t}\left[Y_{j,t}\right]$.
-Assuming that (2) holds, the following linear regression model is applicable:
+We show that $\widehat{\bm{w}}^{\mathrm{LS}}$ is asymptotically biased due to endogeneity. Define $\epsilon_{j,t}=Y_{j,t}-\mathbb{E}_{j,t}[Y_{j,t}]$. Under (2), we obtain
 $Y_{0,t}=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}Y_{j,t}+\nu_{t},\quad\text{where}\quad\nu_{t}=-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}+\epsilon_{0,t}.$
-Here, $Y_{0,t}=\mathbb{E}_{0,t}[Y_{0,t}]+\epsilon_{0,t}=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[Y_{j,t}\right]+\epsilon_{0,t}=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}Y_{j,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}+\epsilon_{0,t}$.
-Note that $\mathbb{E}\left[\nu_{t}Y_{j,t}\right]\neq 0$, implying endogeneity due to a correlation between the regressor and the error term.
-Endogeneity is known to introduce bias in the least-squares estimator.
-Define $Q^{*}$ as a $(J\times J)$-matrix whose $(i,j)$-element is $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\mathbb{E}_{i,t}[Y_{i,t}]\mathbb{E}_{j,t}[Y_{j,t}]$.
-Let $\Sigma$ be a $J\times J$ diagonal matrix whose $(j,j)$-element is $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\sigma^{2}_{j,t}$.
+Noting that $\mathbb{E}[\nu_{t}Y_{j,t}]\neq 0$, we see that there is endogeneity, as the regressor and the error term are correlated.
+Endogeneity is known to induce bias in least-squares estimation.
+Let $Q^{*}$ be a $(J\times J)$ matrix whose $(i,j)$-element is $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\mathbb{E}_{i,t}[Y_{i,t}]\mathbb{E}_{j,t}[Y_{j,t}]$, and let $\Sigma$ be a $J\times J$ diagonal matrix whose $(j,j)$-element is $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\sigma^{2}_{j,t}$.
 Theorem 4.5.
@@
 $\overline{\bm{w}}^{\mathrm{LS}}=\bm{w}^{*}-\left(Q^{*}+\Sigma\right)^{-1}\Sigma\bm{w}^{*}.$
-This theorem is derived from Section 5.6 of Greene (2003).
-The result implies that under the assumption of (2), there is an asymptotic bias in $\widehat{\bm{w}}^{\mathrm{LS}}$; that is, least-squares-type estimators do not converge to the true value $\bm{w}^{*}$.
-This section highlights the identification of treatment effects and the problem of implicit endogeneity.
-Though the estimator $\bm{w}^{\mathrm{Abadie}}$ is widely used in existing studies, we find that it suffers from implicit endogeneity.
-This issue has been explicitly or implicitly discussed in previous studies (Chernozhukov et al., 2019; Ferman and Pinto, 2021; Ferman, 2021).
-Independently of our work, Fry (2024) explicitly addresses the bias as an endogeneity problem and applies the instrumental variable method to resolve the issue.
+This result follows from Section 5.6 of Greene (2003) and implies that, under (2), there is an asymptotic bias in $\widehat{\bm{w}}^{\mathrm{LS}}$; in other words, least-squares-type estimators do not converge to the true $\bm{w}^{*}$.
+Thus, the identification of treatment effects faces a challenge due to implicit endogeneity.
+While the estimator $\bm{w}^{\mathrm{Abadie}}$ remains standard in many applications, it inherits this endogeneity issue.
+Such concerns have been noted in prior works (Chernozhukov et al., 2019; Ferman and Pinto, 2021; Ferman, 2021), and Fry (2024) independently frames the bias as an endogeneity problem, proposing an instrumental variable method to address it.
 5 SCM by Density Matching
-As discussed in Section 4, Assumption 4.1 is considered weak or incompatible for identification of $\bm{w}^{*}$ using least squares.
-Therefore, we consider adopting another assumption for identification.
-Specifically, we propose that the density of $Y^{N}_{0,t}$ can be replicated by a weighted sum of the densities of $Y^{N}_{j,t}$ and introduce a SCM based on the density-matching approach.
+As discussed in Section 4, Assumption 4.1 can be weak or incompatible for identifying $\bm{w}^{*}$ when using least squares.
+We, therefore, adopt a different assumption for identification.
+Specifically, we propose that the density of $Y^{N}_{0,t}$ can be replicated by a weighted combination of the densities of $Y^{N}_{j,t}$, and we introduce an SCM based on density matching.
 5.1 Identification and Estimation Strategy
-Section 4 and the results in Ferman and Pinto (2021) imply that combining the assumption $\mathbb{E}_{0,t}\left[Y_{0,t}\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[Y_{j,t}\right]$ with least squares estimators does not yield consistent estimators for $\bm{w}^{*}$.
-Hence, we consider another more realistic and theoretically favorable assumption.
-We begin with the premise that the distribution of the target unit’s outcome, $Y^{N}_{0,t}$, is a linear combination of those of the control units’ outcomes, $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$.
+Section 4 and the results in Ferman and Pinto (2021) indicate that combining
+$\mathbb{E}_{0,t}\left[Y_{0,t}\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[Y_{j,t}\right]$
+with least squares estimators do not yield consistent estimators of $\bm{w}^{*}$. We thus adopt mixture models as a more theoretically favorable but still realistic assumption.
+We begin by positing that the distribution of the target unit’s outcome $Y^{N}_{0,t}$ can be expressed as a linear combination of the distributions of $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$.
 For simplicity, in this subsection, we assume that $Y^{N}_{j,t}$ is a continuous random variable with a probability density function (PDF).
-We then consider a scenario where the density of $Y^{N}_{0,t}$ can be replicated by a weighted sum of the densities of $Y^{N}_{j,t}$:
+Under this scenario, we suppose that the density of $Y^{N}_{0,t}$ can be replicated by a weighted sum of the densities of $Y^{N}_{j,t}$:
 $p^{N}_{0,t}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p^{N}_{j,t}(y)\qquad\forall t\in\mathcal{T}.$ (3)
-We discuss how to estimate SC weights that satisfy (3).
-One approach is to employ the optimal-transport argument, as used by Gunsilius (2023), which estimates SC weights by minimizing the Wasserstein distance.
-However, in this study, we provide a simpler algorithm for estimating SC weights since the method of Gunsilius (2023) requires a procedure that is more complex than the original SCMs.
-Because the mixture model (3) holds, we establish the following equation:
-$\mathbb{E}_{0,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{0,t}\Big{)}\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{j,t}\Big{)}\right]\quad\forall t\in\mathcal{T},\ \ \forall\xi\in\mathbb{R},$ (4)
-where $\sqrt{-1}$ denotes the imaginary number.
-From this equation, we estimate $w^{*}$ using the following methodology.
-We consider the upper bound of the difference between the expected exponential transforms of $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{j,t}$:
-$\int_{\xi\in(-h,h)}\left|\mathbb{E}_{0,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{0,t}\Big{)}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{j,t}\Big{)}\right]\right|\mathrm{d}\xi$
-$=\int_{\xi\in(-h,h)}\left|\sum_{\gamma\in\{0,1,2,\dots,G\}}\frac{\sqrt{-1}^{\gamma}\xi^{\gamma}}{\gamma!}\left(\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right)+\epsilon|\xi^{G}|\right|\mathrm{d}\xi$
-$\leq\int_{\xi\in(-h,h)}\left(\sum_{\gamma\in\{1,2,\dots,G\}}\frac{|\sqrt{-1}^{\gamma}||\xi|^{\gamma}}{\gamma!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|+\epsilon|\xi^{G}|\right)\mathrm{d}\xi$
-$=\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|+\frac{2\epsilon}{G+1}h^{G+1}.$
-This formula implies that by minimizing $\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|+\frac{2\epsilon}{G+1}h^{G+1}$, we can accurately estimate $\bm{w}^{*}$ such that $\int_{\xi in(-h,h)}\left|\mathbb{E}_{0,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{0,t}\Big{)}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{j,t}\Big{)}\right]\right|\mathrm{d}\xi=0$.
-This estimation strategy can be interpreted as a variant of the method of moments, where we estimate SC weights to match all moments between the treated unit and the weighted sum of the untreated units.
+We next discuss estimating SC weights that satisfy (3).
+One approach employs optimal transport to minimize the Wasserstein distance (Gunsilius, 2023).
+In this study, we propose a simpler estimation procedure, as the approach of Gunsilius (2023) involves more complexity than the original SCMs.
+Because the mixture model (3) holds, we obtain:
+$\mathbb{E}_{0,t}\left[\exp\left(\sqrt{-1}\xi Y^{N}_{0,t}\right)\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[\exp\left(\sqrt{-1}\xi Y^{N}_{j,t}\right)\right]\quad\forall t\in\mathcal{T},\ \forall\xi\in\mathbb{R},$
+where $\sqrt{-1}$ denotes the imaginary unit.
+From this, we can estimate $\bm{w}^{*}$ by considering the upper bound of the difference between the expected exponential transforms of $Y^{N}_{0,t}$ and the corresponding weighted sum for $Y^{N}_{j,t}$.
+Specifically, we have
+$\int_{\xi\in(-h,h)}\left|\mathbb{E}_{0,t}\left[\exp(\sqrt{-1}\xi Y^{N}_{0,t})\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\exp(\sqrt{-1}\xi Y^{N}_{j,t})\right]\right|\mathrm{d}\xi$
+$=\int_{\xi\in(-h,h)}\Biggl{|}\sum_{\gamma\in\{0,1,2,\dots,G\}}\frac{\sqrt{-1}^{\gamma}\xi^{\gamma}}{\gamma!}\left(\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right)+\epsilon\left|\xi^{G}\right|\Biggr{|}\mathrm{d}\xi$
+$\leq\int_{\xi\in(-h,h)}\Biggl{(}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{|\sqrt{-1}^{\gamma}||\xi|^{\gamma}}{\gamma!}\left|\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right|+\epsilon|\xi^{G}|\Biggr{)}\mathrm{d}\xi$
+$=\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right|+\frac{2\epsilon}{G+1}h^{G+1}.$
+Minimizing
+$\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|+\frac{2\epsilon}{G+1}h^{G+1}$
+thus yields an accurate estimate of $\bm{w}^{*}$ that enforces
+$\int_{\xi\in(-h,h)}\left|\mathbb{E}_{0,t}\left[\exp(\sqrt{-1}\xi Y^{N}_{0,t})\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\exp(\sqrt{-1}\xi Y^{N}_{j,t})\right]\right|\mathrm{d}\xi=0.$
+This approach can be seen as a variant of the method of moments, where we match all moments of $Y^{N}_{0,t}$ with those of the weighted sum of the untreated units.
 5.2 Core Assumptions
-We make the following core assumption for identification of the SC weights.
+We now state our core assumption for identifying the SC weights.
 Assumption 5.1.
@@
 Assumption 5.2.
-There exists $\widetilde{\Phi}^{*}\in\Delta^{J}$ such that for all $\bm{w}^{*}\in\widetilde{\Phi}^{*}$, it holds that
+There exists $\widetilde{\Phi}^{*}\subset\Delta^{J}$ such that for all $\bm{w}^{*}\in\widetilde{\Phi}^{*}$,
 $F_{Y^{N}_{0,t}}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F_{Y^{N}_{j,t}}(y)\quad\forall y\in\mathbb{R}.$
-Assumption 5.1 allows us to guarantee the existence of the $G$-th moment of $Y^{N}_{j,t}$ for any $G\in\mathbb{N}$.
-Assumption 5.2 also has been utilized in existing studies such as Wan et al.
-(2018), Gunsilius (2023) and Shi et al.
+Assumption 5.1 ensures that $Y^{N}_{j,t}$ has a finite $G$-th moment for any $G\in\mathbb{N}$.
+Assumption 5.2 has been employed in earlier work, including Wan et al.
+(2018), Gunsilius (2023), and Shi et al.
 (2022).
 Wan et al.
-(2018) points out that the SCMs are based on Assumption 5.2.
-Gunsilius (2023) assumes Assumption 5.2 and estimates $\bm{w}^{*}$ by using minimizing the Wasserstein distance.
-Shi et al.
-(2022) investigates the underlying assumptions in the classical SCMs and then find that the combination of fine-grained models and Assumption 5.2 justifies the standard SCMs.
-We introduce the fine-grained models proposed by Shi et al.
-(2022) in Section 8.1.
-Although the mixture model and linear factor models are different, they are closely related, as pointed out by Shi et al.
-(2022) and Nazaret et al.
-(2023).
-In terms of fine-grained models, Shi et al.
-(2022) identifies implicit assumptions underlying linear factor models, and we argue that the result justifies the use of a mixture model in the context of linear factor models.
-Furthermore, Nazaret et al.
-(2023) independently report this finding.
-For further details, see Section 8.1.
+(2018) notes that SCMs effectively rely on Assumption 5.2.
+Gunsilius (2023) similarly assumes (3) and estimates $\bm{w}^{*}$ by minimizing the Wasserstein distance, while Shi et al.
+(2022) introduces fine-grained models that, combined with Assumption 5.2, can justify standard SCMs.
+We provide further discussion of these fine-grained models in Section 8.1.
+Although mixture models and linear factor models differ, they are closely related (Shi et al., 2022; Nazaret et al., 2023).
+In the fine-grained models of Shi et al.
+(2022), certain implicit assumptions underlie linear factor models, suggesting that mixture model assumptions can also be valid for these settings.
+Independently, Nazaret et al.
+(2023) arrive at a similar conclusion.
+We detail these connections in Section 8.1.
 5.3 Estimation of SC Weights
-Under the assumption of mixture models, we estimate the SC weight to match the characteristic functions (CHFs). We refer to our method as the Density Matching SCM (DMSCM).
-Let $h\in(0,1)$ be a parameter determined and fixed by us. Then, we estimate SC weights as
-$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\coloneqq\left(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0},G}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right)\right|.$ (5)
-By using the estimate $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0}}$, we can estimate the treatment effect and distributional treatment effect.
-Treatment effect estimation.
-Once estimating the weights, as well as the standard SCMs, we can estimate the treatment effect as
+Under the mixture model assumption, we propose estimating SC weights by matching characteristic functions (CHFs). We call our approach the Density Matching SCM (DMSCM).
+Let $h\in(0,1)$ be a fixed parameter. We then estimate the SC weights as
+$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\coloneqq\left(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0},G}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|.$ (4)
+Using $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}$, we can then estimate both the treatment effect and the distributional treatment effect as follows.
+5.3.1 Treatment Effect Estimation
+Once $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}$ is obtained, we estimate the treatment effect similarly to standard SCMs:
 $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0},G}Y_{j,t}.$ (5)
-We refer to this estimator as the Density Matching SCM (DMSCM) estimator. In Algorithm 1, we summarize our proposed DMSCM estimator with the D2MSCM estimator introduced in the following section.
-Distributional treatment effect estimation.
-We can also estimate the distributional treatment effect as the counterfactual distribution of $p_{0,t}(y)$ by a bootstrap method.
-We re-sample $L>1$ observations $\{Y^{l,\dagger}_{0}\}^{L}_{l=1}$ from $\{Y_{0,t},Y_{1,t},\dots,Y_{J,t}\}_{t\in\mathcal{T}_{1}}$ as follows: for each $l\in\mathcal{L}\coloneqq\{1,2,\dots,L\}$, (i) for each $j\in\mathcal{J}$, we re-sample $Y^{l,\dagger}_{j}$ from $\{Y_{j,t}\}_{t\in\mathcal{T}_{1}}$; (ii) we choose $Y^{l,\dagger}_{0}$ from $(Y^{l,\dagger}_{1},\dots,Y^{l,\dagger}_{J})$ with probability $(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0}})_{j\in\mathcal{J}^{U}}$.
-In this case, our focus is on $\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p_{0,t}(y)$.
+We call this estimator the Density Matching SCM (DMSCM) estimator. Algorithm 1 summarizes our proposed estimator and also includes the D2MSCM estimator introduced later.
+5.3.2 Distributional Treatment Effect Estimation
+To estimate the distributional treatment effect, we can approximate the counterfactual distribution $p_{0,t}(y)$ via bootstrap.
+Specifically, we resample $L>1$ observations $\{Y^{l,\dagger}_{0}\}_{l=1}^{L}$ from $\{Y_{0,t},Y_{1,t},\dots,Y_{J,t}\}_{t\in\mathcal{T}_{1}}$ as follows: for each $l\in\mathcal{L}\coloneqq\{1,2,\dots,L\}$,
+  1. 1.
+    For each $j\in\mathcal{J}$, resample $Y^{l,\dagger}_{j}$ from $\{Y_{j,t}\}_{t\in\mathcal{T}_{1}}$.
+  2. 2.
+    Choose $Y^{l,\dagger}_{0}$ from $\left(Y^{l,\dagger}_{1},\dots,Y^{l,\dagger}_{J}\right)$ with probability $\left(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0}}\right)_{j\in\mathcal{J}^{U}}$.
+Because our primary focus is often on $\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p_{0,t}(y)$, this bootstrap procedure enables us to approximate the distribution of the counterfactual outcomes.
 Algorithm 1 DMSCM.
@@
 6 Convergence Analysis
-We provide the consistency results about our proposed SC weight estimator. The following theorem describes the results including the consistency for the treatment effect $\tau_{0}$.
+We now provide consistency results for our proposed SC weight estimator. The following theorem presents these results, including consistency for the treatment effect $\tau_{0}$.
 Assumption 6.1.
-Assume that the density function of $\varepsilon_{j,t}$ exists and is bounded for each $j\in\mathcal{J}$ and $t\in\mathcal{T}_{0}$, and that the stochastic process $\{\varepsilon_{j,t}\}^{T_{0}}_{t=1}$ is stationary, strongly mixing, with sum of mixing coefficient bounded by $M$.
+Assume that the density function of $\varepsilon_{j,t}$ exists and is bounded for each $j\in\mathcal{J}$ and $t\in\mathcal{T}_{0}$, and that the stochastic process $\{\varepsilon_{j,t}\}_{t=1}^{T_{0}}$ is stationary, strongly mixing, with the sum of mixing coefficients bounded by $M$.
 Theorem 6.2.
 Suppose that Assumptions 5.15.2 and 6.1 hold. Fix $h\in(0,1)$. Define
-$\widetilde{\Phi}^{\dagger}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|$
-Then, it holds that $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}$ as $T_{0}\to\infty$.
-Additionally, we have $\bm{w}^{*}_{G}\to\bm{w}^{*}\in\widetilde{\Phi}^{*}$ as $G\to\infty$ and
-$\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}\xrightarrow{\mathrm{p}}\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right),$
-as $T_{0}\to\infty$ and $G\to\infty$.
+$\widetilde{\Phi}^{\dagger}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right|.$
+Then, it holds that
+$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}\quad\text{as }T_{0}\to\infty.$
+Additionally, $\bm{w}^{*}_{G}\to\bm{w}^{*}\in\widetilde{\Phi}^{*}$ as $G\to\infty$, and
+$\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}\xrightarrow{\mathrm{p}}\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)\quad\text{as }T_{0}\to\infty\text{ and }G\to\infty.$
 The proof is shown in Appendix A.
-Note that this result can be applied for $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ as a special case.
-Thus, $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ and $\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}$ are asymptotically unbiased under each assumption.
+Note that this result also applies to $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ as a special case.
+Consequently, under the respective assumptions, $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ and $\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}$ are asymptotically unbiased.
 7 Inference
@@
 7.1 Conformal Inference on the Treatment Effect
-We consider statistical inference on the treatment effect using our proposed DMSCM and D2SCM to test a sharp null hypothesis, $H_{0}:\tau_{0,t}=\alpha_{t}$ for some $t\in\mathcal{T}_{t}$ and $\bm{\alpha}=(\alpha_{t})_{t\in\mathcal{T}_{1}}$.
-Here, we focus on the conformal inference of Chernozhukov et al.
-(2021), which has been utilized in existing studies, such as Ben-Michael et al.
+We consider statistical inference for the treatment effect using our proposed DMSCM and D2SCM, focusing on testing a sharp null hypothesis, $H_{0}:\tau_{0,t}=\alpha_{t}$ for some $t\in\mathcal{T}_{t}$ and $\bm{\alpha}=(\alpha_{t})_{t\in\mathcal{T}_{1}}$.
+Specifically, we adopt the conformal inference approach of Chernozhukov et al.
+(2021), which has been used by, for example, Ben-Michael et al.
 (2021) and Ferman and Pinto (2021).
-Our goal is to establish a confidence interval whose $p$-value $\widehat{p}$ satisfies that for any $a\in(0,1)$,
-$\mathrm{Pr}\left(\widehat{p}\leq a\right)\to a$ (7)
-as $T_{0}\to\infty$ and $T_{1}$ is fixed. To achieve this goal, the conformal inference takes the following three steps. First, for the sharp null hypothesis
+Our goal is to construct a confidence interval whose $p$-value $\widehat{p}$ satisfies
+$\mathrm{Pr}\left(\widehat{p}\leq a\right)\to a\quad\text{for all }a\in(0,1)\quad\text{as }T_{0}\to\infty\text{ and }T_{1}\text{ is fixed}.$ (6)
+To achieve this, conformal inference proceeds as follows. First, for the sharp null hypothesis
 $H_{0}:\tau_{0,t}=\alpha_{t},$
-we create adjusted post-treatment outcomes for the treated unit $Y^{N\sharp}_{0,t}=Y^{I}_{0,t}-\alpha_{t}$ and add it to the original dataset as $\left(Y_{0,1},\dots,Y_{0,T_{0}},Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$.
-Then, we apply our proposed method to the extended datasets consisting of
-$\left(Y_{0,1},Y_{0,2},\dots,Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$
-and $(Y_{j,1})_{t\in\mathcal{T},j\in\mathcal{J}^{U}}$ to obtain adjusted weights $\widehat{\bm{w}}(\alpha)$ and a corresponding predictor of the counterfactual outcome, denoted by $\widehat{Y}^{N}_{0,t}(\bm{\alpha})$.
-Finally, we compute the $p$-value by assessing whether the adjusted residual “conforms” with the pretreatment residuals.
-Let $u_{t}(\bm{\alpha})$ be the residual defined as $u_{t}(\bm{\alpha})=Y_{0,t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{0}$ and $u_{t}(\bm{\alpha})=Y^{I}_{0,t}-\alpha_{t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{1}$.
-For $j\in\mathcal{T}$, let $\pi_{j}$ be a permutation of indexes defined as $\pi_{j}(t)=t+j$ if $t+j\leq T$ and $\pi_{j}(t)=t+j-T$ otherwise.
+we create adjusted post-treatment outcomes for the treated unit $Y^{N\sharp}_{0,t}=Y^{I}_{0,t}-\alpha_{t}$ and combine them with the original dataset as
+$\left(Y_{0,1},\dots,Y_{0,T_{0}},Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right).$
+We then apply our proposed method to this augmented dataset (comprising
+$\left(Y_{0,1},\dots,Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$
+and $(Y_{j,1})_{t\in\mathcal{T},j\in\mathcal{J}^{U}}$) to obtain adjusted weights $\widehat{\bm{w}}(\alpha)$ and an adjusted predictor of the counterfactual outcome, denoted $\widehat{Y}^{N}_{0,t}(\bm{\alpha})$.
+Finally, we compute the $p$-value by assessing whether the adjusted residual “conforms” with the pre-treatment residuals.
+Define $u_{t}(\bm{\alpha})=Y_{0,t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{0}$, and $u_{t}(\bm{\alpha})=Y^{I}_{0,t}-\alpha_{t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{1}$.
+For $j\in\mathcal{T}$, let $\pi_{j}$ be a permutation of indices defined by $\pi_{j}(t)=t+j$ if $t+j\leq T$, and $\pi_{j}(t)=t+j-T$ otherwise.
 Let $\Pi=\{\pi_{1},\dots,\pi_{T-1}\}$.
-Then, the $p$-value is defined as $\widehat{p}(\bm{\alpha})=1-\widehat{F}(\bm{\alpha})$, where
+Then the $p$-value is defined as $\widehat{p}(\bm{\alpha})=1-\widehat{F}(\bm{\alpha})$, where
 $\widehat{F}(\bm{\alpha})=\frac{1}{T}\sum_{j\in\mathcal{T}}\mathbbm{1}\left[\frac{1}{T}\sum_{t\in\mathcal{T}}\left|u_{t}(\bm{\alpha})\right|<\frac{1}{T}\sum_{\pi_{j}\in\mathcal{T}}\left|u_{\pi_{j}}(\bm{\alpha})\right|\right].$
-Because $Y^{N}_{0,t}$ is random, by inverting this test, we can construct a confidence interval for $\tau_{0,t}$, which is equivalent to constructing a conformal prediction set (Vovk et al., 2005) for $Y^{N}_{0,T}$.
-The $(1-\xi)$ confidence interval is given as
-$\widehat{C}^{\mathrm{conf}}_{1-\xi}=\left\{\bm{\alpha}\in\mathcal{A}\mid\widehat{p}(\bm{\alpha})>\xi\right\},$
-where $\mathcal{A}$ is a set of candidates of $\bm{a}$. Theorem 1 of Chernozhukov et al. (2021) guarantees (7).
+Because $Y^{N}_{0,t}$ is random, inverting this test yields a confidence interval for $\tau_{0,t}$, which is equivalent to constructing a conformal prediction set (Vovk et al., 2005) for $Y^{N}_{0,T}$.
+The $(1-\xi)$ confidence interval is
+$\widehat{C}^{\mathrm{conf}}_{1-\xi}=\left\{\bm{\alpha}\in\mathcal{A}\bigm{|}\widehat{p}(\bm{\alpha})>\xi\right\},$
+where $\mathcal{A}$ is a set of candidate values. Theorem 1 of Chernozhukov et al. (2021) guarantees (6).
 7.2 Two-Sample Homogeneity Test
-We define the null and alternative hypotheses as $H_{0}:p^{I}_{t}=p^{N}_{t}$ and $H_{1}:p^{I}_{t}\neq p^{N}_{t}$.
-Then, we can conduct hypothesis testing by applying two-sample homogeneity test, and our distributional treatment effect estimator.
-For example, we can employ maximum mean discrepancy (MMD) for testing (Gretton et al., 2012).
+To test distributional treatment effects, consider the null and alternative hypotheses $H_{0}:p^{I}_{t}=p^{N}_{t}$ versus $H_{1}:p^{I}_{t}\neq p^{N}_{t}$.
+One can perform a two-sample homogeneity test in combination with our distributional treatment effect estimator.
+For instance, one may employ maximum mean discrepancy (MMD) (Gretton et al., 2012) to implement this hypothesis test.
 8 Discussion
-This section discusses some remaining problems.
+We conclude by discussing some remaining issues.
 8.1 Justification of Mixture Models
-To elucidates assumptions in SCMs, Shi et al.
-(2022) posits fine-grained models, where each unit has its components, and the outcomes of treated and untreated units are averages of the outcomes of the components.
-For example, in the empirical study of the tobacco control program in California, the outcome of interest is annual per capita cigarette consumption at the state level, which is an average of the consumption of people in the state.
-Based on their arguments, we can justify the assumption of mixture models by connecting them to linear factor models.
-Assume that there are $U_{j}$ components in unit $j\in\mathcal{J}$.
-For each $u\in\mathcal{U}_{j}\coloneqq\{1,2,\dots,U_{j}\}$, let $(Y^{I}_{j,t,u},Y^{N}_{j,t,u})$ an outcome of a component $u$ in unit $j$ at time $t$.
-We also suppose an existence of a $d_{W}$-dimensional unobserved random variable $W_{j,t,u}\in\mathbb{Z}^{d_{W}}$, contributing to individual $u$’s outcome, such as age, education level, or income.
-We refer to $W_{j,t,u}$ as causes.
-Assume that $(Y^{I}_{j,t,u},Y^{N}_{j,t,u},W_{j,t,u})$ is an i.i.d.
-random variable across $u\in\mathcal{U}_{j}$.
-The relationship between the causes $W_{j,t,u}$ and potential outcomes $(Y^{I}_{j,t,u},Y^{U}_{j,t})$ can be linear or non-linear, and can also be time-variant.
-Let $p_{j,t}(w,y)$ be the PDF of $(W_{j,t},Y^{N}_{j,t})$, $p_{j,t}(y\mid w)$ be the PDF of $Y^{N}_{j,t}$ given $W_{j,t}$, and $p_{j,t}$ be the marginalized PDF of $W_{j,t}$.
-The authors assume that interventions are made at a group-level and individuals in each group comply with their group-level intervention.
-Under this setting, Shi et al. (2022) finds that the following assumptions are required for the linearity of SCMs (2).
+To clarify assumptions in SCMs, Shi et al.
+(2022) proposes fine-grained models wherein each unit is composed of multiple components, and the outcomes of treated and untreated units are averages of the outcomes across these components.
+For example, in the empirical study of California’s tobacco control program, the outcome of interest (annual per capita cigarette consumption) at the state level is an average of individuals’ consumption in the state.
+Building on this framework, one can justify mixture model assumptions by relating them to linear factor models.
+Assume there are $U_{j}$ components in unit $j\in\mathcal{J}$.
+For each $u\in\mathcal{U}_{j}\coloneqq\{1,2,\dots,U_{j}\}$, let $(Y^{I}_{j,t,u},Y^{N}_{j,t,u})$ be the potential outcomes of component $u$ in unit $j$ at time $t$.
+Suppose there is a $d_{W}$-dimensional unobserved random variable $W_{j,t,u}\in\mathbb{Z}^{d_{W}}$ capturing individual characteristics (e.g., age or income).
+We assume $(Y^{I}_{j,t,u},Y^{N}_{j,t,u},W_{j,t,u})$ is i.i.d.
+across $u\in\mathcal{U}_{j}$.
+The relationship between $W_{j,t,u}$ and potential outcomes $(Y^{I}_{j,t,u},Y^{N}_{j,t,u})$ can be linear or nonlinear, and possibly time-varying.
+Let $p_{j,t}(w,y)$ be the PDF of $(W_{j,t},Y^{N}_{j,t})$, and let $p_{j,t}(y\mid w)$ be the conditional PDF of $Y^{N}_{j,t}$ given $W_{j,t}$, while $p_{j,t}$ is the marginal PDF of $W_{j,t}$.
+We assume that the intervention is administered at a group level and that individuals within each group comply with their group-level treatment.
+Under these conditions, Shi et al. (2022) shows that the following assumptions are required for the linearity of SCMs in (2):
 Assumption 8.1 (Independent Causal Mechanism; ICM).
 Conditional on the causes $W_{j,t,u}$, the potential outcome $Y^{N}_{j,t,u}$ is independent of the unit $j\in\mathcal{J}$.
-For population distribution $j\in\in\mathcal{J}$ at time $t\leq\mathcal{T}$, the joint distribution of $(Y^{N}_{j,t,u},W_{j,t,u})$ is given as $p_{j,t}(w,y)=r_{j,t}(w)p_{t}(y|w)$.
+For the population distribution $j\in\mathcal{J}$ at time $t\leq\mathcal{T}$, the joint distribution of $(Y^{N}_{j,t,u},W_{j,t,u})$ is given by $p_{j,t}(w,y)=r_{j,t}(w)p_{t}(y\mid w)$.
 Assumption 8.2 (Stable distributions).
-Decompose the causes into two subsets $W_{j,t,u}=\{U_{j,t,u},S_{j,t,u}\}$.
-Let $S$ denote the subset that differentiates the treated unit from the selected untreated units (minimal invariant set); that is, its distribution in the target unit is different from its distribution in the selected untreated units.
-We assume that for all groups, the distribution of $S$ does not change for all time periods $t\leq T$, $r_{j,t}(w)=\zeta_{j}(s)\xi_{t}(u\mid s)$, where $\zeta_{j}(s)$ is the PDF of $S_{j,t,u}$, and $\xi_{t}(u\mid s)$ is the PDF of $U_{j,t,u}$ given $S_{j,t,u}$.
+Decompose the causes as $W_{j,t,u}=\{U_{j,t,u},S_{j,t,u}\}$.
+Let $S$ denote the subset of causes that differentiates the treated unit from the selected untreated units (the minimal invariant set).
+We assume that, for all groups, the distribution of $S$ does not change for all time periods $t\leq T$.
+Hence, $r_{j,t}(w)=\zeta_{j}(s)\xi_{t}(u\mid s)$, where $\zeta_{j}(s)$ is the PDF of $S_{j,t,u}$, and $\xi_{t}(u\mid s)$ is the PDF of $U_{j,t,u}$ given $S_{j,t,u}$.
 Assumption 8.3 (Sufficiently similar untreated units).
 Let $S$ be the minimal invariant set for the target and untreated units.
-The untreated units are sufficiently similar if the cardinality of the minima invariant set, $\big{|}\mathcal{J}^{U}\big{|}\geq|S|$.
+The untreated units are sufficiently similar if the cardinality of this minimal invariant set satisfies $\left|\mathcal{J}^{U}\right|\geq\lvert S\rvert$.
 Assumption 8.4 (Common support).
-Let $(s_{1},\dots,s_{R})$ be the support of $S$. For each $s$, there exists at least one untreated unit $j$, where $\zeta_{j}(S=s)>0$.
+Let $(s_{1},\dots,s_{R})$ be the support of $S$. For each $s$, there is at least one untreated unit $j$ with $\zeta_{j}(S=s)>0$.
 Proposition 8.5 (Causal identifiability. From Theorem 1 of Shi et al. (2022)).
-Suppose that Assumptions 8.18.4 hold. Then, there exists $\{w^{*}_{j}\}_{j\in[J]}$ such that $p^{N}_{0,t}(y)=\sum_{j\in\mathcal{U}^{U}}w^{*}_{j}p^{N}_{j,t}(y)$ holds.
-This proposition also implies that the mixture model holds (Nazaret et al., 2023), which is consistent to the linear factor models.
-Therefore, based on this result, we can obtain the following result for linear factor models as well as Ferman and Pinto (2021).
+Suppose Assumptions 8.18.4 hold. Then there exist weights $\{w^{*}_{j}\}_{j\in[J]}$ such that $p^{N}_{0,t}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p^{N}_{j,t}(y)$ holds.
+This result also implies that the mixture model (3) holds (Nazaret et al., 2023), aligning it with linear factor models.
+Hence, we obtain the following corollary for linear factor models similar to Ferman and Pinto (2021):
 Corollary 8.6.
-Suppose that Assumptions 5.1, 6.1, and 8.18.4 hold.
-Then, $\bm{w}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}$ holds as $T_{0}\to\infty$, and $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$.
-Additionally, $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ converge to $\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)$ in probability as $T_{0}\to\infty$ and $G\to\infty$.
-The proof is straightforward. From Proposition 8.5, under Assumptions 8.18.4, (3) holds. Therefore, from Theorem 6.2, the claim holds under (1) with $c_{j}=0$ for $j\in\mathcal{J}$.
+Suppose that Assumptions 5.1, 6.1, and 8.18.4 hold. Then
+$\bm{w}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\quad\text{as }T_{0}\to\infty,$
+and $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ converges to
+$\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)$
+in probability as $T_{0}\to\infty$ and $G\to\infty$.
+Proof Sketch. From Proposition 8.5, under Assumptions 8.18.4, (3) holds. Therefore, applying Theorem 6.2 yields the desired result under (1) with $c_{j}=0$ for each $j\in\mathcal{J}$.
 8.2 Auxiliary Covariates
-We can consider a case where auxiliary covariates are observable. For each unit $j\in\mathcal{J}$, let
-$X_{j,t}\coloneqq(X_{j,t,1},X_{j,t,1},\dots,X_{j,t,K})\in\mathbb{R}^{K}$
-be $K$-dimensional covariates, which is not affected by treatment for all $j\in\mathcal{J}$. Let
-^$\widehat{\ell}_{k}(\bm{w})=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left\{X_{0,t,k}-\sum_{j\in\mathcal{J}^{U}}w_{j}X_{j,t,k}\right\}.$
-When the auxiliary covariates are observable, we extend the DMSCM estimator as
-$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0}}\coloneqq v_{0}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right)\right|+\sum^{K}_{k=1}v_{k}\widehat{\ell}^{2}_{k}(\bm{w}),$
-where $\bm{v}\coloneqq(v_{k})_{k\in\{0,1,2,\dots,K\}}$ are some weights such that $\sum^{K}_{k=0}v_{0}=1$.
-8.3 Distributional SCMs and Optimal Transports
-Gunsilius (2023) establies the framework of distributional SCMs. In the framework, that study assumes the following:
-$F^{-1}_{Y_{0,t},N}(q)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F^{-1}_{Y_{j,t},N}(q),\ \ \ \forall q\in(0,1),$
-where $F^{-1}_{Y_{j,t},N}(q)$ is the quantile function defined as $F^{-1}_{Y_{j,t},N}(q)\coloneqq\inf_{y\in\mathbb{R}}\big{\{}F_{Y_{j,t},N}(q)\geq q\big{\}}$.
-Then, Gunsilius (2023) estimates the SC weight to match the distribution under the Wasserstein distance.
-Their method is called DiSCo.
-Specifically, they estimate the SC weights as
-$\widehat{\bm{w}}^{\mathrm{DiSCo}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\frac{1}{M}\sum^{M}_{m=1}\left|\widehat{F}^{-1}_{Y^{N}_{0,t}}(V_{m})-\sum_{j\in\mathcal{J}^{U}}w_{j}\widehat{F}^{-1}_{Y^{N}_{j,t}}(V_{m})\right|,$
-where $\widehat{F}^{-1}_{Y^{N}_{j,t}}(V_{m})$ is an empirical quantile function of $Y^{N}_{j,t}$, and $\{V_{m}\}$ is a set of $M$ independent draws from the uniform distribution with support $[0,1]$.
+We can also extend our method to incorporate auxiliary covariates if they are observable. For each unit $j\in\mathcal{J}$, let
+$X_{j,t}\coloneqq(X_{j,t,1},X_{j,t,2},\dots,X_{j,t,K})\in\mathbb{R}^{K}$
+be $K$-dimensional covariates that are unaffected by treatment. Define
+^$\widehat{\ell}_{k}(\bm{w})=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(X_{0,t,k}-\sum_{j\in\mathcal{J}^{U}}w_{j}X_{j,t,k}\right).$
+When these covariates are available, we can extend the DMSCM estimator as
+$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0}}\coloneqq v_{0}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|+\sum_{k=1}^{K}v_{k}\widehat{\ell}^{2}_{k}(\bm{w}),$
+where $\bm{v}\coloneqq(v_{k})_{k=0}^{K}$ are weights such that $v_{k}\geq 0$ and $\sum_{k=0}^{K}v_{k}=1$.
+8.3 Distributional SCMs and Optimal Transport
+Gunsilius (2023) propose distributional SCMs, which assume
+$F^{-1}_{Y_{0,t},N}(q)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F^{-1}_{Y_{j,t},N}(q)\quad\forall q\in(0,1),$
+where $F^{-1}_{Y_{j,t},N}$ is the quantile function defined by
+$F^{-1}_{Y_{j,t},N}(q)\coloneqq\inf\left\{y\in\mathbb{R}\colon F_{Y_{j,t},N}(y)\geq q\right\}.$
+To estimate $\bm{w}^{*}$, Gunsilius (2023) minimize the Wasserstein distance, referring to their method as DiSCo. Concretely, the DiSCo estimator is
+$\widehat{\bm{w}}^{\mathrm{DiSCo}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\frac{1}{M}\sum_{m=1}^{M}\left|\widehat{F}^{-1}_{Y^{N}_{0,t}}(V_{m})-\sum_{j\in\mathcal{J}^{U}}w_{j}\widehat{F}^{-1}_{Y^{N}_{j,t}}(V_{m})\right|,$
+where $\widehat{F}^{-1}_{Y^{N}_{j,t}}(V_{m})$ is the empirical quantile function of $Y^{N}_{j,t}$, and $\{V_{m}\}_{m=1}^{M}$ are i.i.d. draws from the uniform distribution on $[0,1]$.
 8.4 Demeaned Density Matching SCM (D2MSCM)
-Next, we consider a case where the target distribution does not follow mixture models, but the deviation is zero when it is marginalized.
-Instead of (3), We consider the following another mixture model
+Finally, we consider the case where the target distribution does not precisely satisfy (3), yet any deviation disappears after demeaning. Concretely, instead of (3), we assume
 $p^{N}_{0,t}(y-c_{0})=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p^{N}_{j,t}(y-c_{j})\qquad\forall t\in\mathcal{T}.$
-From the viewpoints of linear factor models 1 and fine-grained models (Shi et al., 2022), this assumption allows us $c_{j}\neq 0$ for some $j\in\mathcal{J}$.
-Under the previous mixture model, we only consider a case where $c_{j}=0$ for all $j\in\mathcal{J}$, which is restrictive in practice.
-To estimate the SC weight, we first estimate $c_{j}$ by taking the average of $(Y_{j,t})_{t\in\mathcal{T}_{0}}$ and subtract it from $Y_{j,t}$.
-This manipulation corresponds to the demeaned estimation in Ferman and Pinto (2021).
-Let $h\in(0,1)$ be a parameter determined and fixed by us.
-Then, we estimate SC weights as
-$\widehat{\bm{w}}^{\mathrm{D2MSCM}}_{T_{0}}\coloneqq\left(\widehat{w}^{\mathrm{D2MSCM}}_{j,T_{0}}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{0,t}-\overline{Y}^{N}_{0}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}-\overline{Y}^{N}_{0}\Big{)}^{\gamma}\right)\right|.$
-We refer to this method as the demeaned DMSCM (D2SCM). By using the D2MSCM estimator $(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0}})_{j\in\mathcal{J}^{U}}$, we estimate the treatment effect as
-$\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\left(\overline{Y}^{N}_{0}+\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{D2SCM}}_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)\right).$ (8)
-9 Experiments
-We conduct experiments with simulated and real-world data. Additional results are in Appendix 9.1.
-9.1 Simulation Studies
-In this section, we perform simulation studies to show how our proposed SCMs compared to traditional SCMs.
-Let $J\in\{10,30,60\}$ and $G\in\{2,3,5,10,20,30,40,50,60,70,80,90,100\}$, $T_{0}=30$, $T_{1}=100$, and $K=5$.
-We perform a simulation on SCMs for $p_{0,t}(\bm{z})=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p_{0,t}(\bm{z})$, where $\bm{z}$ is ($K+1$)-dimensional ($6$-dimensional) vector, where the first element corresponds to $Y_{j,t}$ and the rest corresponds to $X_{j,t}$.
-For each $j\in\mathcal{J}$, let $p_{j,t}$ be a density of a multinomial normal distribution $\mathcal{N}(\bm{\mu}_{j,t},\bm{\sigma}^{2}_{j,t})$, where $\bm{\mu}_{j,t}=(\mu_{j,t,1},\mu_{j,t,2},\mu_{j,t,3},\mu_{j,t,4},\mu_{j,t,5},\mu_{j,t,6})$, and $\bm{\sigma}^{2}_{j,t}$ is a $(6\times 6)$-diagonal matrix with a $(k,k)$-element denoted by $\sigma^{2}_{j,t,k}$.
-Each mean parameter is drawn from a standard normal; distribution.
-Each variance parameter is drawn from a uniform distribution with a support $[1,20]$.
-Let us assume that for each $j\in\mathcal{J}$, $\mu_{j,t,1}=\mu_{j,t,1}+\epsilon_{j,t}$ and $\sigma^{2}_{j,t,1}=\sigma^{2}_{j,t,1}+\widetilde{\epsilon}_{j,t}$, where $\epsilon_{j,t}$ and $\widetilde{\epsilon}_{j,t}$ follow a normal densities $\mathcal{N}(0,10)$, respectively.
-Let us correct the variance as $\widetilde{\epsilon}_{j,t}=0.1$ if $\widetilde{\epsilon}_{j,t}\leq 0.1$ occurs.
-For the treated unit, we assume that $Y^{N}_{0,t}\sim p_{0,t}(\bm{z})$ for $t\in[t_{0},T_{1}]$ and $Y^{I}_{0,t}=20+Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$; that is $\tau_{0,t}=20$.
-We draw $\bm{w}^{*}$ from a uniform distribution and normalize the sum to be $1$.
-We apply DMSCM and Abadie $1,000$ times and compute the error of the treatment effect estimation.
-We show the result with boxplot in Figure 4.
-Our proposed method successfully estimate the treatment effect with smaller estimation error.
-The estimation decreases as $G$ increases.
-We conduct simulation studies to illustrate the comparative performance of our proposed SCMs and traditional SCMs within a setting analogous to Section 9.1.
-The only modifications introduced involve the parameters.
-Specifically, $J\in\{1,5,10,15,20,\dots,35,40,45,50\}$, $G\in\{2,3,5,10\}$, $T_{0}=30$, $T_{1}=1000$, and $K=5$.
-All other conditions align with those detailed in Section 9.1.
-We apply DMSCM and Abadie $1,000$ times and compute the error of the treatment effect estimation in Figure 7.
-Our focus lies on the fluctuation in mean squared estimation errors as the number of units, $J$, increases.
-Despite the method by Abadie (2002) incurring escalating estimation errors with the increase of $J$, estimation errors do not exhibit a similar growth under our proposed methods.
-Furthermore, our proposed methods exhibit lower estimation errors compared to those under the method of Abadie (2002).
-A comparable phenomenon is evident in distributional treatment effect estimation.
-We compare the true and estimated distributions using the Maximum Mean Discrepancy (MMD, Gretton et al., 2012), a probability distance computed through kernel embedding.
-The upper figure in Figure 8 displays how MMDs alter as $J$ increases.
-As illustrated, our proposed methods outperform even in this task.
-The lower figure in Figure 8 presents results for $G\in\{2,10,50,100\}$.
-As the results indicate, the predictive capability augments as $G$ (the number of moment conditions) increases.
-Our findings might have a relationship with those of Spiess et al. (2023), who investigates high-dimensional SCMs. Future investigations will be devoted to exploring these connections.
+From the perspective of the linear factor model (1) and fine-grained models (Shi et al., 2022), this assumption permits $c_{j}\neq 0$ for some $j\in\mathcal{J}$.
+Under the previous mixture model, we implicitly required $c_{j}=0$ for all $j\in\mathcal{J}$, which may be restrictive in practice.
+To estimate the SC weights in this setting, we first estimate $c_{j}$ by taking the average of $(Y_{j,t})_{t\in\mathcal{T}_{0}}$ and subtracting it from each $Y_{j,t}$.
+This demeaned approach is inspired by the proposal of Ferman and Pinto (2021) but is applied here to density matching.
+Let $h\in(0,1)$ be a fixed parameter.
+We estimate the weights as
+$\widehat{\bm{w}}^{\mathrm{D2MSCM}}_{T_{0}}$ $\coloneqq\left(\widehat{w}^{\mathrm{D2MSCM}}_{j,T_{0}}\right)_{j\in\mathcal{J}^{U}}$
+$\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\biggl{|}\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}-\overline{Y}^{N}_{0}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)^{\gamma}\right)\biggr{|}.$
+We call this procedure demeaned DMSCM (D2SCM). Using these weights, we can estimate the treatment effect as
+$\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\biggl{(}\overline{Y}^{N}_{0}+\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{D2MSCM}}_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)\biggr{)}.$
 Figure 2: Results of simulation studies.
-Figure 3: Counterfactual outcomes in empirical analysis.
-Figure 4: Counterfactual distributions in empirical anaylsis.
-Table 2: Results of conformal inference. Estimated treatment effect $\widehat{\tau}_{0,t}$ and $90\%$ confidence interval $(\underline{\tau}_{0,t},\overline{\tau}_{0,t})$.
-Terrorist conflict in Basque country Tobacco control program in California Reunification of Germany
-Abadie DMSCM Abadie DMSCM Abadie DMSCM
-$\gamma$ $2$ $5$ $10$ $100$ $2$ $5$ $10$ $100$ $2$ $5$ $10$ $100$
-$\widehat{\tau}_{0,t}$0.800 -0.845 -1.054 -0.279 -0.138 -26.148 -27.744 -32.306 -35.080 -36.879 -1492.374 -1253.616 -1419.120 -1324.448 -510.765
-$\overline{\tau}_{0,t}$ 0.522 0.699 0.495 0.521 0.578 8.698 7.852 8.898 9.544 11.748 2635.583 -1008.822 -591.420 -716.376 2327.944
-$\underline{\tau}_{0,t}$2.587 -2.150 -2.605 -2.616 -2.590 -36.798 -29.268 -34.783 -37.308 -35.243 -2390.413 -2738.232 -1989.323 -2018.878 -2625.129
 Figure 3: Results of simulation studies on ATE prediction. The $x$-axis is the number of untreated units $J$, and the $y$-axis is the mean squared prediction error.
 Figure 4: Results of simulation studies on distributional treatment effect prediction. The $x$-axis is the number of untreated units $J$, and the $y$-axis is the MMD.
-Figure 7: Results of simulation studies on ATE prediction. The $x$-axis is the number of untreated units $J$, and the $y$-axis is the mean squared prediction error.
-Figure 8: Results of simulation studies on distributional treatment effect prediction. The $x$-axis is the number of untreated units $J$, and the $y$-axis is the MMD.
+9 Experiments
+We conduct experiments using both simulated and real-world data. Additional results can be found in Appendix 9.1.
+9.1 Simulation Studies
+We first perform simulation studies to compare our proposed SCMs with traditional SCMs. Set $J\in\{10,30,60\}$ and $G\in\{2,3,5,10,20,30,40,50,60,70,80,90,100\}$, $T_{0}=30$, $T_{1}=100$, and $K=5$.
+We consider SCMs for
+$p_{0,t}(\bm{z})=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p_{j,t}(\bm{z}),$
+where $\bm{z}$ is a $(K+1)$-dimensional (i.e., $6$-dimensional) vector: the first element corresponds to $Y_{j,t}$ and the remaining elements correspond to $X_{j,t}$.
+For each $j\in\mathcal{J}$, let $p_{j,t}$ be the density of a multivariate normal distribution
+$\mathcal{N}\left(\bm{\mu}_{j,t},\bm{\sigma}^{2}_{j,t}\right),$
+where $\bm{\mu}_{j,t}=(\mu_{j,t,1},\dots,\mu_{j,t,6})$ and $\bm{\sigma}^{2}_{j,t}$ is a $(6\times 6)$ diagonal matrix whose $(k,k)$-element is $\sigma^{2}_{j,t,k}$.
+Each $\mu_{j,t,k}$ is drawn from a standard normal distribution, and each $\sigma^{2}_{j,t,k}$ is drawn from a uniform distribution on $[1,20]$.
+For each $j\in\mathcal{J}$, we modify $\mu_{j,t,1}$ and $\sigma^{2}_{j,t,1}$ by adding noise:
+$\mu_{j,t,1}\leftarrow\mu_{j,t,1}+\epsilon_{j,t},\quad\sigma^{2}_{j,t,1}\leftarrow\sigma^{2}_{j,t,1}+\widetilde{\epsilon}_{j,t},$
+where $\epsilon_{j,t}$ and $\widetilde{\epsilon}_{j,t}$ follow $\mathcal{N}(0,10)$.
+If $\widetilde{\epsilon}_{j,t}$ is less than or equal to $0.1$, we set $\widetilde{\epsilon}_{j,t}=0.1$ to avoid near-zero variance.
+For the treated unit, we let $Y^{N}_{0,t}\sim p_{0,t}(\bm{z})$ for $t\in[t_{0},T_{1}]$ and set $Y^{I}_{0,t}=20+Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$, implying $\tau_{0,t}=20$.
+We draw $\bm{w}^{*}$ from a uniform distribution over $[0,1]^{J}$ and then normalize it so that its entries sum to $1$.
+We apply DMSCM and the method of Abadie and Gardeazabal (2003) (referred to here as Abadie) $1000$ times and evaluate the estimation error of the treatment effect.
+We plot the estimation error box plots in Figure 3, indicate that our proposed method achieves smaller estimation errors.
+Moreover, the estimation error declines as $G$ increases.
+Next, we consider another set of simulation studies under conditions similar to those above, but with the following changes: $J\in\{1,5,10,15,20,\dots,35,40,45,50\}$, $G\in\{2,3,5,10\}$, $T_{0}=30$, $T_{1}=1000$, and $K=5$.
+All other settings remain as in Section 9.1.
+We again apply DMSCM and Abadie $1000$ times and compute the error of the treatment effect estimation (Figure 3).
+We focus on how the mean squared estimation error changes as $J$ grows.
+While the estimation error under the method of Abadie (2002) grows with $J$, our proposed method does not exhibit such growth.
+Additionally, our method maintains lower estimation errors than the method of Abadie (2002).
+A similar pattern appears in distributional treatment effect estimation.
+We measure the discrepancy between the true and estimated distributions using the Maximum Mean Discrepancy (MMD, Gretton et al., 2012), a probability metric based on kernel embeddings.
+In the top panel of Figure 4, we show how MMD varies with $J$.
+As before, our proposed methods outperform the method of Abadie (2002) in this setting.
+In the bottom panel of Figure 4, we plot MMD for $G\in\{2,10,50,100\}$.
+As $G$ grows (i.e., as more moment conditions are imposed), the predictive performance improves.
+Our observations may connect with findings by Spiess et al. (2023), who examine high-dimensional SCMs. Exploring these connections will be part of our future research.
+Figure 5: Counterfactual outcomes in empirical analysis.
+Figure 6: Counterfactual distributions in empirical analysis.
+Table 2: Results of treatment effect estimation and conformal inference.
+Estimated treatment effect $\widehat{\tau}_{0,t}$ and $90\%$ confidence interval $(\underline{\tau}_{0,t},\overline{\tau}_{0,t})$.
+Terrorist conflict in Basque country.
+Abadie DMSCM
+$\gamma$- $2$ $5$ $10$ $100$
+$\widehat{\tau}_{0,t}$ $-0.800$ $-0.845$ $-1.054$ $-0.279$ $-0.138$
+$\overline{\tau}_{0,t}$ $0.522$ $0.699$ $0.495$ $0.521$ $0.578$
+$\underline{\tau}_{0,t}$ $-2.587$ $-2.150$ $-2.605$ $-2.616$ $-2.590$
+Tobacco control program in California.
+Abadie DMSCM
+$\gamma$- $2$ $5$ $10$ $100$
+$\widehat{\tau}_{0,t}$ $-26.148$ $-27.744$ $-32.306$ $-35.080$ $-36.879$
+$\overline{\tau}_{0,t}$ $8.698$ $7.852$ $8.898$ $9.544$ $11.748$
+$\underline{\tau}_{0,t}$ $-36.798$ $-29.268$ $-34.783$ $-37.308$ $-35.243$
+Reunification of Germany.
+Abadie DMSCM
+$\gamma$- $2$ $5$ $10$ $100$
+$\widehat{\tau}_{0,t}$ $-1492.374$ $-1253.616$ $-1419.120$ $-1324.448$ $-510.765$
+$\overline{\tau}_{0,t}$ $2635.583$ $-1008.822$ $-591.420$ $-716.376$ $2327.944$
+$\underline{\tau}_{0,t}$ $-2390.413$ $-2738.232$ $-1989.323$ $-2018.878$ $-2625.129$
 9.2 Empirical Analysis
-We conduct empirical studies discussed in Abadie and Gardeazabal (2003), Abadie et al. (2010), and Abadie et al. (2015).
+We now present empirical studies based on Abadie and Gardeazabal (2003), Abadie et al. (2010), and Abadie et al. (2015).
 Terrorist conflict in the Basque Country.
-Abadie and Gardeazabal (2003) investigates the effect of terrorist conflict in the Basque Country on gross domestic product (GDP).
+Abadie and Gardeazabal (2003) investigate the effect of terrorist conflict in the Basque Country on gross domestic product (GDP).
 In this example, $J=16$, $K=14$, and $(t_{0},T_{0},T_{1})=(1955,1970,1997)$.
-Basque Euskadi Ta Askatasuna (Basque Homeland and Liberty; ETA) was founded in 1959 as a Basque separatist organization in Spain.
-ETA promoted an independent Basque state by carrying out terrorist activities mainly in the Basque Country region of Spain.
-Abadie and Gardeazabal (2003) utilizes annual Spanish regional-level panel data from 1955 to 1997, with a particular focus on the 1970s when ETA conducted significant terrorist activity.
-In the data, 15 years of pre-intervention data are included.
-The set of untreated units comprises $17$ other Spanish regions.
-Each region’s datum contains the outcome of interest (GDP per capita) at the regional level, with covariates including investment rate (percentage), population density, five sectoral productions, and four human capitals, all averaged over the period from 1955 to 1968.
+Basque Euskadi Ta Askatasuna (Basque Homeland and Liberty; ETA) was founded in 1959 as a Basque separatist organization in Spain, seeking an independent Basque state primarily through terrorist activities in the Basque Country region.
+Abadie and Gardeazabal (2003) use annual Spanish regional-level panel data from 1955 to 1997, with a particular focus on the 1970s when ETA’s terrorist activity was significant.
+The dataset includes fifteen years of pre-intervention observations.
+The set of untreated units consists of $17$ other Spanish regions.
+Each region’s record contains the outcome of interest (GDP per capita) and additional covariates (investment rate, population density, five sectoral output measures, and four measures of human capital), all averaged from 1955 to 1968.
 Tobacco control program in California.
 Abadie et al.
-(2010) study the effect of a large tobacco control program adopted in California in $1988$, which we explained in our introduction.
-In this example, $J=20$, $K=5$, and $(t_{0},T_{0},T_{1})=(1970,1989,2000)$.
-Abadie et al.
-(2015) investigates the treatment effect of the reunification of Germany.
-East Germany was absorbed into West Germany, the result being a reunited Germany in 1990.
-They utilize annual panel data at the country level from 1960 to 2003.
-The outcome of interest is GDP per capita at the country level.
-The pre-intervention period spans $30$ years.
-The set of untreated units comprises $16$ OECD countries that make up synthetic West Germany.
-Covariates include inflation rate, industry share, investment rate, schooling, and a measure of trade openness.
+(2010) study the effect of a large tobacco control program adopted in California in 1988 (described in the Introduction).
+Here, $J=20$, $K=5$, and $(t_{0},T_{0},T_{1})=(1970,1989,2000)$.
 Reunification of Germany.
 Abadie et al.
-(2015) studies the effect of the reunification of Germany on GDP in 1990.
-Because the countries were dissimilar from any one country to make a comparison group, they apply SCMs to create a composite comparison group.
+(2015) examine the impact of Germany’s reunification in 1990.
+Because no single country provided a suitable comparison, SCMs were used to create a composite unit.
 In this example, $J=8$, $K=31$, and $(t_{0},T_{0},T_{1})=(1960,1990,2003)$.
-Abadie et al.
-(2010) investigate the causal effect of Proposition 99, a comprehensive tobacco control program adopted by California.
-They utilize annual state-level panel data from 1970 to 2000.
-The program was implemented in January 1989, providing 18 years of pre-intervention data.
-Four states that have adopted large-scale tobacco control programs (Massachusetts, Arizona, Oregon, and Florida) and states that increased state cigarette taxes by $50$ cents or more between 1989 and 2000 (Alaska, Hawaii, Maryland, Michigan, New Jersey, New York, and Washington) and the District of Columbia are excluded from the untreated units of 38 states.
-The outcome of interest is annual per capita cigarette consumption at the state level, measured as per capita cigarette sales in packs.
-Covariates include the average retail price of cigarettes, state income per capita (log), percentage of the population aged 15-24, and beer consumption per capita; all averaged over the period from 1980 to 1988.
-Results.
-We show the result of the estimated treatment effect and $90$% confidence interval in Table 2.
-We show the results of treatment effect estimation in Figure 4 and distributional treatment effect estimation in Figure 4.
-For each dataset, we plot the counterfactual outcomes and distributions in Figures 4 and 4, respectively.
-We also conduct conformal inference on the sharp null hypothesis $H_{0}:\tau_{0,(T_{0}+1)}=\cdots=\tau_{0,T_{1}}=0$ for the treatment effect.
-For the results of Abadie, DMSCM, We show the $90\%$ confidence interval of the treatment effect in Table 2.
-Our proposed method has a good fit in terms of the observed outcomes in pretreatment periods $t\in\mathcal{T}_{0}$.
-Therefore, we consider that our proposed method also predicts counterfactual outcomes and distributions well.
-In conformal inference, our method does not reject the null hypothesis as with Abadie, except for the DMSCM in the tobacco control problem in California.
+Separately, Abadie et al.
+(2010) also investigate the causal effect of Proposition 99, a comprehensive tobacco control program adopted by California.
+They use annual state-level panel data from 1970 to 2000.
+The program was implemented in January 1989, providing eighteen years of pre-intervention observations.
+Four states that adopted large-scale tobacco control programs (Massachusetts, Arizona, Oregon, and Florida) and states that raised cigarette taxes by at least \$0.50 between 1989 and 2000 (Alaska, Hawaii, Maryland, Michigan, New Jersey, New York, and Washington) as well as the District of Columbia are excluded from the set of 38 control states.
+The outcome of interest is annual per capita cigarette sales in packs; covariates include the average retail price of cigarettes, log of state income per capita, the percentage of the population aged 15–24, and per capita beer consumption, all averaged between 1980 and 1988.
+Results. Table 2 reports the estimated treatment effects and 90% confidence intervals. Figures 6 and 6 show the estimated treatment effects and the distributional treatment effects, respectively.
+In each dataset, we plot the counterfactual outcomes and distributions in Figures 6 and 6.
+We also perform conformal inference on the sharp null hypothesis $H_{0}:\tau_{0,(T_{0}+1)}=\cdots=\tau_{0,T_{1}}=0$ for the treatment effect.
+For both Abadie’s method and DMSCM, we present the 90% confidence interval of the treatment effect in Table 2.
+Our proposed method fits the observed outcomes well in the pre-treatment period ($t\in\mathcal{T}_{0}$) and, accordingly, produces plausible counterfactual outcomes and distributions.
+In the conformal inference analysis, our method does not reject the null hypothesis in the same manner as Abadie’s estimator, except for the DMSCM in the California tobacco control example.
 10 Conclusion
-In this study, we identified the problem of implicit endogeneity in the linear regression models commonly used in existing SCM studies.
-To address this issue, we proposed a novel approach to SCMs using density matching.
-Our proposed method enables consistent estimation of SC weights and improves the accuracy of counterfactual outcome predictions compared to existing methods.
-We validated the effectiveness of our approach through simulation experiments and empirical analyses.
+We have highlighted the problem of implicit endogeneity in the linear regression models commonly employed by existing SCMs.
+To address this issue, we proposed assuming mixture models among outcome distributions and developed a density-matching SCM.
+By matching densities, our method consistently estimates SC weights and reinforces the theoretical foundation for using SCMs in comparative studies.
+We demonstrated the effectiveness of this approach through both simulation experiments and empirical applications.
 Appendix A Proof of Theorem 6.2
-First, we confirm the basic fact about the uniqueness of the CHF.
+First, we confirm the basic facts about the uniqueness of the CHF.
 Proposition A.1 (Theorem 26.2 in Billingsley (1995)).

### v5

@@
-Asymptotically Unbiased Synthetic Control Methods by Density Matching
+Asymptotically Unbiased Synthetic Control Methods by Moment Matching
 Abstract
@@
 In this study, we highlight a key endogeneity issue in existing SCMs—namely, the correlation between the outcomes of untreated units and the error term of the synthetic control, which leads to bias in both counterfactual outcome prediction and treatment effect estimation.
-To address this issue, we propose a novel SCM based on density matching, assuming that the outcome density of the treated unit can be approximated by a weighted mixture of the joint density of untreated units.
+To address this issue, we propose a novel SCM based on moment matching, assuming that the outcome distribution of the treated unit can be approximated by a weighted mixture of the distributions of untreated units.
 Under this assumption, we estimate SC weights by matching the moments of the treated outcomes with the weighted sum of the moments of the untreated outcomes.
-Our method offers three advantages: first, under the mixture model assumption, our estimator is asymptotically unbiased; second, this asymptotic unbiasedness reduces the mean squared error in counterfactual predictions; and third, our method provides full densities of the treatment effect rather than just expected values, thereby broadening the applicability of SCMs.
+Our method offers three advantages: first, under the mixture model assumption, our estimator is asymptotically unbiased; second, this asymptotic unbiasedness reduces the mean squared error in counterfactual predictions; and third, our method provides full distributions of the treatment effect rather than just expected values, thereby broadening the applicability of SCMs.
 Finally, we present experimental results that demonstrate the effectiveness of our approach.
-[^1] This study was presented at the Workshop on Counterfactuals in Minds and Machines at the International Conference on Machine Learning in July 2023 and at the International Conference on Econometrics and Statistics in August 2023.
 1 Introduction
-Synthetic control methods (SCMs, Abadie and Gardeazabal, 2003; Abadie et al., 2010) have emerged as a crucial tool for program evaluation in comparative case studies across disciplines such as economics, statistics, and machine learning.
-As noted by Athey and Imbens (2017), SCMs represent “arguably the most important innovation in the policy evaluation literature in the last 15 years” due to their “simplicity of the idea, and the obvious improvement over standard methods.”
-In SCMs, there are multiple units, where one unit receives a policy intervention (treated unit) at a certain time while the remaining units remain untreated.
+Synthetic control methods (SCMs, Abadie & Gardeazabal, 2003; Abadie et al., 2010) have emerged as a crucial tool for program evaluation in comparative case studies across disciplines such as economics, statistics, and machine learning.
+In SCMs, there are multiple units, where one unit receives a policy intervention (treated unit) at a certain time while the remaining unitts remain untreated.
 The objective is to estimate the treatment effect for the treated unit, defined as the difference between its observed factual outcome and its unobserved counterfactual outcome.
@@
 See Figure 1 for an illustration.
-d SCMs have been widely applied in various empirical studies, such as evaluating the impact of terrorism on GDP (Abadie and Gardeazabal, 2003) and the decriminalization of indoor prostitution (Cunningham and Shah, 2017).
+SCMs have been widely applied in various empirical studies, such as evaluating the impact of terrorism on GDP (Abadie & Gardeazabal, 2003) and the decriminalization of indoor prostitution (Cunningham & Shah, 2017).
 For instance, Abadie et al.
@@
 A key challenge in SCMs is the assumption, made in most existing studies (Abadie, 2002), that the expected outcome of the treated unit can be represented as a weighted sum of the outcomes of untreated units.
-Ferman and Pinto (2021) highlights that this assumption leads to a non-negligible bias in the estimator.
-According to Ferman and Pinto (2021), this bias arises from endogeneity in the linear regression model implied by the SCM assumption due to the correlation between the outcomes of untreated units and the error term.
+Ferman & Pinto (2021) highlights that this assumption leads to a non-negligible bias in the estimator.
+According to Ferman & Pinto (2021), this bias arises from endogeneity in the linear regression model implied by the SCM assumption due to the correlation between the outcomes of untreated units and the error term.
 This issue is analogous to the endogeneity caused by measurement error (Greene, 2003).
@@
 To address this issue, we propose an alternative approach based on a mixture model assumption rather than a linear relationship among expected outcomes.
-Specifically, we assume that the density of the treated unit’s outcome can be approximated by a weighted combination of the densities of the untreated units.
+Specifically, we assume that the distribution of the treated unit’s outcome can be approximated by a weighted combination of the distributions of the untreated units.
 This assumption further implies that each moment of the treated unit’s outcome can be represented as a weighted sum of the moments of the untreated units.
 Leveraging this insight, we estimate SC weights by matching the moments of the treated unit’s outcome to those of the untreated units.
-We refer to our method as the Density Matching SCM (DMSCM).
+We refer to our method as the Moment Matching SCM (MMSCM).
 Unlike existing methods, our proposed estimator is asymptotically unbiased if the mixture model assumption is correct.
@@
 This discrepancy arises because, under the mixture model assumption, Gunsilius (2023) provides a consistent estimator for the treatment effect, whereas existing SCM methods do not.
-We establish a connection between distributional SCMs and the asymptotic bias problem and propose a simplified estimation method based on density matching.
+We establish a connection between distributional SCMs and the asymptotic bias problem and propose a simplified estimation method based on moment matching.
 In summary, our contributions are as follows:
   • Reexamining the assumptions necessary for treatment effect estimation in SCMs.
-  • Proposing a density-matching approach for asymptotically unbiased SCMs.
-  • Simplifying the implementation of density matching in SCMs.
+  • Proposing a moment-matching approach for asymptotically unbiased SCMs.
+  • Simplifying the implementation of moment matching in SCMs.
 We formulate the problem in Section 2.
@@
 In Section 5, we then propose our method under the mixture model assumption.
-We investigate the asymptotic properties of this method in Section 6 and present a statistical inference method in Section 7.
+We investigate the asymptotic properties of this method in Section 5.5 and present a statistical inference method in Section 5.6.
 In Section 6, we discuss related topics and existing studies.
@@
 Related work.
-Following the pioneering work by Abadie and Gardeazabal (2003), numerous studies have advanced the use of SCMs in comparative studies (Abadie et al., 2010, 2015).
-These SCMs, including their variants such as those by Doudchenko and Imbens (2016) and Ben-Michael et al.
+Following the pioneering work by Abadie & Gardeazabal (2003), numerous studies have advanced the use of SCMs in comparative studies (Abadie et al., 2010, 2015).
+These SCMs, including their variants such as those by Doudchenko & Imbens (2016) and Ben-Michael et al.
 (2019), are based on minimizing the pre-treatment sum of squared errors between the outcome of the treated unit and the weighted sum of those of the untreated units, with some regularization and constraints regarding the weights.
@@
 (2010) demonstrates that under a pre-treatment fit, the bias of the LS-SCM estimator converges to zero asymptotically as the length of the pre-treatment periods $T_{0}$ increases.
-However, Ferman and Pinto (2021) finds that if the pre-treatment fit does not hold, then the LS-SCM estimator has a bias and is not consistent.
+However, Ferman & Pinto (2021) finds that if the pre-treatment fit does not hold, then the LS-SCM estimator has a bias and is not consistent.
 The cause of this bias is the correlation between the outcomes of the untreated units and the error term in the posited linear model, which we refer to as implicit endogeneity.
@@
 Ferman (2021) demonstrates that we can obtain an asymptotically unbiased estimator under the regime where both the number of untreated units ($J\to\infty$) and the length of the pre-intervention periods ($T_{0}$) grow large.
-Concurrently with us, Fry (2024) also discusses the endogeneity problem in SCMs as a cause of the asymptotic bias reported by Ferman and Pinto (2021).
+Concurrently with us, Fry (2024) also discusses the endogeneity problem in SCMs as a cause of the asymptotic bias reported by Ferman & Pinto (2021).
 [^1] Note that his study became public on arXiv after our study became public on arXiv and at the ICML workshop, and he and we conducted our studies completely independently.
 While our formulation and his regarding endogeneity are similar, our approach differs significantly.
-We address the problem through density matching, similar to the strategy in Gunsilius (2023), whereas Fry (2024) employs an instrumental variable approach using outcomes of untreated units not used in SCMs as instruments.
+We address the problem through moment matching, similar to the strategy in Gunsilius (2023), whereas Fry (2024) employs an instrumental variable approach using outcomes of untreated units not used in SCMs as instruments.
 In this study, we focus on mixture models among outcome distributions, an assumption also employed in Wan et al.
@@
 While Gunsilius (2023) suggests that standard least squares SCMs try to match the first moment of the outcome distribution (and therefore differ from distributional approaches), we stress that least squares approaches are additionally vulnerable to implicit endogeneity, which can generate bias.
-Our density-matching approach avoids this problem by matching a broader set of distributional features.
+Our moment-matching approach avoids this problem by matching a broader set of distributional features.
 Table 1: Comparison of SCMs. See the following sections for the definitions of variables and methods.
 Method Assumption Asymptotic unbiasedness Remark
-Ours DMSCM $F_{Y^{N}_{0,t}}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F_{Y^{N}_{j,t}}(y)\ \ \ \forall y\in\mathbb{R}$ Unbiased
-Abadie and Gardeazabal (2003) Constrained least squares $\mathbb{E}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}[Y^{N}_{j,t}]$ Biased
+Ours MMSCM $F_{Y^{N}_{0,t}}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F_{Y^{N}_{j,t}}(y)\ \ \ \forall y\in\mathbb{R}$ Unbiased
+Abadie & Gardeazabal (2003) Constrained least squares $\mathbb{E}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}[Y^{N}_{j,t}]$ Biased
 Ferman (2021) Constrained least squares $\mathbb{E}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}[Y^{N}_{j,t}]$ Unbiased $J\to\infty$
@@
 2 Problem Setting
-Suppose there are $J+1$ units, indexed by $j\in\mathcal{J}\coloneqq\{0,1,2,\dots,J\}$, and a time series indexed by $t\in\mathcal{T}=\{1,2,\dots,T\}$, where $3\leq T\in\mathbb{N}$.
-Without loss of generality, assume that the first unit ($j=0$) is the treated unit, i.e., the unit affected by the policy intervention of interest.
-The set of potential control units is $\mathcal{J}^{U}\coloneqq\mathcal{J}\backslash\{0\}=\{1,2,\dots,J\}$, consisting of units not affected by the intervention.
-We also assume that our data span $T$ periods and that the intervention occurs at $t=T_{0}$, with $2\leq T_{0}<T$.
-Observations in period $t\in\mathcal{T}_{0}\coloneqq\{1,2,\dots,T_{0}\}$ are from before the intervention, and observations in period $t\in\mathcal{T}_{1}\coloneqq\mathcal{T}\backslash\mathcal{T}_{0}$ are from after the intervention.
-Let $T_{1}\coloneqq|\mathcal{T}_{1}|=T-T_{0}$.
+[…]
 2.1 Potential Outcomes
@@
 Depending on the application, we may also adjust the target density, for example, by averaging across all post-intervention periods: $\overline{p}^{N}_{0}(y)=\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p^{N}_{0,t}(y)$.
-We refer to the treatment effects related to these distributions as distributional treatment effects (Park et al., 2021; Kallus and Oprescu, 2022; Chikahara et al., 2022).
+We refer to the treatment effects related to these distributions as distributional treatment effects (Park et al., 2021; Kallus & Oprescu, 2022; Chikahara et al., 2022).
 In the context of SCMs, Gunsilius (2023) and Chen (2020) have proposed methods for estimating such distributional treatment effects.
 3 Recap of SCMs
-This section reviews the original SCM proposed by Abadie and Gardeazabal (2003).
+This section reviews the original SCM proposed by Abadie & Gardeazabal (2003).
 To estimate the treatment effect $\tau_{0,t}$, we predict the counterfactual outcome $Y^{N}_{0,t}$ using a weighted sum of $\{Y^{N}_{j,t}\}_{j=1}^{J}$ for $t\in\mathcal{T}_{1}$.
@@
 Several methods exist for estimating the SC weights.
-In the SCM proposed by Abadie and Gardeazabal (2003), these weights are determined by minimizing the squared error between $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$.
+In the SCM proposed by Abadie & Gardeazabal (2003), these weights are determined by minimizing the squared error between $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$.
 Define
 $\Delta^{J}\coloneqq\left\{\bm{w}=(w_{1}\ w_{2}\ \dots\ w_{J})^{\top}\in[0,1]^{J}\mid\sum_{j\in\mathcal{J}^{U}}w_{j}=1\right\}.$
-Then, Abadie and Gardeazabal (2003) estimate the SC weights as
+Then, Abadie & Gardeazabal (2003) estimate the SC weights as
 $\bm{w}^{\mathrm{Abadie}}=\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{0,t}-\sum_{j\in\mathcal{J}^{U}}w_{j}Y^{N}_{j,t}\right)^{2}.$
@@
 $\widehat{\tau}^{\mathrm{Abadie}}_{t}=Y^{I}_{0,t}-\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}.$
-The literature on inference for SCMs has grown significantly, extending the original proposals by Abadie and Gardeazabal (2003) and Abadie et al.
+The literature on inference for SCMs has grown significantly, extending the original proposals by Abadie & Gardeazabal (2003) and Abadie et al.
 (2010).
-Notable contributions include works by Li (2020), Shaikh and Toulis (2021), Cattaneo et al.
+Notable contributions include works by Li (2020), Shaikh & Toulis (2021), Cattaneo et al.
 (2021), and Chernozhukov et al.
 (2021).
-Doudchenko and Imbens (2016) proposes estimating the SC weight as
+Doudchenko & Imbens (2016) proposes estimating the SC weight as
 $\widetilde{\bm{w}}^{\mathrm{DI}}=\operatorname*{arg\,min}_{\widetilde{\bm{w}}=(w_{0}\ w_{1}\ \cdots\ w_{J})^{\top}\in\mathbb{R}^{J+1}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{0,t}-w_{0}-\sum_{j\in\mathcal{J}^{U}}w_{j}Y_{j,t}\right)^{2}+\lambda_{1}\|\widetilde{\bm{w}}\|_{1}+\lambda_{2}\|\widetilde{\bm{w}}\|_{2},$
@@
 $Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}.$
-This assumption, often called the perfect pre-treatment fit (Ferman and Pinto, 2021), implies that the stochastic process $\left(Y^{N}_{0,t}\right)_{t\in\mathcal{T}_{0}}$ can be fully replicated by $\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$ for $t\in\mathcal{T}_{0}$.
+This assumption, often called the perfect pre-treatment fit (Ferman & Pinto, 2021), implies that the stochastic process $\left(Y^{N}_{0,t}\right)_{t\in\mathcal{T}_{0}}$ can be fully replicated by $\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$ for $t\in\mathcal{T}_{0}$.
 In linear factor models where $Y^{N}_{j,t}=c_{j}+\delta_{t}+\lambda_{t}\mu_{j}+\epsilon_{j,t}$, this assumption might imply $\epsilon_{0,t}=\sum_{j\in[J]}w^{*}_{j}\epsilon_{j,t}$, which is often considered unrealistic.
@@
 Assumption 4.2 is a specific case of Assumption 4.1; if Assumption 4.1 holds, then Assumption 4.2 also holds under the linear factor model in Definition 2.1.
-Under these assumptions, Ferman and Pinto (2021) shows that the SC estimator $\widehat{\tau}^{\mathrm{Abadie}}_{0,t}$ suffers from bias and is not consistent for $\tau_{0,t}$ when $c_{j}$ is non-zero.
+Under these assumptions, Ferman & Pinto (2021) shows that the SC estimator $\widehat{\tau}^{\mathrm{Abadie}}_{0,t}$ suffers from bias and is not consistent for $\tau_{0,t}$ when $c_{j}$ is non-zero.
 We also adopt the following assumption regarding the linear factor model.
@@
   • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\epsilon_{j,t}\lambda_{t}\xrightarrow{\mathrm{p}}0$.
-Under Assumptions 4.2 and 4.3, Ferman and Pinto (2021) concludes that the LS-SCM proposed by Abadie and Gardeazabal (2003) does not yield a consistent estimator for $\tau_{0,t}$, as the following proposition.
-Proposition 4.4 (Proposition 1 in Ferman and Pinto (2021)).
+Under Assumptions 4.2 and 4.3, Ferman & Pinto (2021) concludes that the LS-SCM proposed by Abadie & Gardeazabal (2003) does not yield a consistent estimator for $\tau_{0,t}$, as the following proposition.
+Proposition 4.4 (Proposition 1 in Ferman & Pinto (2021)).
 Under Assumptions 4.34.2, we have
@@
 as $T_{0}\to\infty$.
-To alleviate this bias problem, Ferman and Pinto (2021) proposes the demeaned SC estimator:
+To alleviate this bias problem, Ferman & Pinto (2021) proposes the demeaned SC estimator:
 $\widehat{\bm{w}}^{\mathrm{FP}}=\operatorname*{arg\,min}_{\widehat{\bm{w}}^{\mathrm{FP}}\in\Delta^{J}}\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left\{\left(Y^{N}_{0,t}-\overline{Y}^{N}_{0}\right)-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)\right\}^{2},$
@@
 4.3 Implicit Endogeneity
-We now interpret the asymptotic bias found by Ferman and Pinto (2021) as an endogeneity problem.
+We now interpret the asymptotic bias found by Ferman & Pinto (2021) as an endogeneity problem.
 We consider (2), that is, $\mathbb{E}_{0,t}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}[Y^{N}_{j,t}]$, which is more general than the linear factor model.
@@
 Thus, the identification of treatment effects faces a challenge due to implicit endogeneity.
-While the estimator $\bm{w}^{\mathrm{Abadie}}$ remains standard in many applications, it inherits this endogeneity issue.
-Such concerns have been noted in prior works (Chernozhukov et al., 2019; Ferman and Pinto, 2021; Ferman, 2021), and Fry (2024) independently frames the bias as an endogeneity problem, proposing an instrumental variable method to address it.
-5 SCM by Density Matching
+While the estimator $\widehat{\bm{w}}^{\mathrm{Abadie}}_{n}$ remains standard in many applications, it inherits this endogeneity issue.
+Such concerns have been noted in prior works (Chernozhukov et al., 2021; Ferman & Pinto, 2021; Ferman, 2021), and Fry (2024) independently frames the bias as an endogeneity problem, proposing an instrumental variable method to address it.
+Bias is caused by a measurement error.
+As we showed above, the source of the implicit endogeneity is the measurement error.
+Since we assume $\mathbb{E}_{0,t}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}[Y^{N}_{j,t}]$ as a regression model, the regressors are $(\mathbb{E}_{j,t}[Y^{N}_{j,t}])_{j\in\mathcal{J}^{U}}$.
+However, in practice, we can only use $(Y_{j,t})$ as the regressors.
+Because the regressors contain measurement errors $(\mathbb{E}_{j,t}[Y^{N}_{j,t}]-Y_{j,t})_{j\in\mathcal{J}^{U}}$, endogeneity arises, as is commonly discussed in elementary textbooks (Greene, 2003).
+5 SCM by Moment Matching
 As discussed in Section 4, Assumption 4.1 can be weak or incompatible for identifying $\bm{w}^{*}$ when using least squares.
 We, therefore, adopt a different assumption for identification.
-Specifically, we propose that the density of $Y^{N}_{0,t}$ can be replicated by a weighted combination of the densities of $Y^{N}_{j,t}$, and we introduce an SCM based on density matching.
+Specifically, we propose that the distribution of $Y^{N}_{0,t}$ can be replicated by a weighted combination of the distributions of $Y^{N}_{j,t}$, and we introduce an SCM based on moment matching.
 5.1 Identification and Estimation Strategy
-Section 4 and the results in Ferman and Pinto (2021) indicate that combining
+Section 4 and the results in Ferman & Pinto (2021) indicate that combining
 $\mathbb{E}_{0,t}\left[Y_{0,t}\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[Y_{j,t}\right]$
 with least squares estimators do not yield consistent estimators of $\bm{w}^{*}$. We thus adopt mixture models as a more theoretically favorable but still realistic assumption.
+Identification.
 We begin by positing that the distribution of the target unit’s outcome $Y^{N}_{0,t}$ can be expressed as a linear combination of the distributions of $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$.
@@
 $p^{N}_{0,t}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p^{N}_{j,t}(y)\qquad\forall t\in\mathcal{T}.$ (3)
-We next discuss estimating SC weights that satisfy (3).
+Estimation.
+We next discuss how to estimate the SC weights that satisfy (3).
 One approach employs optimal transport to minimize the Wasserstein distance (Gunsilius, 2023).
-In this study, we propose a simpler estimation procedure, as the approach of Gunsilius (2023) involves more complexity than the original SCMs.
-Because the mixture model (3) holds, we obtain:
-$\mathbb{E}_{0,t}\left[\exp\left(\sqrt{-1}\xi Y^{N}_{0,t}\right)\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[\exp\left(\sqrt{-1}\xi Y^{N}_{j,t}\right)\right]\quad\forall t\in\mathcal{T},\ \forall\xi\in\mathbb{R},$
+In this study, we propose a simpler estimation procedure, as the approach of Gunsilius (2023) is more complex than that of the original SCMs.
+There are various ways to estimate $(w^{*}_{j})_{j\in\mathcal{J}^{U}}$. We consider the following:
+  • Distribution divergence (distance) minimization.
+  • Characteristic function matching.
+  • Moment matching.
+A straightforward approach to estimating $(w^{*}_{j})_{j\in\mathcal{J}^{U}}$ is to minimize the divergence or distance between two distributions whose densities are given by $p^{N}_{0,t}(y)$ and $\sum_{j\in\mathcal{J}^{U}}w_{j}p^{N}_{j,t}(y)$, given $(w_{j})_{j\in\mathcal{J}^{U}}$.
+For the divergence or distance measures, we may use, for example, the Kullback-Leibler (KL) divergence or the Wasserstein distance.
+The KL divergence can be estimated nonparametrically using density ratio techniques (assuming the densities exist) (Sugiyama et al., 2012; Kato & Teshima, 2021).
+Alternatively, we may estimate the weights by minimizing the maximum mean discrepancy.
+The second approach is to match characteristic functions.
+Since the mixture model (3) holds, we have $\mathbb{E}_{0,t}\left[\exp\left(\sqrt{-1}\xi Y^{N}_{0,t}\right)\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[\exp\left(\sqrt{-1}\xi Y^{N}_{j,t}\right)\right]$ for all $t\in\mathcal{T}$ and $\xi\in\mathbb{R}$.
 where $\sqrt{-1}$ denotes the imaginary unit.
-From this, we can estimate $\bm{w}^{*}$ by considering the upper bound of the difference between the expected exponential transforms of $Y^{N}_{0,t}$ and the corresponding weighted sum for $Y^{N}_{j,t}$.
-Specifically, we have
+The LHS is the characteristic function (CHF) of $p^{N}_{0,t}(y)$, and the RHS is that of $\sum_{j\in\mathcal{J}^{U}}w_{j}p^{N}_{j,t}(y)$, given $(w_{j})_{j\in\mathcal{J}^{U}}$.
+Since CHFs uniquely determine the corresponding distributions, we can estimate $(w^{*}_{j})_{j\in\mathcal{J}^{U}}$ by matching the CHFs.
+Such matching can be implemented using the method of moments, for instance, the approach proposed in Carrasco & Kotchoni (2017).
+The last approach is to match the moments of the distributions $p^{N}_{0,t}(y)$ and $\sum_{j\in\mathcal{J}^{U}}w_{j}p^{N}_{j,t}(y)$.
+Intuitively, if we find $(w_{j})_{j\in\mathcal{J}^{U}}$ such that all the moments of the distributions match, we can interpret it as an estimate of $(w^{*}_{j})_{j\in\mathcal{J}^{U}}$.
+This idea can also be motivated from the perspective of CHF matching, as noted in Remark 5.2.
+Note that to justify this approach, we must ensure that the moments uniquely determine the distributions.
+This is the classical moment problem (Schmüdgen, 2020), which has been extensively studied.
+It is known that certain conditions are required to guarantee uniqueness.
+Among the three approaches, we focus on the last one, moment matching.
+As we show later, it is sufficient to estimate $(w^{*}_{j})_{j\in\mathcal{J}^{U}}$ by matching sample means of polynomials of $Y_{a,j}$.
+Compared to the other approaches, the implementation of moment matching is significantly simpler.
+Note that it cannot be applied to all general distributions without additional assumptions since the moments may not be uniquely determined for each distribution.
+However, if we assume boundedness of $Y_{a,j}$, we can ensure that the distributions are uniquely determined by their moments (Hausdorff, 1921).
+In this study, for simplicity, we adopt this moment matching approach under the assumption of bounded random variables.
+Remark 5.1 (Distribution SCMs using divergence or distance minimization.).
+Since we are interested in finding $(w_{j})_{j\in\mathcal{J}^{U}}$ such that two distributions match, it is sufficient to use a measure that guarantees the distributions are identical when the divergence or distance is zero.
+That is, we do not need to consider symmetry or deficient support, which often motivate the use of the Wasserstein distance instead of the KL divergence.
+Therefore, while Gunsilius (2023) introduces distributional SCM using the Wasserstein distance, we believe that there is no strong advantage to using it over other divergence or distance measures.
+Remark 5.2 (From CHF matching to moment matching).
+We can estimate $\bm{w}^{*}$ by bounding the difference between the CHFs of $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{j,t}$. Specifically, we have
 $\int_{\xi\in(-h,h)}\left|\mathbb{E}_{0,t}\left[\exp(\sqrt{-1}\xi Y^{N}_{0,t})\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\exp(\sqrt{-1}\xi Y^{N}_{j,t})\right]\right|\mathrm{d}\xi$
-$=\int_{\xi\in(-h,h)}\Biggl{|}\sum_{\gamma\in\{0,1,2,\dots,G\}}\frac{\sqrt{-1}^{\gamma}\xi^{\gamma}}{\gamma!}\left(\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right)+\epsilon\left|\xi^{G}\right|\Biggr{|}\mathrm{d}\xi$
+$=\int_{\xi\in(-h,h)}\Biggl|\sum_{\gamma\in\{0,1,2,\dots,G\}}\frac{\sqrt{-1}^{\gamma}\xi^{\gamma}}{\gamma!}\left(\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right)+\epsilon|\xi^{G}|\Biggr|\mathrm{d}\xi$
 $\leq\int_{\xi\in(-h,h)}\Biggl(\sum_{\gamma\in\{1,2,\dots,G\}}\frac{|\sqrt{-1}^{\gamma}||\xi|^{\gamma}}{\gamma!}\left|\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right|+\epsilon|\xi^{G}|\Biggr)\mathrm{d}\xi$
@@
 $\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|+\frac{2\epsilon}{G+1}h^{G+1}$
-thus yields an accurate estimate of $\bm{w}^{*}$ that enforces
+thus yields an accurate estimate of $\bm{w}^{*}$ that ensures
 $\int_{\xi\in(-h,h)}\left|\mathbb{E}_{0,t}\left[\exp(\sqrt{-1}\xi Y^{N}_{0,t})\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\exp(\sqrt{-1}\xi Y^{N}_{j,t})\right]\right|\mathrm{d}\xi=0.$
-This approach can be seen as a variant of the method of moments, where we match all moments of $Y^{N}_{0,t}$ with those of the weighted sum of the untreated units.
+This approach can be viewed as a variant of the method of moments, where we match all moments of $Y^{N}_{0,t}$ with those of the weighted sum of the untreated units.
+Remark 5.3 (Reports in Gunsilius (2023)).
+Distributional SCMs are also considered in Gunsilius (2023).
+However, the author does not emphasize the fact that distributional SCMs can estimate the treatment effect in an asymptotically unbiased manner, whereas original SCMs such as Abadie & Gardeazabal (2003) yield biased estimators.
+For example, Gunsilius (2023) states that “the distributional synthetic controls method finds optimal weights that replicate all moments of the target distribution as closely as possible.
+In contrast, applying the classical method to averages of the distributions will find the optimal weights based on replicating the first moment.” This claim may be accurate, as the classical method cannot yield unbiased and consistent estimates of the optimal weights for replicating the first moment unless, for example, we assume that the outcomes are nonrandom variables.
 5.2 Core Assumptions
@@
 Assumption 5.4.
-The outcomes $(Y^{N}_{j,t})_{j\in\mathcal{J},t\in\mathcal{T}}$ are bounded.
+The outcomes $(Y^{N}_{j,t})_{j\in\mathcal{J},t\in\mathcal{T}}$ are bounded as $Y^{N}_{j,t}\in[0,1]$.
 Assumption 5.5.
@@
 Assumption 5.4 ensures that $Y^{N}_{j,t}$ has a finite $G$-th moment for any $G\in\mathbb{N}$.
+If $Y^{N}_{j,t}$ is bounded but not in $[0,1]$, we can normalize $Y^{N}_{j,t}$ to satisfy this assumption without loss of generality.
 Assumption 5.5 has been employed in earlier work, including Wan et al.
@@
 5.3 Estimation of SC Weights
-Under the mixture model assumption, we propose estimating SC weights by matching characteristic functions (CHFs). We call our approach the Density Matching SCM (DMSCM).
-Let $h\in(0,1)$ be a fixed parameter. We then estimate the SC weights as
-$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\coloneqq\left(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0},G}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|.$ (4)
-Using $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}$, we can then estimate both the treatment effect and the distributional treatment effect as follows.
-5.3.1 Treatment Effect Estimation
-Once $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}$ is obtained, we estimate the treatment effect similarly to standard SCMs:
-$\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0},G}Y_{j,t}.$ (5)
-We call this estimator the Density Matching SCM (DMSCM) estimator. Algorithm 1 summarizes our proposed estimator and also includes the D2MSCM estimator introduced later.
-5.3.2 Distributional Treatment Effect Estimation
-To estimate the distributional treatment effect, we can approximate the counterfactual distribution $p_{0,t}(y)$ via bootstrap.
-Specifically, we resample $L>1$ observations $\{Y^{l,\dagger}_{0}\}_{l=1}^{L}$ from $\{Y_{0,t},Y_{1,t},\dots,Y_{J,t}\}_{t\in\mathcal{T}_{1}}$ as follows: for each $l\in\mathcal{L}\coloneqq\{1,2,\dots,L\}$,
-  1. 1.
-    For each $j\in\mathcal{J}$, resample $Y^{l,\dagger}_{j}$ from $\{Y_{j,t}\}_{t\in\mathcal{T}_{1}}$.
-  2. 2.
-    Choose $Y^{l,\dagger}_{0}$ from $\left(Y^{l,\dagger}_{1},\dots,Y^{l,\dagger}_{J}\right)$ with probability $\left(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0}}\right)_{j\in\mathcal{J}^{U}}$.
-Because our primary focus is often on $\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p_{0,t}(y)$, this bootstrap procedure enables us to approximate the distribution of the counterfactual outcomes.
-Algorithm 1 DMSCM.
+Under the mixture model assumption, we propose estimating SC weights by matching moment functions. We refer to our approach as the Momemnt Matching SCM (MMSCM).
+Let $v_{\gamma}\in(0,\infty)$ for $\gamma=1,2,\dots,G$ be a fixed weight sequence. We then estimate the SC weights as
+$\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0},G}\coloneqq\left(\widehat{w}^{\mathrm{MMSCM}}_{j,T_{0},G}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma=1}^{G}v_{\gamma}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|.$ (4)
+We may use any weights $v_{\gamma}\in(0,\infty)$ for $\gamma=1,2,\dots,G$.
+For example, based on Remark 5.2, we can choose $v_{\gamma}=\frac{2h^{\gamma+1}}{(\gamma+1)!}$.
+In this case, we estimate $(w^{*}_{j})_{j\in\mathcal{J}^{U}}$ as
+$\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0},G}\coloneqq\left(\widehat{w}^{\mathrm{MMSCM}}_{j,T_{0},G}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|.$
+Using $\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0},G}$, we can then estimate both the treatment effect and the distributional treatment effect, as discussed in the following subsections.
+5.4 Treatment Effect Estimation
+Once $\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0},G}$ is obtained, we estimate the treatment effect similarly to standard SCMs:
+$\widehat{\tau}^{\mathrm{MMSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{MMSCM}}_{j,T_{0},G}Y_{j,t}.$ (5)
+We call this estimator the Moment Matching SCM (MMSCM) estimator. Algorithm 1 summarizes our proposed estimator.
+Remark 5.6 (Multiple treatments).
+We can extend our method to the case with multiple treatments $K\geq 2$.
+Assume that there are units $\widetilde{J},\cdots,-1$ that receive treatments $\widetilde{J},\cdots,-1$, while unit 0 receives treatment $1$.
+In such cases, how we assume the DGP of the outcomes is crucial.
+For example, if we assume that the outcomes of the treated units follow the same distribution, it is sufficient to take the average over the treated group; that is, we solve $\min_{\bm{w}\in\Delta^{J}}\sum_{\gamma=1}^{G}v_{\gamma}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\frac{1}{\widetilde{J}+1}\sum_{\widetilde{j}\in\{0,-1,\dots,\widetilde{J}\}}\left(Y^{N}_{\widetilde{j},t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|$.
+We can also consider other situations in this setting, such as differential treatment timing.
+These extensions remain an open issue in the general context of SCM, and are not limited to the method proposed in this study.
+Algorithm 1 MMSCM.
 Parameter: The number of moments $G$.
-  Obtain trained weights $\widehat{\bm{w}}^{\mathrm{DMSCM}}$ as (4).
+ Obtain trained weights $\widehat{\bm{w}}^{\mathrm{MMSCM}}$ as (4).
  Estimate the treatment effect as (5).
-6 Convergence Analysis
+5.5 Convergence Analysis
 We now provide consistency results for our proposed SC weight estimator. The following theorem presents these results, including consistency for the treatment effect $\tau_{0}$.
 Assumption 5.7.
-Assume that the density function of $\varepsilon_{j,t}$ exists and is bounded for each $j\in\mathcal{J}$ and $t\in\mathcal{T}_{0}$, and that the stochastic process $\{\varepsilon_{j,t}\}_{t=1}^{T_{0}}$ is stationary, strongly mixing, with the sum of mixing coefficients bounded by $M$.
+The stochastic process $\{\varepsilon_{j,t}\}_{t=1}^{T_{0}}$ is stationary, strongly mixing, with the sum of mixing coefficients bounded by $M$.
 Theorem 5.8.
-Suppose that Assumptions 5.15.2 and 6.1 hold. Fix $h\in(0,1)$. Define
-$\widetilde{\Phi}^{\dagger}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right|.$
+Suppose that Assumptions 5.45.5 and 5.7 hold. Let $v_{\gamma}\in(0,\infty)$ for $\gamma=1,2,\dots,G$ be a fixed weight sequence. Define
+$\widetilde{\Phi}^{\dagger}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}v_{\gamma}\left|\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right|.$
 Then, it holds that
-$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}\quad\text{as }T_{0}\to\infty.$
+$\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}\quad\text{as }T_{0}\to\infty.$
 Additionally, $\bm{w}^{*}_{G}\to\bm{w}^{*}\in\widetilde{\Phi}^{*}$ as $G\to\infty$, and
-$\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}\xrightarrow{\mathrm{p}}\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)\quad\text{as }T_{0}\to\infty\text{ and }G\to\infty.$
-The proof is shown in Appendix A.
-Note that this result also applies to $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ as a special case.
-Consequently, under the respective assumptions, $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ and $\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}$ are asymptotically unbiased.
-7 Inference
-This section presents methods for statistical inference using our proposed DMSCM estimator.
-7.1 Conformal Inference on the Treatment Effect
-We consider statistical inference for the treatment effect using our proposed DMSCM and D2SCM, focusing on testing a sharp null hypothesis, $H_{0}:\tau_{0,t}=\alpha_{t}$ for some $t\in\mathcal{T}_{t}$ and $\bm{\alpha}=(\alpha_{t})_{t\in\mathcal{T}_{1}}$.
+$\widehat{\tau}^{\mathrm{MMSCM}}_{0,t}\xrightarrow{\mathrm{p}}\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)\quad\text{as }T_{0}\to\infty\text{ and }G\to\infty.$
+The proof is shown in Appendix A. Note that this result also applies to $\widehat{\tau}^{\mathrm{MMSCM}}_{0,t}$ as a special case.
+Consequently, because $\mathbb{E}\left[\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right]=0$, $\widehat{\tau}^{\mathrm{MMSCM}}_{0,t}$ is asymptotically unbiased.
+Relationship to the classical SCMs.
+We next consider the relationship between our MMSCM and the classical SCMs proposed in Abadie & Gardeazabal (2003).
+Abadie & Gardeazabal (2003) assumes that $Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$, which corresponds to a specific case of our mixture model assumption, $F_{Y^{N}_{0,t}}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F_{Y^{N}_{j,t}}(y)\quad\forall y\in\mathbb{R}$.
+Therefore, under the assumptions in Abadie & Gardeazabal (2003), by using the MMSCMs, we can estimate the weights in an asymptotically unbiased way.
+Corollary 5.9.
+Suppose that Assumptions 5.4 and 5.7 hold. Assume that
+$Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$
+holds. Let $v_{\gamma}\in(0,\infty)$ for $\gamma=1,2,\dots,G$ be a fixed weight sequence. Define
+$\widetilde{\Phi}^{\dagger}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}v_{\gamma}\left|\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right|.$
+Then, it holds that
+$\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}\quad\text{as }T_{0}\to\infty.$
+Additionally, $\bm{w}^{*}_{G}\to\bm{w}^{*}\in\widetilde{\Phi}^{*}$ as $G\to\infty$, and
+$\widehat{\tau}^{\mathrm{MMSCM}}_{0,t}\xrightarrow{\mathrm{p}}\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)\quad\text{as }T_{0}\to\infty\text{ and }G\to\infty.$
+Remark 5.10 (Section 4.1 in Gunsilius (2023)).
+The proposed method is an extension of the classical method in a rigorous sense: if we apply it to probability measures supported on one point, that is, Dirac measures of the form $\delta_{y}(A)$, taking the value $1$ if $y\in A$ and 0 otherwise, then we obtain the same results as the classical method.
+This means the proposed method reduces to the classical estimator when we are only given aggregate values and not distributions.
+5.6 Inference
+This section presents methods for statistical inference using our proposed MMSCM estimator.
+We consider statistical inference for the treatment effect using our proposed MMSCM and D2SCM, focusing on testing a sharp null hypothesis, $H_{0}:\tau_{0,t}=\alpha_{t}$ for some $t\in\mathcal{T}_{t}$ and $\bm{\alpha}=(\alpha_{t})_{t\in\mathcal{T}_{1}}$.
 Specifically, we adopt the conformal inference approach of Chernozhukov et al.
 (2021), which has been used by, for example, Ben-Michael et al.
-(2021) and Ferman and Pinto (2021).
+(2021) and Ferman & Pinto (2021).
+Note that under our linear factor model, $\tau_{0,t}=Y^{N}_{0,t}-Y^{I}_{0,t}=\alpha_{t}$ holds, that is, the ATE equals the individual treatment effect.
+Therefore, our null is sharp.
+This formulation is also adopted in other works, such as Chernozhukov et al.
+(2021).
 Our goal is to construct a confidence interval whose $p$-value $\widehat{p}$ satisfies
@@
 $\left(Y_{0,1},\dots,Y_{0,T_{0}},Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right).$
-We then apply our proposed method to this augmented dataset (comprising
-$\left(Y_{0,1},\dots,Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$
-and $(Y_{j,1})_{t\in\mathcal{T},j\in\mathcal{J}^{U}}$) to obtain adjusted weights $\widehat{\bm{w}}(\alpha)$ and an adjusted predictor of the counterfactual outcome, denoted $\widehat{Y}^{N}_{0,t}(\bm{\alpha})$.
+We then apply our proposed method to this augmented dataset, comprising $\left(Y_{0,1},\dots,Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$ and $(Y_{j,1})_{t\in\mathcal{T},j\in\mathcal{J}^{U}}$, to obtain adjusted weights $\widehat{\bm{w}}(\alpha)$ and an adjusted predictor of the counterfactual outcome, denoted $\widehat{Y}^{N}_{0,t}(\bm{\alpha})$.
 Finally, we compute the $p$-value by assessing whether the adjusted residual “conforms” with the pre-treatment residuals.
@@
 $\widehat{C}^{\mathrm{conf}}_{1-\xi}=\left\{\bm{\alpha}\in\mathcal{A}\bigm|\widehat{p}(\bm{\alpha})>\xi\right\},$
-where $\mathcal{A}$ is a set of candidate values. Theorem 1 of Chernozhukov et al. (2021) guarantees (6).
-7.2 Two-Sample Homogeneity Test
+where $\mathcal{A}$ is a set of candidate values. Theorem 1 of Chernozhukov et al. (2021) guarantees (6). We summarize the procedure in Algorithm 2
+Remark 5.11 (Conformal inference and randomization inference).
+Conformal inference provides confidence intervals under a stationarity assumption and a prediction model for the counterfactual outcome.
+In contrast, randomization inference requires random assignment and prediction models for post treatment outcomes of all units.
+These differences are recognized advantages of conformal inference (Chernozhukov et al., 2021), and the method has recently been widely used in SCMs (Ben-Michael et al., 2021).
+Furthermore, conformal inference can itself be interpreted as a variant of randomization inference (Ritzwoller et al., 2025).
+Algorithm 2 Conformal prediction.
+Sharp null hypothesis: $H_{0}:\tau_{0,t}=\alpha_{t}$ for some $t\in\mathcal{T}_{t}$ and $\bm{\alpha}=(\alpha_{t})_{t\in\mathcal{T}_{1}}$.
+ Make an augmented dataset $\left(Y_{0,1},\dots,Y_{0,T_{0}},Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$.
+ We apply the MMSCM by regarding the datasets $\left(Y_{0,1},\dots,Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$ and $(Y_{j,1})_{t\in\mathcal{T},j\in\mathcal{J}^{U}}$ as the pretreatment dataset.
+ Using the estimated weight, we obtain estimate of the counterfactual outcome $\widehat{Y}^{N}_{0,t}(\bm{\alpha})$.
+ Obtain $u_{t}(\bm{\alpha})=Y_{0,t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{0}$, and $u_{t}(\bm{\alpha})=Y^{I}_{0,t}-\alpha_{t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{1}$.
+ Compute the $p$-value as $\widehat{p}(\bm{\alpha})=1-\widehat{F}(\bm{\alpha})$ and confidence interval as $\widehat{C}^{\mathrm{conf}}_{1-\xi}=\left\{\bm{\alpha}\in\mathcal{A}\bigm|\widehat{p}(\bm{\alpha})>\xi\right\}$.
+5.7 Distributional Treatment Effect Estimation
+To estimate the distributional treatment effect, we approximate the counterfactual distribution $p_{0,t}(y)$ using a bootstrap-like simulation-based method.
+Specifically, we resample $L>1$ observations $\{Y^{l,\dagger}_{0}\}_{l=1}^{L}$ from $\{Y_{0,t},Y_{1,t},\dots,Y_{J,t}\}_{t\in\mathcal{T}_{1}}$ as follows: for each $l\in\mathcal{L}\coloneqq\{1,2,\dots,L\}$,
+  1. 1.
+    For each $j\in\mathcal{J}$, resample $Y^{l,\dagger}_{j}$ from $\{Y_{j,t}\}_{t\in\mathcal{T}_{1}}$.
+  2. 2.
+[…]
+The empirical distribution of $\{Y^{l,\dagger}_{0}\}_{l=1}^{L}$ serves as an estimator of the distribution of $\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}Y_{0,t}$, as shown in the following theorem.
+The proof is provided in Appendix B.
+Thus, when our primary focus is on $\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p_{0,t}(y)$, the above procedure enables us to achieve this goal.
+Theorem 5.12.
+Suppose that the assumptions in Theorem 5.8 hold. Define
+  • $G_{T_{0},L}\coloneqq\frac{1}{L}\sum^{L}_{l=1}\mathbbm{1}[Y^{l,\dagger}_{0}\leq y]$.
+  • $F_{Y^{N}_{j,t},T_{0}}(y)\coloneqq\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\mathbbm{1}[Y^{N}_{j,t}\leq y]$.
+  • $\widehat{F}_{Y^{N}_{0,t},T_{0}}(y)\coloneqq\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{MMSCM}}_{j,T_{0},G}F_{Y^{N}_{j,t},T_{0}}(y)$.
+Then, the following holds:
+$\sup_{y}\big|G_{T_{0},L}(y)-F_{Y^{N}_{0,t}}(y)\big|\xrightarrow{\mathrm{p}}0\quad(L\to\infty,T_{0}\to\infty),$
+where we recall that $F_{Y^{N}_{0,t}}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F_{Y^{N}_{j,t}}(y)\quad\forall y\in\mathbb{R}$.
+Two-sample homogeneity test.
 To test distributional treatment effects, consider the null and alternative hypotheses $H_{0}:p^{I}_{t}=p^{N}_{t}$ versus $H_{1}:p^{I}_{t}\neq p^{N}_{t}$.
@@
 This result also implies that the mixture model (3) holds (Nazaret et al., 2023), aligning it with linear factor models.
-Hence, we obtain the following corollary for linear factor models similar to Ferman and Pinto (2021):
+Hence, we obtain the following corollary for linear factor models similar to Ferman & Pinto (2021):
 Corollary 6.6.
-Suppose that Assumptions 5.1, 6.1, and 8.18.4 hold. Then
-$\bm{w}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\quad\text{as }T_{0}\to\infty,$
-and $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ converges to
+Suppose that Assumptions 5.4, 5.7, and 6.16.4 hold. Then
+$\bm{w}^{\mathrm{MMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\quad\text{as }T_{0}\to\infty,$
+and $\widehat{\tau}^{\mathrm{MMSCM}}_{0,t}$ converges to
 $\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)$
@@
 be $K$-dimensional covariates that are unaffected by treatment. Define
-^$\widehat{\ell}_{k}(\bm{w})=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(X_{0,t,k}-\sum_{j\in\mathcal{J}^{U}}w_{j}X_{j,t,k}\right).$
-When these covariates are available, we can extend the DMSCM estimator as
-$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0}}\coloneqq v_{0}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|+\sum_{k=1}^{K}v_{k}\widehat{\ell}^{2}_{k}(\bm{w}),$
-where $\bm{v}\coloneqq(v_{k})_{k=0}^{K}$ are weights such that $v_{k}\geq 0$ and $\sum_{k=0}^{K}v_{k}=1$.
+$\widehat{\ell}_{k}(\bm{w})=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(X_{0,t,k}-\sum_{j\in\mathcal{J}^{U}}w_{j}X_{j,t,k}\right).$
+When these covariates are available, we can extend the MMSCM estimator as
+$\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}b_{0}\sum_{\gamma=1}^{G}v_{\gamma}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|+\sum_{k=1}^{K}b_{k}\widehat{\ell}^{2}_{k}(\bm{w}),$
+where $(b_{k})_{k=0}^{K}$ are weights such that $b_{k}\geq 0$ and $\sum_{k=0}^{K}b_{k}=1$.
 6.3 Distributional SCMs and Optimal Transport
@@
 where $\widehat{F}^{-1}_{Y^{N}_{j,t}}(V_{m})$ is the empirical quantile function of $Y^{N}_{j,t}$, and $\{V_{m}\}_{m=1}^{M}$ are i.i.d. draws from the uniform distribution on $[0,1]$.
-8.4 Demeaned Density Matching SCM (D2MSCM)
-Finally, we consider the case where the target distribution does not precisely satisfy (3), yet any deviation disappears after demeaning. Concretely, instead of (3), we assume
-$p^{N}_{0,t}(y-c_{0})=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p^{N}_{j,t}(y-c_{j})\qquad\forall t\in\mathcal{T}.$
-From the perspective of the linear factor model (1) and fine-grained models (Shi et al., 2022), this assumption permits $c_{j}\neq 0$ for some $j\in\mathcal{J}$.
-Under the previous mixture model, we implicitly required $c_{j}=0$ for all $j\in\mathcal{J}$, which may be restrictive in practice.
-To estimate the SC weights in this setting, we first estimate $c_{j}$ by taking the average of $(Y_{j,t})_{t\in\mathcal{T}_{0}}$ and subtracting it from each $Y_{j,t}$.
-This demeaned approach is inspired by the proposal of Ferman and Pinto (2021) but is applied here to density matching.
-Let $h\in(0,1)$ be a fixed parameter.
-We estimate the weights as
-$\widehat{\bm{w}}^{\mathrm{D2MSCM}}_{T_{0}}$ $\coloneqq\left(\widehat{w}^{\mathrm{D2MSCM}}_{j,T_{0}}\right)_{j\in\mathcal{J}^{U}}$
-$\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\biggl{|}\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}-\overline{Y}^{N}_{0}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)^{\gamma}\right)\biggr{|}.$
-We call this procedure demeaned DMSCM (D2SCM). Using these weights, we can estimate the treatment effect as
-$\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\biggl{(}\overline{Y}^{N}_{0}+\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{D2MSCM}}_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)\biggr{)}.$
 Figure 2: Results of simulation studies.
+Each result represents estimation error ($\tau_{0,t}-\widehat{\tau}$ for an estimator $\widehat{\tau}$).
+The orange horizontal line denotes the median, and the triangle mark denotes the mean.
 Figure 3: Results of simulation studies on ATE prediction. The $x$-axis is the number of untreated units $J$, and the $y$-axis is the mean squared prediction error.
@@
 7 Experiments
-We conduct experiments using both simulated and real-world data. Additional results can be found in Appendix 9.1.
+We conduct experiments using both simulated and real-world data. In all experiments, we set $v_{\gamma}=1$ for the MMSCM.
 7.1 Simulation Studies
@@
 For the treated unit, we let $Y^{N}_{0,t}\sim p_{0,t}(\bm{z})$ for $t\in[t_{0},T_{1}]$ and set $Y^{I}_{0,t}=20+Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$, implying $\tau_{0,t}=20$.
-We draw $\bm{w}^{*}$ from a uniform distribution over $[0,1]^{J}$ and then normalize it so that its entries sum to $1$.
-We apply DMSCM and the method of Abadie and Gardeazabal (2003) (referred to here as Abadie) $1000$ times and evaluate the estimation error of the treatment effect.
-We plot the estimation error box plots in Figure 3, indicate that our proposed method achieves smaller estimation errors.
+We draw $\bm{w}^{*}$ from a uniform distribution over $0,1]^{J}$ and then normalize it so that its entries sum to $1$.
+We apply MMSCM, the method proposed in Abadie et al.
+(2010) (referred to here as Abadie), and DiSCo in Gunsilius (2023).
+We iterate trials $100$ times and evaluate the estimation error of the treatment effect.
+We plot the estimation error box plots in Figure 2 with medians and means, indicate that our proposed method achieves smaller estimation errors.
 Moreover, the estimation error declines as $G$ increases.
-Next, we consider another set of simulation studies under conditions similar to those above, but with the following changes: $J\in\{1,5,10,15,20,\dots,35,40,45,50\}$, $G\in\{2,3,5,10\}$, $T_{0}=30$, $T_{1}=1000$, and $K=5$.
+Next, we consider another set of simulation studies under conditions similar to those above, but with the following changes: $J\in\{1,5,10,15,20,25,30\}$, $G\in\{2,5,10\}$, $T_{0}=30$, $T_{1}=1000$, and $K=5$.
 All other settings remain as in Section 7.1.
-We again apply DMSCM and Abadie $1000$ times and compute the error of the treatment effect estimation (Figure 3).
+We again apply MMSCM and Abadie $100$ times and compute the error of the treatment effect estimation (Figure 3).
 We focus on how the mean squared estimation error changes as $J$ grows.
@@
 We measure the discrepancy between the true and estimated distributions using the Maximum Mean Discrepancy (MMD, Gretton et al., 2012), a probability metric based on kernel embeddings.
-In the top panel of Figure 4, we show how MMD varies with $J$.
+In the top panel of Figure 4, we show how MMD varies with $J\in\{1,5,10,15,20,\dots,35,40,45\}$.
 As before, our proposed methods outperform the method of Abadie (2002) in this setting.
@@
 As $G$ grows (i.e., as more moment conditions are imposed), the predictive performance improves.
+Note that in this experiment, we included the results of Abadie et al.
+(2010) for reference only, since it is not a method for distributional SCM.
+Our focus is on the gains induced by choosing a larger $G$.
 Our observations may connect with findings by Spiess et al. (2023), who examine high-dimensional SCMs. Exploring these connections will be part of our future research.
@@
 Figure 6: Counterfactual distributions in empirical analysis.
-Table 2: Results of treatment effect estimation and conformal inference.
-Estimated treatment effect $\widehat{\tau}_{0,t}$ and $90\%$ confidence interval $(\underline{\tau}_{0,t},\overline{\tau}_{0,t})$.
-Terrorist conflict in Basque country.
-Abadie DMSCM
-$\gamma$- $2$ $5$ $10$ $100$
-$\widehat{\tau}_{0,t}$ $-0.800$ $-0.845$ $-1.054$ $-0.279$ $-0.138$
-$\overline{\tau}_{0,t}$ $0.522$ $0.699$ $0.495$ $0.521$ $0.578$
-$\underline{\tau}_{0,t}$ $-2.587$ $-2.150$ $-2.605$ $-2.616$ $-2.590$
-Tobacco control program in California.
-Abadie DMSCM
-$\gamma$- $2$ $5$ $10$ $100$
-$\widehat{\tau}_{0,t}$ $-26.148$ $-27.744$ $-32.306$ $-35.080$ $-36.879$
-$\overline{\tau}_{0,t}$ $8.698$ $7.852$ $8.898$ $9.544$ $11.748$
-$\underline{\tau}_{0,t}$ $-36.798$ $-29.268$ $-34.783$ $-37.308$ $-35.243$
-Reunification of Germany.
-Abadie DMSCM
-$\gamma$- $2$ $5$ $10$ $100$
-$\widehat{\tau}_{0,t}$ $-1492.374$ $-1253.616$ $-1419.120$ $-1324.448$ $-510.765$
-$\overline{\tau}_{0,t}$ $2635.583$ $-1008.822$ $-591.420$ $-716.376$ $2327.944$
-$\underline{\tau}_{0,t}$ $-2390.413$ $-2738.232$ $-1989.323$ $-2018.878$ $-2625.129$
 7.2 Empirical Analysis
-We now present empirical studies based on Abadie and Gardeazabal (2003), Abadie et al. (2010), and Abadie et al. (2015).
+We now present empirical studies based on the datasets and settings in Abadie & Gardeazabal (2003), Abadie et al.
+(2010), and Abadie et al.
+(2015).
+In this analysis, we also compare our method with the prediction intervals for SCMs proposed by Cattaneo et al.
+(2021), which provide a robust uncertainty evaluation for SCMs.
+Although this prediction interval cannot be applied to our method due to differences in the underlying models, it serves as a promising alternative to conformal prediction.
+Therefore, we include its results for reference.
+Note that while we used our own implementation for our method, Abadie’s method, and DiSCo, we used the pcsi library for the prediction intervals, as published by Cattaneo et al.
+(2022).
 Terrorist conflict in the Basque Country.
-Abadie and Gardeazabal (2003) investigate the effect of terrorist conflict in the Basque Country on gross domestic product (GDP).
+Abadie & Gardeazabal (2003) investigate the effect of terrorist conflict in the Basque Country on gross domestic product (GDP).
 In this example, $J=16$, $K=14$, and $(t_{0},T_{0},T_{1})=(1955,1970,1997)$.
 Basque Euskadi Ta Askatasuna (Basque Homeland and Liberty; ETA) was founded in 1959 as a Basque separatist organization in Spain, seeking an independent Basque state primarily through terrorist activities in the Basque Country region.
-Abadie and Gardeazabal (2003) use annual Spanish regional-level panel data from 1955 to 1997, with a particular focus on the 1970s when ETA’s terrorist activity was significant.
+Abadie & Gardeazabal (2003) use annual Spanish regional-level panel data from 1955 to 1997, with a particular focus on the 1970s when ETA’s terrorist activity was significant.
 The dataset includes fifteen years of pre-intervention observations.
@@
 Abadie et al.
-(2010) study the effect of a large tobacco control program adopted in California in 1988 (described in the Introduction).
+(2010) also investigate the causal effect of Proposition 99, a comprehensive tobacco control program adopted by California.
+They use annual state-level panel data from 1970 to 2000.
+The program was implemented in January 1989, providing eighteen years of pre-intervention observations.
+Four states that adopted large-scale tobacco control programs (Massachusetts, Arizona, Oregon, and Florida) and states that raised cigarette taxes by at least \$0.50 between 1989 and 2000 (Alaska, Hawaii, Maryland, Michigan, New Jersey, New York, and Washington) as well as the District of Columbia are excluded from the set of 38 control states.
+The outcome of interest is annual per capita cigarette sales in packs; covariates include the average retail price of cigarettes, log of state income per capita, the percentage of the population aged 15–24, and per capita beer consumption, all averaged between 1980 and 1988.
 Here, $J=20$, $K=5$, and $(t_{0},T_{0},T_{1})=(1970,1989,2000)$.
@@
 In this example, $J=8$, $K=31$, and $(t_{0},T_{0},T_{1})=(1960,1990,2003)$.
-Separately, Abadie et al.
-(2010) also investigate the causal effect of Proposition 99, a comprehensive tobacco control program adopted by California.
-They use annual state-level panel data from 1970 to 2000.
-The program was implemented in January 1989, providing eighteen years of pre-intervention observations.
-Four states that adopted large-scale tobacco control programs (Massachusetts, Arizona, Oregon, and Florida) and states that raised cigarette taxes by at least \$0.50 between 1989 and 2000 (Alaska, Hawaii, Maryland, Michigan, New Jersey, New York, and Washington) as well as the District of Columbia are excluded from the set of 38 control states.
-The outcome of interest is annual per capita cigarette sales in packs; covariates include the average retail price of cigarettes, log of state income per capita, the percentage of the population aged 15–24, and per capita beer consumption, all averaged between 1980 and 1988.
-Results. Table 2 reports the estimated treatment effects and 90% confidence intervals. Figures 6 and 6 show the estimated treatment effects and the distributional treatment effects, respectively.
-In each dataset, we plot the counterfactual outcomes and distributions in Figures 6 and 6.
-We also perform conformal inference on the sharp null hypothesis $H_{0}:\tau_{0,(T_{0}+1)}=\cdots=\tau_{0,T_{1}}=0$ for the treatment effect.
-For both Abadie’s method and DMSCM, we present the 90% confidence interval of the treatment effect in Table 2.
+Results.
+For each dataset, Figures 5 and 6 show the estimated treatment effects and the distributional treatment effects, respectively.
+In Appendix C, Tables 24 report the estimated treatment effects and 90% confidence bands obtained via conformal inference.
+We also plot the confidence bands in Figures 79.
 Our proposed method fits the observed outcomes well in the pre-treatment period ($t\in\mathcal{T}_{0}$) and, accordingly, produces plausible counterfactual outcomes and distributions.
-In the conformal inference analysis, our method does not reject the null hypothesis in the same manner as Abadie’s estimator, except for the DMSCM in the California tobacco control example.
+In the conformal inference analysis, when we set the sharp null hypothesis as $H_{0}:\tau_{0,(T_{0}+1)}=\cdots=\tau_{0,T_{1}}=0$ for the treatment effect, none of the methods reject the null hypothesis, except for the “California tobacco control” dataset.
+Prediction intervals proposed in Cattaneo et al.
+(2021) shows tight intervals for every settings.
+Note that we cannot directly compare the performance of their method and ours due to differences in the underlying models.
 8 Conclusion
-We have highlighted the problem of implicit endogeneity in the linear regression models commonly employed by existing SCMs.
-To address this issue, we proposed assuming mixture models among outcome distributions and developed a density-matching SCM.
-By matching densities, our method consistently estimates SC weights and reinforces the theoretical foundation for using SCMs in comparative studies.
+This study investigated the asymptotically unbiased estimation of the treatment effect in the setting of SCMs.
+We first pointed out the problem of implicit endogeneity in the linear regression models commonly employed by existing SCMs, which can be interpreted as a form of measurement error bias.
+To address this issue, we focused on distributional SCMs, in which we assume mixture models for the outcome distributions and developed a moment-matching SCM.
+By matching the moments between the outcome of the treated unit and the weighted sum of the outcomes of the untreated units, our method consistently estimates SC weights and reinforces the theoretical foundation for using SCMs in comparative studies.
 We demonstrated the effectiveness of this approach through both simulation experiments and empirical applications.
 Appendix A Proof of Theorem 5.8
-First, we confirm the basic facts about the uniqueness of the CHF.
-Proposition A.1 (Theorem 26.2 in Billingsley (1995)).
-Distinct measures cannot have the same characteristic function.
-Let us define
-$\widehat{Q}_{G}(\bm{w})\coloneqq\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right)\right|,$
-$Q_{G}(\bm{w})\coloneqq\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|.$
+This section provides the proof for Theorem 5.8. Before presenting the proof, we introduce some existing results related to the moment problem and additional notations.
+Moment problem.
+The question of whether a sequence of moments uniquely determines a distribution is known as the moment problem (Schmüdgen, 2020).
+In particular, when the random variable is bounded, the problem is referred to as the Hausdorff moment problem.
+For this case, Hausdorff (1921) provides the following result: if a random variable has support on $[0,1]$, then the sequence of its moments is uniquely determined by its distribution.
+Proposition A.1 (Hausdorff (1921), Satz III).
+Let $\{\mu_{n}\}_{n\geq 0}$ be a real sequence. Define the forward differences recursively by
+$\Delta\mu_{n}\coloneqq\mu_{n+1}-\mu_{n},\qquad\Delta^{k}\mu_{n}\coloneqq\Delta(\Delta^{k-1}\mu_{n})\quad(k\geq 1).$
+The following are equivalent:
+  • (Moment representation) There exists a finite positive Borel measure $\nu$ supported on $[0,1]$ such that
+    $\mu_{n}=\int^{1}_{0}u^{n}d\nu(u)\qquad(n=0,1,2,\dots).$
+  • (Complete monotonicity) The sequence is completely monotone, that is
+    $(-1)^{k}\Delta^{k}\mu_{n}\geq 0\qquad\text{for all }k,n\in\mathbb{N}.$
+Moreover, when these conditions hold, the representing measure $\nu$ is unique.
+Additional notations. Let us define
+$\widehat{Q}_{G}(\bm{w})\coloneqq\sum_{\gamma\in\{1,2,\dots,G\}}v_{\gamma}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big(Y^{N}_{j,t}\Big)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big(Y^{N}_{j,t}\Big)^{\gamma}\right)\right|,$
+$Q_{G}(\bm{w})\coloneqq\sum_{\gamma\in\{1,2,\dots,G\}}v_{\gamma}\left|\mathbb{E}_{0,t}\left[\Big(Y^{N}_{j,t}\Big)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big(Y^{N}_{j,t}\Big)^{\gamma}\right]\right|.$
+Proof. Then, we prove Theorem 5.8 as follows:
 Proof.
 First, from the definition of $\Delta^{J}$, $\Delta^{J}$ is a compact parameter space.
-Next, $\widehat{Q}(\bm{w})$ converges uniformly to $Q(\bm{w})$ as $T_{0}\to\infty$ under Assumption 6.1.
+Next, $\widehat{Q}(\bm{w})$ converges uniformly in probability to $Q(\bm{w})$ as $T_{0}\to\infty$ under Assumptions 5.4 and 5.7 (Pötscher & Prucha, 1989).
 Furthermore, $Q(\bm{w})$ is continuous with respect to $\bm{w}$.
@@
   1. 1.
-    $\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>\widehat{Q}(\bm{w}_{0})-\varepsilon/3$;
+    $\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{MMSCM}}\right)<\widehat{Q}(\bm{w}_{0})+\varepsilon/3$;
   2. 2.
-    $\widehat{Q}_{G}(\widehat{\bm{w}}^{\mathrm{DMSCM}})>\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)-\varepsilon/3$;
+    $Q_{G}(\widehat{\bm{w}}^{\mathrm{MMSCM}})<\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{MMSCM}}\right)+\varepsilon/3$;
   3. 3.
-    $\widehat{Q}_{G}\left(\bm{w}_{0}\right)>Q_{G}(\bm{w}_{0})-\varepsilon/3$.
+    $\widehat{Q}_{G}\left(\bm{w}_{0}\right)<Q_{G}(\bm{w}_{0})+\varepsilon/3$.
 Therefore, with probability approaching one,
-$Q_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)-\varepsilon/3>\widehat{Q}_{G}\left(\bm{w}_{0}\right)-2\varepsilon/3>Q_{G}\left(\bm{w}_{0}\right)-\varepsilon.$
-Thus, for any $\varepsilon>0$, $Q_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>Q_{G}\left(\bm{w}_{0}\right)-\varepsilon$ with probability approaching one.
+$Q_{G}\left(\widehat{\bm{w}}^{\mathrm{MMSCM}}\right)<\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{MMSCM}}\right)+\varepsilon/3<\widehat{Q}_{G}\left(\bm{w}_{0}\right)+2\varepsilon/3<Q_{G}\left(\bm{w}_{0}\right)+\varepsilon.$
+Thus, for any $\varepsilon>0$, $Q_{G}\left(\widehat{\bm{w}}^{\mathrm{MMSCM}}\right)<Q_{G}\left(\bm{w}_{0}\right)+\varepsilon$ holds with probability approaching one.
 Let $\mathcal{N}$ be any open subset of $\Delta^{J}$ containing $\widetilde{\Phi}^{*}$.
-Since $\Theta\cap\mathcal{N}^{c}$ is compact, and $Q_{G}(\bm{w})$ is continuous with respect to $\bm{w}$, we have
-$\sup_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w})=Q_{G}(\bm{w}^{\dagger})<Q_{G}(\bm{w}_{0})$
-for some $\bm{w}^{\dagger}\in\Theta\cap\mathcal{N}^{c}$. Thus, choosing $\varepsilon=Q_{G}(\bm{w}_{0})-\sup_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w})$, it follows that
-$Q_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>\sup_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w}),$
-and hence $\widehat{\bm{w}}^{\mathrm{DMSCM}}\in\mathcal{N}$, where recall that $\mathcal{N}$ is any open subset of $\Delta^{J}$ containing $\widetilde{\Phi}^{*}$.
-Thus, $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}$ holds as $T_{0}\to\infty$.
+Since $\Delta^{J}\cap\mathcal{N}^{c}$ is compact, and $Q_{G}(\bm{w})$ is continuous with respect to $\bm{w}$, we have
+$\inf_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w})=Q_{G}(\bm{w}^{\dagger})>Q_{G}(\bm{w}_{0})$
+for some $\bm{w}^{\dagger}\in\Delta^{J}\cap\mathcal{N}^{c}$. Thus, choosing $\varepsilon=Q_{G}(\bm{w}_{0})-\inf_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w})$, it follows that
+$Q_{G}\left(\widehat{\bm{w}}^{\mathrm{MMSCM}}\right)<\inf_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w}),$
+with probability approaching one.
+This implies that $\widehat{\bm{w}}^{\mathrm{MMSCM}}\notin\Delta^{J}\cap\mathcal{N}^{c}$, and hence $\widehat{\bm{w}}^{\mathrm{MMSCM}}\in\mathcal{N}$, where recall that $\mathcal{N}$ is any open subset of $\Delta^{J}$ containing $\widetilde{\Phi}^{*}$.
+Therefore, $\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}$ holds as $T_{0}\to\infty$.
 Additionally, from Proposition A.1, we have $\bm{w}^{*}_{G}\to\bm{w}^{*}\in\widetilde{\Phi}^{*}$ as $G\to\infty$.
 ∎
+Appendix B Proof of Theorem 5.12
+Proof.
+Each draw $Y^{l,\dagger}_{0}$ is independent and distributed according to $\widehat{F}_{Y^{N}_{0,t},T_{0}}(y)$. Therefore, for any fixed $T_{0}$, we have
+$G_{T_{0},L}\xrightarrow{\mathrm{a.s.}}\widehat{F}_{Y^{N}_{0,t},T_{0}}(y)$
+as $B\to\infty$, from the Glivenko–Cantelli theorem.
+Next, from Theorem 5.8, $\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}$ holds as $T_{0}\to\infty$.
+Lastly, $F_{Y^{N}_{j,t},T_{0}}(y)$ converges uniformly in probability to $F_{Y^{N}_{0,t}}(y)$ as $T_{0}\to\infty$ under Assumptions 5.4 and 5.7 (Pötscher & Prucha, 1989).
+By combining above results, we have
+$\sup_{y}\big|G_{T_{0},L}(y)-F_{Y^{N}_{0,t}}(y)\big|\xrightarrow{\mathrm{p}}0\quad(L\to\infty,n\to\infty),$
+Appendix C Confidence bands
+In Tables 24 and Figures 79, we present the results of the conformal inference analysis described in Section 7.2. We omitted MMSCM ($G=100$) due to the space limitation.